A Bibliography of Publications in the Journal of FisheriesSciences.com

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

18 January 2025 Version 0.04

Title word cross-reference

23 ± 4 [Alp13]. ^{+2} [Bah07]. ^{210} [DFDRMGT17]. [ÖÖK07]. _{2} [Bah07]. α [SK13]. × [Ünv12]. [Erk13].

-3 [Erk13]. -Methyltestosterone [SK13].

icerikleri [Pol12]. icin [Yeş13]. iki [Aka11b, Dağ08]. iklim [Oğu11]. ile [Özd11, Öİ08, Yeş11]. İlhan [Baş11b, BSİU12]. ili [YÖ10]. iliskileri [PiYB09]. iliskisi [Aka11b]. immunostimulanların [Erg12]. ’in [Oğu11]. İnceburun [Gön10]. incelenmesi [Çol14]. isgalci [Dev11]. iskenderun [Pol12, AKÖ09, Aka10, Dem08b, Duy08, Özy09, Pol08b]. iskorpit [Aka11b]. isleminin [Alp13]. isleme [MU10]. isletmelerinin [BAU^{+}12]. istavrit [Atı12]. İzmir [Tad09, Aka11b, Taş11, Tay11].

/H [Bah07].

12th [Tiw19]. 1810’nın [Taş11]. 19 [MM21].

2021 [Cha21]. 21st

[VdV16, vGdVW16, vGdV16].

353-nonylphenol [Kle13].

4 [ÖÖK07]. 4-hexylresorcinol [Sel14].

5ec [Gün09].

A. [Ünv12]. Abeokuta [AF11]. Ability [BS18]. Abnormal [Dağ08]. Above [Ayd08a]. Abundance [Bra20, TM16]. Acari [Fer14]. According [ÇKB^{+}13]. Accumulation [Gün08a]. Ache [Çap11]. Acid [Çağ11, Har13, Kay14a, MK14, MKS17, NLM16, Öks10, Öks12, PÇD^{+}15b, Sid13, Gül07, Özp10, PÇD^{+}15a, Tur07, Yıl10]. Acidification [MKK15]. Acids [Kay08, Khi11, Dun10, NCA13]. Acinetobacter [Yon10]. Acipenser [Akb13]. Acisindan [Sür11]. ACP [FE15]. Actinopterygii [Özv14b]. Activities [AEU18, Naz12a, Yab09a, UACÖ13]. Activity [Bul12, DFDRMGT17]. Acute [Gün08b, Küç08]. Adana [Çev08, Erg10, EG12, Gel09, Özo10, Şah09, Sib08, Usl11, Oğu11]. Adaptation [Kar10, SKB^{+}18]. Additive [SAA^{+}17]. Additives [ASTH21, KST16, Yil14]. adhesion [KAA^{+}17]. Adigüzel [BSİU12, Baş11b, BSİU12]. Adiyaman [Bay12, Bay12]. Administration [Baş09]. Adult [ZPW^{+}17]. aduncus [LWP^{+}16]. Advances [Sin15]. Advantages [BMK17]. Aegean [Aky11a, Aky11b, Gür12, İşi09, Özc09, Şen11, Tan12, Tos07, Yab09a, AS10, Fıl09]. Aeration [Usl12b]. Aerobic [Kub09]. Aeromonas [DVSS12, Str14]. aestuarii [Özc09]. affecting [DED^{+}10]. affinis [Gün09, Gün10]. Africa [LRMO19]. African [Ade09, AEU18, GAN17, KG16, OKA16, OAE^{+}18, SIAF16, Tur11, UEA18]. Afrin [Dal08]. After [Baş09]. Afyonkarahisar [Açı13, Özt13]. against [Bra19, JK16]. Agar [MP17]. Agar-Digesting [MP17]. Agardh [CŞA10]. Age [Ayd09, Fıl09, Gür11, Pol08a, PiYB09, Bos08, Bos09, PiYB09]. ageing [YPY07]. ages [Bos07]. Agır [KMK09]. Agırlık [Aka11b, PiYB09]. aglarin [BG10]. Aglarinda [ŞHG^{+}08]. Agriculture [ACSHvdH17]. ait [Tay11]. Akdeniz [Kes12a, ÖL08, Pol12]. akdeniz’de [BG10].

Akobo [TWT19]. Aktivitesine [Çap11]. Akyatan [Çev08, Gel09, Özo10]. Al [HSAE16]. Al-Arab [HSAE16]. Alabaligi [YÖ10, Yiğ11b]. alabalıgın [Bau13]. Alabalik [BAU^{+}12]. Alabaliklarin [Çap11]. Alanindaki [Kuş13]. albacares [Gar20]. Albatinah [Alh12]. Alberta [RGB^{+}16]. Algae [ÇKAE10, JK16, Kar07, MKS17, OAA15, CŞA10, SYG^{+}08, Soy10, Taş07b]. Algeria [BMKS20]. Algicidal [JK16]. Alginate [Naz12b]. Algleri [ÇKAE10]. Alien [Özv13b]. Alimentation [AKD^{+}17]. Alkaline [BMBS17]. allometric [Çob09]. Alloy [BMK17]. Along [Kes12b, RGO13]. Alosa [CFFM16, Duy12]. Alterations [PMOBF18]. Alternative [Erk14, Gök09, Meg14]. Altinkaya [YPY07, YPY07]. altivelis [Ayd10b]. amarus [Lha14]. Amazon [Mat14]. Amazonas [Lem12]. Ambalaj [Sür11]. Amberjack [Öks12]. Ambient [Alp13]. American [LZ21]. Amerti [Tek15]. amet1601 [RPCC15]. Amino [Çağ11, MK14, Sid13]. amnicola [GC16]. Among [Kes12b, BA16]. Amount [Gel09]. amphioxys [Kut10]. Amphipoda [Taş10b]. Amphiprion [KTT16]. Anabas [RCR^{+}16]. Anaesthetics [AEU18]. analis [MMdA^{+}15]. Analizi [Öİ08]. Analysis [Cha20, HBPH19, HBP^{+}18, KLPK16, LWP^{+}16, Öİ08, RBT^{+}17, SPC^{+}17, YG08a, Zla17, Doğ09b, Doğ10, Mar11, Uzm08]. Anambra [IEN^{+}19]. Anatolia [GUA14]. Anchovy [Kay08, UAH^{+}07, Bat07, İÖÇK10, Lev11]. Androgen [Yue18]. androstenedione [Tur11]. Androstenedionun [Tur11]. anesthetic [UACÖ13]. Angel [Ürk13a]. Anglers [KBD17]. angorae [GUA14]. Anguilla [Deg16, Rad13, SCEA08]. anguillarum [Did13]. Animal [SMJO15]. Annular [Ayd14a]. annularis [Ayd14a]. anormal [Dağ08]. Antalya [Özv14b, Gök14, Yel11, Yeş11]. Antalya/Turkey [Gök14]. antennarius [Özc13]. Antibacterial [TYB^{+}13]. Antibiotics [AF11]. Antinutritional [Doğ09a]. Antioxi [Yon14]. Antioxi-dant [Yon14]. antioxidan [UACÖ13]. Antioxidant [Kay14a]. API [Özo10]. Aplysiidae [Özv14a]. Apolyont [ÇKB^{+}13, Ber09]. Application [AKV^{+}17, ANM19, Alt13, BL17, Dur10, RSSG16]. Applications [MMdA^{+}15, Mol11b, NS17, Kil11, KDED11]. Aqua [VGP^{+}17]. Aquaculture [Akb09, Ayd10a, BMK17, BAU^{+}12, Cha20, Dem08a, FE15, KDED11, LZ21, LRMO19, Nik15, Ram17, SKB^{+}18, SBW^{+}17, SMG^{+}18, Sin15, Taş07a, Tiw19, Yab09a, Meg14, Tos13, Cha21]. Aquafeeds [Yil14]. Aquaponic [SBL^{+}18]. Aquaponics [PKMB17]. Aquaria [Opi14]. aquarium [Kar11]. Aquatic [OF18, PRI16, Taş12, Ade12, AAÖ11, KMK09, Kay10, Top09, Uzm08]. Arab [HSAE16]. Arabian [Meh13]. Arabibarbus [PÇD^{+}15a, PÇD^{+}15b]. araclarının [Bay12]. Arakanga [AF11]. aral [Abu13]. Arasındaki [Kes12a, Bau13]. Arastirilmasi [ŞHG^{+}08, Yük11, Açı13]. Archaeogastropoda [Aya10]. Archipelago [FA14]. Area [Gül09, MM21, BDAK10]. Areas [Özt12]. Arel [RSSG16]. Argulus [Açı13, Özt13]. Argyrosomus [Par13]. aromatic [Ade12]. Arrenurus [Fer14]. arriving [Cha21]. Artemia [RPCC15]. Arunachal [AKD^{+}17, DDAD17, KAGD16, AKDD17]. Asa [OF18]. Ascidiacea [Dın10]. Asetilkolinesteraz [Çap11]. asetic [Özp10]. Asi [Dem09, Boz10b, DO15]. Asidi [Erk13]. Asiti [Pol12]. Aslantas [Boz09b, Boz10a, Boz09a]. asota [Mat14]. Aspects [Abt14, OOB16, RSA^{+}21, RCR^{+}16, BDAK10]. aspersa [Çağ11]. Assam [DAM18, DADL17, GDAD16, SAMN16, Abu13]. assemblage [RGO13]. assesment [Ayd10a]. Assessing [RSSG16]. Assessment [ATAA17, AB15, AASASG^{+}15, Bra19, Has14, PCV^{+}16, Tar10, ÜBİG07, Ade12]. Asso [Par13]. Associated [Jal14, Kor08, MP17]. associates [Doğ10]. assurance [Erk08]. Astacidae [DED^{+}10]. Astacus [Ber09, Yük11]. Astaxanthin [Usl12b]. Astronotus [Mat14]. Asymmetry [Jaw12b]. Atatürk [Bay12, Bay12, DAS14]. Atherina [ELKESA19, Gür12, Taş11]. Atlantic [RCV^{+}17, Dun10, Man08, NCA13]. Atrazine [GAN17]. Attention [PC16]. Attributes [ASTH21]. aurata [AKÖ09, Aka10, Yt12]. auratus [Özt10a]. Autochthonous [BL17]. Av [Özd11, ŞHG^{+}08, Bay12]. Avciligin [ÇKB^{+}13]. avlanabilir [Gön13]. Avlanan [Özd11, Aks11a]. Avlanilan [Yeş11]. Ayancik [Kar07]. Aylik [Yeş11]. Ayu [Ayd10b].

Bir [Çan11]. Birden [Alp13]. Biyolojisi [Kuş13]. Babadillimani [Man08]. Bacillus [AKV^{+}17, MP17, RPCC15]. Background [Opi14]. Bacoor [SRB^{+}15]. Bacteria [AF11, BL17, Did13, LL07, NLM16]. Bacterial [AOO16, Kub09, LL08, Özo10, RPCC15, TYB^{+}13, Ürk13a]. Bacteriocin [AKV^{+}17]. Bacterium [DVSS12]. BAD [Gül07]. BAD-UTF-8 [Gül07]. Bafra [Soy10, YYBP07]. Bagridae [DDAD17]. Baird [Gün09, Gün10]. Bakteriyel [Ürk13b]. Baligi [YYBP07, YPY07, Alp13, Taş11, Ayd11d]. Baligi’nin [ÖL08]. Balik [BSİU12, ÇKB^{+}13, İSU14, Yeş11, Yiğ11a, YYBP07, Baş11b, Erg12, Bau13]. Balikciligi [Yeş11, BG10, Bay12]. Balıkcılıgında [Ayd14a]. Balıkesir [İSU14, İSU14]. Baliklar [AC08]. Baliklarin [AKA11a, Gön13]. Balıklarında [Ürk13b]. Baliklarinin [Özd11, Aks11a].

Balkaya [Uzm08]. Balkaya-Kırklareli [Uzm08]. Balls [KÖÇ12]. Baltic [Dev11]. baltık [Dev11]. Banana [Law14]. Bangladesh [RAB^{+}17, Akt14, HBPH19, HBP^{+}18, Sid13]. Banks [Ero09]. Baraj [Baş11b, BSİU12, Öİ08, YPY07, Açı13, Bay12, Çob11, Yük11]. Barb [DAM18]. barbata [CŞA10]. barbatus [Aks11b, Aks11a, Özd11, Ayd14a]. Barbel [BSİU12]. Barbell [ZPW^{+}17]. Barbunya [Aks11a, Ayd14a, Özd11]. Barbus [DAS14]. barroisi [DO15]. Bartail [Has14]. Based [MKS17, Pep21, Zla17, vGdV16, ATAA17, Dur10, NS17]. Basil [SBL^{+}18]. Basin [DO15, KAGD16, Özc13, ÖTGG07, Saç14]. Bass [Alp13, Jal14, Kor08]. Bata [PRI16]. Batasio [DDAD17]. batatas [SOO15]. Batı [Çol14]. Bay [AKÖ09, Aka10, Aky11b, Dem08b, Duy08, Gür12, Pol08b, Şen11, SRB^{+}15, Tan12, Taş07a, Aya11a, Aya11b, Ayd14b, Dın10, Özy09, Yel11]. Bayes [Aka13, Sed14]. Bazar [Akt14]. Bazi [BSİU12, KÖÇ12, Atı12, Baş11b, KMK09, Pol12, Aks11a]. Be [G16, BA16, Cha21]. Beacons [ATEA^{+}16]. Beel [GDAD16]. Behaviors [NLM16]. Behaviour [Baş09]. Being [Bay12]. Belirlenmesi [Duy12, KÖÇ12, YPY07]. belone [KÖÇ12, KÖÇ12, PiYB09]. belonging [CFFM16, Kar10]. beluga [Akb10]. benekli [Bau13]. bengalensis [Jaw12b]. Benin [SIAF16]. Benthic [MG16, SYG^{+}08]. Benue [OOB16]. berolinensis [Fer14]. Beslenme [YYBP07]. Beslenmesi [Şen13b]. Besleyici [Usl12a]. Beta [SBW^{+}17]. Beta-1-3-Glucan [SBW^{+}17]. Between [Bau13, Jaw12a, Tar08, RCV^{+}17]. bicirrhosum [Lem12]. Bigeye [RCV^{+}17]. Bight [Man08]. Billion [vGdV16]. Bio [OF18]. Bio-indicators [OF18]. Bioaccumulation [GC16, Gün10, IEN^{+}19, SRB^{+}15]. Biochemical [Naz12b, Öks08].

Biodiversity [AİGS11]. biofloc [Meg14]. Biological [ATEA^{+}16, Tar10, BDAK10, KMK09, Kay10]. Biology [Kuş13, Nik15, OOB16, RCR^{+}16, Ünv12]. Biomarker [Har15]. Biomarkers [OAE^{+}18]. Biomass [Gön10]. Biomedical [SMJO15]. Biometric [Gür12, Rad13, Tar08]. biopolymer [Dur10]. Biopreservation [AKV^{+}17]. Bird [Cha14]. Bitterling [Lha14]. Biyikli [BSİU12, Baş11b]. biyokütle [Gön13]. Biyolojik [KMK09]. biyolojisi [Ünv12]. Biyometrik [Taş11]. Blaasop [Özv14b]. Black [Akb12, Ayd11a, Ayd11b, Ayd11c, Baç12, Baş11a, Fıl09, Özt12, ŞHG^{+}08, Şah10, Taş07a, YKŞ15, Akb09, Akb10, Aks11b, Bat07, Dev11, Erd10, Gön10, Lev11, Pol08a, PiYB09, Tur07, UAH^{+}07, Zla17]. Blainville [Özv13a, Özv14a, Tan12]. Bleeding [YOH^{+}16]. Bloch [Abu13, Bos07, Lha14, YYBP07]. Blood [ATG^{+}14, OAE^{+}18, Kel10]. Blooms [CGZ21]. Blue [AAP^{+}17, Aya11a, Aya11b, Gel09, GC16, Gül09, Kes12b, Oeh08, Özo10, TMGY16]. Bluefin [DÖD^{+}13]. board [Özk10, Özk14a]. Boat [BJSN19, Doğ09b]. Bodo [GC16]. Body [DAS^{+}16, Köp12, Lem12, MKS17, PCV^{+}16, Tar08]. Bölgesinde [ŞHG^{+}08]. bolluk [Tay11]. Bone [Tar08]. Bonito [Öks08]. bony [Bos07, Bos08, Bos09, Pol08a, YPY07]. Bottom [Ayd08a, Ayd14a, Ayd14b, Gön10, Yel11]. box6 [Kle13]. Boy [Aka11b, Ayd14a, Özd11, PiYB09]. Boy-Agırlık [Aka11b, PiYB09]. boyeri [Gür12, Taş11]. Boylamanin [Dik11]. boyu [Bau13]. boyutları [Bau13]. boyutları-balık [Bau13]. Bozcaada [Aky11a]. Brackish [MUK17]. Brahmaputra [DADL17]. Brain [EN17]. Branchiura [Açı13, Özt13]. Brazil [Lem12]. Breaking [YOH^{+}16]. Bream [Ayd14a, Yt12, AKÖ09, Aka10]. Broken [Ayd09]. Brood [Sim14]. Brook [ACSHvdH17, Dal08]. Brown [BDB20, CŞA10, Sim14]. Buccal [AF11]. Bulgaria [Zla17]. Buoys [KKP16]. Burchell [AEU18, AAP18, KAA^{+}17, Law14, OKA16, Tur11]. Burger [ELKESA19]. Burnt [Ayd09]. Bursa [Ber09, Doğ09b]. Bursa/Turkey [Doğ09b]. Bursatella [Özv14a, Tan12]. Bushehr [RGO13]. Business [BA16]. Businesses [BA16]. Büyüklügünün [Yük11]. büyüklükleri [Gön13]. Büyüme [Baş11b, BSİU12, Çob11, Çol14, Öİ08]. Büyümesi [Yiğ11b]. Büyümesi [ÖL08]. Büyümesine [Tür09]. Buzdolabı [Duy12]. By-products [G16]. Bycatch [ŞHG^{+}08].

C [Oğu11, Alp13, Gül07, Öza09, ÖÖK07, Usl11]. C-Fikosiyanin [Oğu11]. C-phycocyanin [Usl11]. C. [CŞA10]. Caddisflies [DB08]. Cadmium [Utk08]. Caglarca [YÖ10, YÖ10]. Calamary [Öza09]. Calcitrans [RPCC15]. Calcium [Law13, SOA^{+}15]. calisma [ÖL08, ŞHG^{+}08]. Callinectes [Aya11a, Gel09, GC16, Gül09, Kes12a, Kes12b, Özo10, TMGY16]. Can [BM17, Ber17b, Gar20, Ber15]. Candidate [Did13, SBW^{+}17, VGP^{+}17]. candir [BAU^{+}12, BAU^{+}12]. Canned [Gar20, MK14, Oeh08]. canola [Der10]. Capilaria [Ürk13b]. Capillaria [Ürk13a]. capoeta [SGH^{+}16, Ayd12, Çob11, DO15, SGH^{+}16]. Captive [AKDD17, DÖD^{+}13, KTT16]. Captivity [Şen13a]. Captured [Gel09, Akb10, NCA13]. Carangid [Jaw12a]. Carangoides [Jaw12a]. Carapace [Aya11a]. Carassius [Bos07, Özt10a, Tar08, Utk08, YYBP07]. Carcinus [Özc09]. Cardiovascular [Gar20]. Carena [Sağ11]. Caribbean [FA14]. Carinaria [Özv13a]. caring [Evr09].

Carotenoid [KTT16]. Carp [ASTH21, Ayd09, Fro21b, Gök09, HDS^{+}16, Köp12, Küç08, MKS17, STFO15, Yt09, Bos07, YYBP07]. carpio [Ayd09, Kay14a, Küç08, Yt09]. Carps [AK17, Ani14]. Case [AASASG^{+}15, JZ15, KANO21, ML16, AAÖ11]. Caspian [Dev11]. Catch [BDdF19, KBD17, AKÖ09, Aka10, Ayd08b, Ors10b, ÖEAÖ10]. Catchable [Gön10]. Catching [Ayd14b, Ors10a, Özt12, Özy09]. Categories [MUK17]. Catfish [Ade09, AEU18, DTT^{+}16, GAN17, KG16, Law14, OKA16, OAE^{+}18, SAA^{+}17, Tur11, UEA18]. Caught [Dağ08, Gön10, Özo10, Özt12, Şah09, Aks11b, ÖEAÖ10, Yel11]. Cavity [AF11]. cayi [ÇKAE10, İSU14, İS12, Sal11, İpe12]. Cell [Şah09]. Cells [BM17]. Century [VdV16, vGdVW16, vGdV16]. Cephalopod [Duy08]. Cephalopoda [Şen11]. Cephalopods [Şen13b]. cephalus [Açı13, Bah07, OO13, Özt13, Ünv12]. Certain [ATG^{+}14, DADL17]. Ceyhan [Şah09]. cf. [Sim14]. Chaetognaths [Ter08]. Chain [HBPH19, HBP^{+}18, PRI16, Dun10]. Chalcalburnus [YG08a]. chalcoides [Ünv12]. Challenge [FE15]. Chamelea [Çol14]. Change [Dor12, Har15, SKB^{+}18, Yel11]. changed [Bat07, Lev11]. Changes [GAN17, Has15, Usl12a, YOH^{+}16, AST^{+}12, Bil07, CŞA10]. Changing [Har15]. Channel [Akt14]. Characteristics [Aya10, Baç12, Özy09, Rad13, Tos07, Yük12, Ayd11d, Bos08]. Characterization [BMBS17, BL17, Çel10, DNMS13]. characters [Ter08]. Charting [SKK16]. Cheatoceros [RPCC15]. Check [Beh14]. Chemical [Aya11a, Aya11b, Dum12, ELKESA19, İÖÇK10, MK14, NCA13, Özo12, Sid13, ÜBİG07, Bil07, CŞA10, Duy12, Taş10a]. Chilika [Beh14]. China [CGZ21]. Chinook [FGF20]. Chips [Dum12]. Chironomidae [Taş10b, Tad09]. Chitin [MKK15]. Cholesterol [MK14, PÇD^{+}15a, PÇD^{+}15b, NCA13]. Chondrostoma [Erg10]. Chrysicthys [AF11]. Chub [Özt13]. Cichlidae [PMOBF18]. Cichlids [OAO^{+}15]. cinsiyet [Şen14, Tur11]. Circadian [Nik15]. Cirrhinus [BL17]. Cirrihinus [SSG13]. cirrosa [Baş09]. City [Jaw12b]. civril [Yac09]. civril-Denizli [Yac09]. civril-Denizli/Turkey [Yac09]. Cladocera [Boz09a, Sal09]. Cladoceran [NBBR17]. Cladophora [Esr14]. Clam [Çol14]. Clarias [ATAA17, Ade09, AOO16, AEU18, AAP18, GTO15, GAN17, KAA^{+}17, KG16, Law13, Law14, LVA15, OKA16, OAE^{+}18, SAA^{+}17, SOO15, Tur11, UEA18]. clarkii [ANM19]. clavata [Tur07]. Climate [Dor12, Har15, SKB^{+}18]. Climatic [Usl11]. clinoptilolite [Öz10]. Clove [Has15]. Clown [KTT16]. Cnidaria [AİGS11]. co [Doğ10]. co-operatives [Doğ10]. Coast [AAP^{+}17, Beh14, Kes12b, Özv14a, Tos07, Tur07, NPR16]. Coastal [Aky11a, BDdF19, İşi09, Meh13, AKA11a, BDAK10, RGO13]. coastline [Kar07]. Cobalt [Gün10]. Cochlidium [SRB^{+}15]. Cod [LL07]. Codend [Ayd08a]. coeruleolineatus [AAP^{+}17]. coeruleopinnatus [Jaw12a]. Coi [Kes12a, Kes12b]. coindetti [Öza09]. Colanitida [GTO15]. Cold [Fro21a]. Coldwater [Sin15]. Collected [Akt14, Jaw12a, Jaw12b, Sid13]. Color [RCP^{+}16, Tür09]. Colossoma [DAS^{+}16, dMSD^{+}16]. Colour [Öks12, Opi14, Ors10b, Ors10a]. Column [Yab09a]. comb [Dev11]. Combined [Opi14]. Come [VdV16]. Commentary [LZ21, Pep21]. Commercial [ÇKB^{+}13, Gön10, HDS^{+}16, IEN^{+}19, NS17, Naz12a, UEA18, Zla17]. commercialized [NCA13]. Common [ASTH21, Fro21b, Gün14, Küç08, STFO15, AS10, Çob09].

Communities [İşi09, MG16, Ade12]. Community [Alh12, İS13, MG16]. Comparation [Özt12]. Comparative [KY17, NBBR17, OKA16, STFO15, SGH^{+}16]. Compared [Bra19]. Comparision [AOO16]. Comparison [Ayd12, Gün14, NS17, Öks08, Öks12, Ors10b]. Complexes [Dın10]. Composition [Aya11a, Aya11b, Boz09b, Boz10a, ÇKB^{+}13, Dum12, ELKESA19, GD13, Har13, Kay08, Köp12, MK14, MKS17, Öks12, Özo12, ŞHG^{+}08, Sid13, TM16, Usl12a, Ayd08b, Bil07, CŞA10, Duy12, Gül07, İÖÇK10, NCA13, ÖEAÖ10, Tur07, Yıl10]. Compositions [Naz12b, Öks10, PÇD^{+}15a, PÇD^{+}15b]. Concentration [Bul12, Gün09, KBD17, PP16]. Concentrations [Akt14, AAP18, DFDRMGT17, Gün10, HDS^{+}16, MUK17, RCV^{+}17, SCEA08]. Concrete [OKA16]. Condition [DADL17, eGK14, ÖÖK07, SAMN16, SMG^{+}18, TWT19, Kar10]. Conditions [Alp13, Şen13b, Usl11, Duy12, Şen09]. Confirms [CFFM16]. Confronting [ML16]. Conservation [KANO21]. Consortium [RPCC15]. Consumption [Akb12, Dor12, Gar20, Mol08, Yt09]. Contamination [RSSG16]. Content [MKK15, PÇD^{+}15a, PÇD^{+}15b, Usl12a, NCA13]. Contents [Çağ11]. Contribution [AİGS11, BMKS20]. Contributions [Tad09, Top09]. Control [BS18, Did13]. controlled [Şen09]. Cooked [DFDRMGT17]. Cooking [PP16]. Copepoda [Boz09a, Sal09]. Copper [BMK17, Gün08a, Gün08b, Utk08]. Coral [ER15]. Corn [Yt12]. Correction [KTT16]. Cortisol [Bul12]. cory [ÜA08]. Corydoras [ÜA08]. Cost [KG16, UAH^{+}07, HB07]. coubie [OOB16]. Could [BA16]. Cover [Özk14a]. COVID [MM21]. COVID-19 [MM21]. Cox [Akt14]. cözüm [MU10]. cözündürme [Alp13]. CPUEs [Özt12]. Crab [AO19, Aya11a, Aya11b, BDdF19, Gel09, GC16, Gül09, Kes12b, NPR16, Özc09, Özo10, P19, TMGY16]. Craspedacusta [AİGS11]. Crayfish [ANM19, Ber09, Yük12, DED^{+}10]. Creek [BAU^{+}12, GC16]. Crises [vGdVW16]. crispata [Esr14]. Crocodile [LRMO19, Tos13]. Crocodylus [LRMO19]. Crucian [Bos07]. Crude [BMBS17]. Crustacea [Açı13, Özt13, Özu11, Sal09]. Cryopreserved [dMSD^{+}16]. CSI [RSSG16]. Ctenopharyngodon [Köp12]. Cuddalore [Cha14]. cukurova [Usl11]. cUkurovabölgesi [Oğu11]. Cultivable [CPE^{+}15]. Cultivated [ASTH21]. Cultivation [Kuş13]. Cultural [Gan15]. Culture [KANO21, NPR16, OKA16, RPCC15, Şen13b, CŞA10, DED^{+}10, Kar10]. Cultured [Mat14, Öks12, Çan11, Çob09, Yon10]. cultures [VdFHLAA15]. Cumulative [Zoe11]. Cured [NLM16]. cuttlebone [BDAK10]. Cuttlefish [Aky11b, Meh13, Öza09, Özo12, Şen09]. Cuvier [KANO21]. Cyanophyta [Oğu11, Usl11]. Cyhalotrin [Gün09]. Cynoglossus [RSSG16]. Cyprinid [Dağ08, Ünv12]. Cyprinidae [Şah09, Abt14, BMKS20, DO15, YG08a, YG08b]. Cyprinion [YG08b]. Cyprinus [Ayd09, Kay14a, Küç08, Yt09]. Cyprinuscarpio [STFO15]. Cystoseira [CŞA10]. Cytometry [Gün12]. Cytoprotection [EN17].

Dicle [Çob11]. Dip [Yeş11]. Dagılım [Pol12]. Dagilimi [ÇKB^{+}13, AKA11a, Tay11]. Dagılımları [ÇKAE10]. Daily [Ayd10b]. Dakota [FGF20]. Dalyan [Gül09]. Dam [AST^{+}12, Ayd12, BSİU12, Bay12, Boz09a, Boz09b, Boz10a, Dar10, DAS14, Erg10, EG12, Ero09, Gür11, Özt13, Öİ08, PÇD^{+}15a, PÇD^{+}15b, Sal14, Sib08, Taş07b, Yac09, YPY07, Yük12]. Damages [Dev11]. daniconius [SAMN16]. Danio [Kay14b, Koç10, SMJO15]. danrica [DAM18]. dant [Yon14]. Dariören [ÇKAE10, ÇKAE10]. Dark [Öks08, Kab09]. Darlik [GD13]. date [Cha21]. day [Ayd08b]. De-adhesion [KAA^{+}17]. decadactylus [SIAF16]. Decapoda [Özc13]. Deep [KY17, Dun10]. Deep-Sea [KY17, Dun10]. Deerfield [KDCS20]. Deficiency [KTT16, Usl12b]. Deformities [Ber17b, Ber15, Kle13]. Degisimi [Yeş11]. Degerlendirilmesi [Erk13, Kes12a]. degisikliklerin [Usl12a]. degisimleri [Gön13]. Degradation [ÖÖK07]. Delaying [YOH^{+}16]. Delta [DTT^{+}16, GC16, Rad13, Tad09, Ade12]. Delta/Turkey [Rad13]. demented [Evr09]. Demersal [AKA11a, Özd11, Tos07]. Denizatlarinin [Kuş13]. Denizel [Çan11]. Denizi [Taş11, Dev11]. denizi’ne [Dev11]. Denizli [Baş11b, BSİU12]. Denizli/Turkey [Yac09]. Denoised [LWP^{+}16]. Density [AO19, Opi14, AKA11a]. dentex [AS10, Çob09, AS10, Çob09, Naz12a]. Dependence [KAGD16]. depressive [Zoe11]. Derbent [Taş07b]. Dere [BAU^{+}12]. Deresi [İS12, İpe12]. Deresive [ÇKAE10]. Derleme [Çan11]. Designs [Dem08b]. Destructive [SGH^{+}16]. Detecting [ATEA^{+}16]. Detection [TYB^{+}13, ÇKB^{+}13]. Detergent [MMdA^{+}15]. Determinant [Tek15]. Determination [Ayd09, Baş11a, DAS14, Dum12, Dun10, Erk07, Gel09, Gün09, KÖÇ12, Kes12b, MUK17, Öza09, ÖEAÖ10, PÇD^{+}15a, PÇD^{+}15b, PiYB09, Şah09, TYB^{+}13, Bos08, Pol08a, Sel14, Duy12, YPY07]. Determined [Ayd08a]. Deveined [BS18]. Developed [Gar20]. Developing [FE15]. Development [Akb09, ACSHvdH17, Koç10, RCP^{+}16, Ber15, Göz10, Şen09, Tur11, ÜA08]. DHA [Dun10]. Dhaka [RAB^{+}17]. Di-Calcium [SOA^{+}15]. Diameter [Şen13a]. diatom [Pol08b]. Diatomix [NB19]. Diatoms [NB19]. Dibang [KAGD16]. Dibrugarh [DADL17]. Dicentrarchus [Alp13, Jal14, Kor08]. Diet [CCL18, Fıl09, LVA15, SOO15, Ozt07]. Dietary [Köp12, Law14, Lin14, YYBP07]. Diets [ATAA17, Law13, MKS17, Naz12a, RBT^{+}17, STFO15, UEA18, VBR19, Yt12]. difference [Bos07]. Different [Alt13, AAP18, Ayd14b, Bra19, Bra20, Bul12, CFFM16, Dum12, Gün10, KTT16, KAA^{+}17, Khi11, Law13, MK14, ÖÖK07, Özt12, PP16, Şah09, VHKB20, Bil07, Bos07, Göz10, HB07, İÖÇK10, Kay13, Ors10b, Özp10, Tok08, Tür09, Yıl10]. Differentiation [Deg16]. Digesting [MP17]. dimension [RAB^{+}17]. Dimensions [Bau13, Jaw12a]. Dimensions-Fish [Bau13]. Dimorphism [Ayd11c]. dinoflagellates [Pol08b]. Dinophyceae [JK16]. Dip [Ayd14a, Gön13]. Diplodus [Ayd14a, Ayd14b]. Diptera [Tad09]. Direction [BDdF19]. Dirma [TM16]. dıs [Şen14]. discus [Çel10]. Disease [Ber21c]. Diseased [Jal14, Kor08, SSG13]. Diseases [TYB^{+}13]. Disi [ŞHG^{+}08]. disinfectants [Str14]. disorder [Zoe11]. Distichodus [AF11]. Distinction [Şen14]. Distribution [Aya10, Beh14, Boz09b, Boz10a, ÇKAE10, ÇKB^{+}13, Duy08, ER15, Gön10, Özv13b, Özv13a, Özv14a, RCV^{+}17, Pol08b, Şah10, AKA11a]. distributions [Bal07, BÇÖ07]. District [AKD^{+}17, Cha14, GDAD16, TWT19]. Diversity [GD13, HSAE16, NBBR17, TWT19]. Dizileri [Kes12a]. DNA [NS17, dMSD^{+}16]. DNA-extraction [NS17]. Do [Har15]. Doctor [Bha16]. Dogu [Atı12, ÖL08, ŞHG^{+}08]. Dolphin [Ber21b, LWP^{+}16]. Dominant [LL08]. Doñana [ANM19]. Donax [Çol14]. dönüsümü [Tur11]. Dosage [BL17]. Dose [SK13]. Down [Bra19]. Dried [STFO15]. Drinkable [Tar10]. Drives [FGF20]. Drum [Baş09]. Dry [NLM16]. Dry-Cured [NLM16]. due [KMK09]. Dulakhojiya [GDAD16]. Dumanlanmis [KÖÇ12]. Dumeril [Bil07]. dumerili [Öks12]. Dummeril [Bau13]. Dummy [GMG^{+}20]. Dünyada [Yeş13]. Duration [SK13]. During [ASTH21, Koç10, Kub09, Şen13a, Tek15, UAH^{+}07, ZPW^{+}17, Gül07]. Durumlari [Kuş13]. Durusu [Saç14]. Dusky [Öks10]. Dwelling [Sim14]. Dynamical [Man08]. Dynamics [ANM19, Meh13, Pol08b].

East [Beh14, Dun10]. Eastern

[Baç12, Duy08, ÖL08, Yab09a, Rad13, Şah10, TMGY16]. ebonyicum [AO19]. Ecologically [CPE^{+}15]. Economic [Alh12, BMK17, Bra19, Law13, Uzm08, AKA11a, Bal07, BÇÖ07, Doğ09b, KAGD16, Yel11]. economical [Doğ10]. Economics [NPR16]. Economy [Law14, vGdV16]. ecosystem [Bat07, Lev11]. ecosystems [AAÖ11]. Ecotoxicological [Ade12]. Ecotoxicology [Har15]. ed [Dağ08]. Edible [Çağ11, Dur09]. Edirne [Erd12]. Editor [Ano20b, Ano20a, Ber17a]. Editorial [Ber21a, Fro21a, Fro21b, Jon20, Özk10, Özk14a]. edulis [SOA^{+}15]. edwards [Özc13]. Eel [Abu13, Deg16, Ero09]. Effect [ASTH21, Ade09, Akb13, AO19, Ayd11a, Ayd11b, Ayd11c, Dor12, ELKESA19, Gar20, HB07, JK16, KAA^{+}17, Kel10, KDCS20, KST16, MM21, MKS17, MYMZ15, MKK15, Özt10b, SBW^{+}17, SSG13, SAA^{+}17, SGH^{+}16, SRB^{+}15, Tür09, UAH^{+}07, AAÖ11, Kut10, Zoe11, Alp13]. Effectiveness [AKV^{+}17, KKP16, KG16]. Effects [AEU18, Alt13, AAP18, Ayd08b, Bah07, BA16, Bul12, BAU^{+}12, CCL18, Doğ09a, FY21, Kar11, Kay14b, Köp12, Law14, Man15, MAC19, MS07, Naz12a, OF18, Opi14, Ozt07, PRI16, PCV^{+}16, Şen13a, Sim14, SCEA08, Tar10, UACÖ13, Usl11, Usl12a, Usl12b, Utk08, Yue18, Mol08, Tok08, Yıl10, Tur11]. Efficacy [KG16, RPCC15]. Efficiencies [Bra19]. efficiency [Ors10b, Ors10a]. Ege [Taş11]. Egg [ACSHvdH17, HBP^{+}18, KAA^{+}17, MYMZ15, STFO15, Şen09]. Eggs [DTT^{+}16, Kub09, Ayd11b]. Egirdir [Bal07, BÇÖ07, DB08]. Ehrenberg [Kut10]. Ekmekci [BSİU12, Baş11b, BSİU12]. Ekonomik [AKA11a, Yeş11]. Eksi [ÖL08]. Elazig [İS13, PÇD^{+}15a, PÇD^{+}15b, Sal09, İpe12, İS12, Sal11, İS12]. Elazig-Türkiye [Sal11]. Elazıg-Türkiye [İpe12, İS12]. Eledone [Şen11]. Element [Öks10, Usl12a]. Elemental [RCV^{+}17]. elements [KMK09]. Elevation [FGF20]. Elimination [DVSS12]. Embryo [AASASG^{+}15]. Embryos [FY21, Kle13]. Emphasis [KAGD16, RCP^{+}16]. encrasicolus [İÖÇK10, Kay08]. Endpoints [Har15]. Enfestasyonu [Ürk13b]. enfestasyonunun [Açı13]. Engin [BSİU12, Baş11b]. Engraulis [İÖÇK10, Kay08]. engycephalus [AF11]. Enhancer [NB19]. Enriched [RPCC15]. enterprises [Uzm08]. Enumeration [AKD^{+}17, GDAD16]. environment [Man15]. Environmental [ATEA^{+}16, BMK17, Deg16, KDCS20, MG16, SMG^{+}18, Yab09a]. Enzim [Çap11, Yiğ11a]. Enzyme [AEU18, UACÖ13]. EPA [Dun10]. Epilithic [ÇKAE10]. Epilitik [ÇKAE10]. Epipelic [Soy10]. Eritrosit [Çap11]. Erythrocytes [PMOBF18]. Escaped [Ayd08a]. Eschcholtz [Ber09]. Eschscholtz [Yük11]. Escox [Ber21a]. Esomus [DAM18]. Esox [Yac09]. Especially [PRI16]. Essipov [Aks11b, Aks11a]. Estimate [Bra20]. Estimation [Aks11b, DAS^{+}16]. Estradiol [Yue18]. Estrogen [Yue18]. Estuarine [FY21, MG16]. Ethiopia [JZ15, Tek15, TM16, TWT19]. Etkisi [Alp13, Çap11, Oğu11, Yiğ11b]. Etkileri [BAU^{+}12, Tür09, Tur11]. etkisi [Usl12a]. eugenol [Özp10]. Eurasian [Lin14].

European [Cha20, Deg16, ML16, Şen14]. Euryhaline [VGP^{+}17]. euxini [KÖÇ12]. euxinus [Özd11]. Evacuation [Baş11a]. Evaluate [OAE^{+}18]. Evaluation [HDS^{+}16, KKP16, KG16, Lem12, MYMZ15, Oeh08, PMOBF18, SOO15, Bos07, İÖÇK10, Mar11]. ex [CPE^{+}15]. Example [Ayd10b]. Exercise [VHKB20]. Existence [CFFM16, YÖ10]. Exotic [Ugu07]. expenditures [Mar11]. Experimental [KLPK16, OAA15]. Exploitation [SIAF16]. Exportation [Sağ11]. Exposed [GAN17, Has15]. Expression [Yue18, Kle13]. Extends [EN17]. Extension [FA14]. Extracellular [BMBS17]. Extraction [Ayd10b, Khi11, NS17].

Fikosiyanin [Oğu11]. Face [HDS^{+}16]. Facilitates [Pep21]. Facility [BAU^{+}12]. Factor [DADL17, SAMN16, TWT19]. Factors [ACSHvdH17, Ber15, Doğ09a, KDCS20, Küç08, MG16, RGB^{+}16, Tek15, DED^{+}10, Mar09]. Fair [Lem12]. fallax [CFFM16]. Family [Gan15, Jaw12b, DED^{+}10]. familyası [Tay11]. Farkli [Özd11, Tür09]. Farklılıgın [Kes12a]. Farm [LL08, LL10]. Farming [VGP^{+}17, Yab09b, Man15, Tos13]. Farms [Yıl11, Cha14]. Farnesoate [P19]. Fat [Har13, PÇD^{+}15a, PÇD^{+}15b, Kab09, Yıl10]. Fat-Soluble [PÇD^{+}15b, PÇD^{+}15a]. Fatty [Çağ11, Gül07, Har13, Kay08, MKS17, Öks10, Öks12, PÇD^{+}15a, PÇD^{+}15b, Tur07, Dun10, NCA13, Yıl10]. Fauna [Boz09a, ÇKB^{+}13, Erd12, GD13, İSU14, Özt13, ÖTGG07, Özu11, Saç14, Sal09, Tad09, Taş12, TTÇY08, Dal08, EG12, Taş10b, Top09]. Faunasi [İSU14]. Faunasinin [ÇKB^{+}13]. Fayda [AC08]. Fazla [Alp13]. Fe [Bah07]. Features [BSİU12, Bay12, Gür11, Tar10, Özp10, Taş10a]. Fed [GTO15, Law13, RBT^{+}17, STFO15, SOA^{+}15, UEA18, Akb10]. Feed [Baş11a, Köp12, KST16, P19, UAH^{+}07, Yil14, Yt09, Dem08a, HB07]. Feeding [Abu13, Akb13, Ayd11a, BKB15, BM17, DAM18, DNMS13, DO15, DADL17, MYMZ15, Şen11, Kar11, Şen13b, YYBP07]. Feeds [LRMO19, Der10, HB07, Meg14]. Female [AAP^{+}17, P19, Rad13, Yue18]. Fenton [Bah07]. Fermented [NLM16, VBR19]. Fertilization [KAA^{+}17]. Fertilized [AOO16]. Fichawa [Tek15]. Fillets [ASTH21, LL07]. films [Dur09, Dur10]. Finfish [Yab09b]. Finfish-Seaweed [Yab09b]. Finger [ELKESA19]. Fingerling [SOO15, Öz10]. Fingerlings [AK17]. Finike [Ayd14b]. Firat [Çob11, Kar10]. First [DDAD17, Fer14, FA14, Gök14, Özv14b, Tan12]. Fish [AKD^{+}17, AB15, Ayd10b, BSİU12, Bau13, BMK17, Ber21c, Bha16, ÇKB^{+}13, Dağ08, DADL17, EN17, ELKESA19, FY21, GDAD16, Gön10, Gül08, Har13, Har15, IEN^{+}19, İSU14, Jaw12a, KTT16, KÖÇ12, KY17, KANO21, KST16, KAGD16, LL08, MK14, MCAA19, MAC19, NLM16, ÖTGG07, PRI16, PP16, RAB^{+}17, RCR^{+}16, RCP^{+}16, Saç14, Sak15, SSG13, Sim14, Tek15, TM16, TTÇY08, ÜBİG07, Ürk13a, YKŞ15, Bal07, BÇÖ07, Bos08, Çel10, Dal08, Dem08a, Der10, Dun10, EG12, Kab09, Kar11, Kay13, eGK14, LL10, Man15, Mol08, Mol11a, RGO13, Tok08, Ugu07, Ünl09, Yel11, YV07, YYBP07, Erg12]. fishburger [VdFHLAA15]. fishburger-Type [VdFHLAA15]. Fisheries [Alh12, Bay12, Cha20, Dem09, Fro21a, Fro21c, Gün14, JZ15, KAGD16, MM21, Sed14, Sin15, Tiw19, Aky11a, Bat07, KHK^{+}13, Lev11, Meg14]. FisheriesSciences [Ber17a]. FisheriesSciences.com [Ano20b, Ano20a, Jon20]. Fisherman [AKD^{+}17, JZ15, MM21]. fishermen [Doğ09b]. Fishery [BA16, G16, KKP16, Zla17, AS10, BG10, Doğ10]. Fishes [AKDD17, Beh14, Doğ09a, ER15, PC16, RCP^{+}16, SPC^{+}17, TWT19, AKA11a, Ber15, Bos09, Dog08, NCA13, ÖEAÖ10]. Fishing [BJSN19, Bay12, ÇKB^{+}13, KLPK16, ML16, SIAF16, Aky11a, Ayd08b, Dar10, Doğ09b, Erd10, Yel11]. fishmeal [Meg14]. Fitness [SBW^{+}17]. Five [Gül08, Saç14, eGK14]. Flagged [ML16]. Flathead [Has14]. flavescens [RGB^{+}16, VBR19]. Flavobacterium [YÖ10]. Flesh [Öks12]. flesus [Ayd11d, Şah10, BKB15]. Flora [AOO16, Kub09, Özo10, Kar07]. Flotow [Usl12b]. Flounder [BKB15, SBW^{+}17, Ayd11d, Şah10]. Flow [Gün10, Gün11, Gün12]. Fluctuations [BDdF19]. fluviatalis [Lin14]. Flying [DAM18]. foliaceus [Açı13, Özt13]. Food [Abu13, Akb13, DFDRMGT17, Gar20, PRI16, RCP^{+}16, Erk08, Kar11]. Foods [Naz12a, Kar11, Kil11]. foods-packagings [Kil11]. Foodstuffs [Doğ09a]. Formaldehyde [Bul12, Yon14]. Formulated [UEA18]. Forsskål [Öks10]. Forum [Tiw19]. Four [Bra20]. Fractions [MYMZ15]. fragrans [SAA^{+}17]. Frailes [FA14]. Frameworks [Ber21c]. Frequencies [PMOBF18]. Frequency [Ayd11a, AOO16, Pol08a]. Fresh [SMG^{+}18, dMSD^{+}16]. Freshwater [AO19, BMKS20, Gül08, RCR^{+}16, SKB^{+}18, DED^{+}10, Ugu07]. Freshwaters [Lha14]. frisii [Tar08]. frontinalis [ACSHvdH17]. Frozen [ASTH21, ELKESA19, Oeh08, Tok08]. Fry [DTT^{+}16, KAA^{+}17, OKA16, OO13, Opi14, STFO15, Kar11, Kel10]. Full [Har13]. Full-Fat [Har13]. Functional [Gar20, Öza09]. Fungicide [Kay14a]. Future [Akb09, RCP^{+}16, SMJO15].

G [Ünv12]. Galeoides [SIAF16]. galileus [Bra20]. gallina [Çol14]. galloprovincialis [Öza09]. Gambella [TWT19]. Gambusia [Gün09, Gün10]. Gametes [PCV^{+}16, SPC^{+}17]. Ganges [Ber21b]. gangetica [Ber21b]. Gar [KÖÇ12]. Garden [Çağ11]. Garfish [PiYB09]. gariepinus [ATAA17, Ade09, AOO16, AEU18, AAP18, GTO15, GAN17, KAA^{+}17, KG16, Law13, Law14, LVA15, OKA16, OAE^{+}18, SAA^{+}17, SOO15, Tur11, UEA18]. Garra [Bha16]. garvieae [ATG^{+}14]. Gasoline [HDS^{+}16]. Gastric [Baş11a]. Gastropod [Özv13b]. Gastropoda [Tan12]. Gears [Bay12]. Geban [İpe12, İS12]. Gediz [Tad09]. Gelatine [Khi11]. gelisimi [Tur11]. Gelistirilmesi [Ayd14a]. Gen [Kes12a]. Gender [Evr09]. gene [Kle13]. Genera [LL08]. general [Dem08a]. Genes [Yue18, ZPW^{+}17]. Genetic [Ber17b, DAS^{+}16, Kes12b]. Genetics [CFFM16]. Genetik [Kes12a]. Genus [FA14]. Geographical [RCV^{+}17]. Geometry [BDdF19]. Gerado [TM16]. Ghana [Bra19, Bra20, MCAA19]. gibelio [Bos07, Tar08, YYBP07]. Gıcı [Soy10]. Gida [Sür11]. Gida-Ambalaj [Sür11]. Gill [AF11, SIAF16, Ors10b, Ors10a]. Gillnet [Bra19, Bra20]. Gillnets [Bra20, Özt12, Ayd08b, Dar10]. Gills [AK17, Has15, Mat14]. Gilo [TWT19]. gilt [AKÖ09, Aka10]. Gilthead [Yt12]. Girard [Gün09]. Girars [Gün10]. Giresun [Özt12]. Girgir [ŞHG^{+}08]. GIS [ATEA^{+}16]. Gizzard [GMG^{+}20]. Glass [Gök14]. Global [Tiw19, Yeş13, vGdV16, Tos13]. Glucan [SBW^{+}17]. Glucose [Bul12]. Gluten [Yt12]. gobiidae [Tay11]. Gobius [Fıl09]. Goblet [BM17]. Goby [Fıl09]. Gökkusagi [Çap11, Yiğ11b, YÖ10]. Göksu [BAU^{+}12, Rad13, BAU^{+}12]. Golden [ZPW^{+}17]. Golden-Line [ZPW^{+}17]. goldfish [Özt10a]. Golleri’nde [YYBP07]. Gölü [Baş11b, BSİU12, Çob11, YPY07, Yük11, Açı13, Bay12]. Gölü’ndek [Öİ08]. Gölü’Ndeki [ÇKB^{+}13, Ünv12]. Gonad [RMKM17, Tur11]. gonadal [Tur11]. Gönen [İSU14, İSU14]. Göre [ÇKB^{+}13]. Görgüsan [İpe12, İS12, İS12]. görünüsünden [Şen14]. Gösteren [Pol12].

Government [BA16]. Gracilariopsis [MP17]. gracilis [SGH^{+}16]. grahami [ZPW^{+}17]. Grass [Köp12, Özc13]. Gravading [MS07]. Greece [PKMB17]. Green [Akt14, Özc09]. Green-lipped [Akt14]. Grey [Bah07, OO13, Gül07]. Grimm [Duy12]. Group [FE15]. Growing [Gök09]. Growth [Akb10, Akb13, AO19, Ayd11a, Ayd11c, Ayd12, BSİU12, BMKS20, CCL18, Çol14, Deg16, DADL17, Erg10, FY21, Fıl09, GTO15, Köp12, KST16, Law13, Law14, MKS17, MYMZ15, OKA16, Opi14, P19, RGB^{+}16, SBW^{+}17, Sib08, SAA^{+}17, STFO15, SIAF16, TMGY16, UEA18, Usl12b, UAH^{+}07, Yt09, Çob09, DED^{+}10, HB07, Kar11, Kut10, Meg14, ÖL08, Ozt07, Tür09, Öİ08]. Grunow [Kut10]. grypus [DAS14, PÇD^{+}15a, PÇD^{+}15b]. Gujarat [VGP^{+}17]. Gulf [Has14, Gök14, Özv14b, RGO13]. Gümüs [Taş11]. Gümüshane [Ayd10a]. Günther [KÖÇ12]. Guppies [Kar11]. guppy [YV07]. Gussevia [Mat14]. Gut [RBT^{+}17]. Güvenilir [YPY07]. Güvenligi [Yeş13].

H [Bah07, DAS14]. H. [Özc13]. Haar [LWP^{+}16]. Habit [Abu13]. Habitat [OAA15, Tar10]. Habits [DAM18, DNMS13, DO15]. Haematococcus [Usl12b]. Haematogical [SSG13]. Haematological [GAN17, MAC19]. Haematology [Ade09, SOA^{+}15]. Haemocytes [Gel09]. Hairtail [KLPK16]. Hamilton [DAM18, DNMS13, MKS17, SAMN16]. Handling [Ade09]. Hanging [Bra19, Ors10b, Ors10a]. Hantzchia [Kut10]. Hapa [OKA16]. Harbour [AB15]. Hard [MG16]. Hare [Özv14a, Tan12]. Harmful [JK16]. Harvest [KBD17]. Harvesting [ANM19, Per10, Tek15]. harveyi [Çan11]. Haryana [Ani14]. Hatay [Boz10b, Dem09, Gül08]. hatcheries [YÖ10].

Hatchery [Kub09, STFO15, Ayd11d]. hatchery-reared [Ayd11d]. Hatching [Ayd11b, FY21, KAA^{+}17]. Havuz [YYBP07]. Hazar [Çob11, Dev11]. head [AKÖ09, Aka10]. Heading [BJSN19]. Headline [Ayd08a]. Heads [Khi11]. Health [Bha16, Erg12, Gar20, VdV16, Mol08]. heat [Kay10]. Heavy [Esr14, GC16, Gün08a, IEN^{+}19, Kay14b, RSSG16]. Heckel [Ayd12, Çob11, Erg10, PÇD^{+}15a, PÇD^{+}15b]. Hedef [ŞHG^{+}08]. Helix [Çağ11]. Hematocrit [SCEA08]. Hematological [DAS14, GTO15, SGH^{+}16, Yon14]. hepsetus [ELKESA19]. Herbicide [PMOBF18]. heteroclada [MP17]. heteropods [Özv13a]. Heterotis [KANO21, KANO21]. hexylresorcinol [Sel14]. Hg [AB15, PP16]. hibritlerinin [Ünv12]. Hidrobiyolojide [Gün11]. High [PRI16, Pep21, Erk14]. High-Throughput [Pep21]. Hill [AKDD17]. Hilsa [HBP^{+}18]. Hippocampus [Kuş13]. Hirfanli [Gür11, eGK14]. Hirudo [Sağ11]. Histological [Abt14, Dın10]. Histology [BM17]. Histopathological [RSA^{+}21]. Histopathology [AK17]. Historical [Akb09, Has15]. Hoc [Gün14]. hoferi [Özt10a]. Homa [Kut10, Özc09]. Hong [Tiw19]. Hooded [Meh13]. Hook [Ayd14b]. Hormonal [Deg16]. Hormone [Baş09]. Hormones [KG16]. Horse [Baç12, Erd10]. Horticulture [Cha20]. hospital [Mar11]. Hot [ÜBİG07, Bil07]. Hot-Smoked [ÜBİG07]. Hour [Gün09]. HPLC [Sel14]. HSP [Kay10]. HSP70 [EN17]. human [Mol08]. Huso [Akb10]. hybrids [Ünv12]. Hydrachnidia [Fer14]. Hydrobiology [Gün12]. hydrocarbon [Ade12]. hydroelectric [AAÖ11]. hydrophila [DVSS12, Str14]. Hydrozoa [AİGS11]. Hypnea [Sid13]. Hypophthalmichthys [HDS^{+}16]. hypophthalmus [DTT^{+}16].

icerigine [Usl12a]. Ichthyophonus [Özt10a]. Ichthyophthirius [Özt10b]. Icin [YPY07]. idella [Köp12]. Identification [NLM16, Özo10, Ter08]. Identify [NS17]. Identity [NPR16]. ilish [HBPH19]. ilisha [HBPH19, HBP^{+}18]. iliski [Bau13]. Illegal [ML16]. Illex [Öza09]. Image [RMKM17]. Immersion [SK13]. Immune [Alt13, RSA^{+}21]. Immunostilumants [Sak15]. Immunostimulants [Erg12]. Impact [Har13, SKB^{+}18, SMG^{+}18, Bat07, Göz10, Lev11]. Impacts [VHKB20, Yab09a]. Implanted [GMG^{+}20]. Implementation [BA16]. Implication [RCP^{+}16]. Importance [IEN^{+}19, Kay10, Kil11]. Imposex [SRB^{+}15]. Improve [Ber17b, Gar20]. Improvement [BM17]. Improving [Ayd14a]. inceburun [Gön13]. Incidence [BJSN19]. Incorporated [CCL18]. Increased [Yt09]. Increments [Ayd10b]. Incubation [Kub09, Şen09]. Index [RSSG16]. India [NBBR17, AKDD17, AKD^{+}17, Ani14, Beh14, Cha14, DAM18, DDAD17, DADL17, GDAD16, KAGD16, PRI16, SKB^{+}18, SAMN16, SKK16]. Indian [AK17, AB15, AA13, Ani14, DAM18, MKS17, Sin15]. indicators [OF18]. Indices [Lem12, SGH^{+}16, UEA18]. indicus [Has14]. Indigenous [KANO21]. Indirect [Bra20]. Indonesia [MUK17]. Indoor [OKA16, STFO15]. Induced [Has15, PMOBF18]. induces [Kle13]. Induction [Kay14b]. industry [Erk08]. Inexpensive [Pep21]. Infected [ATG^{+}14]. Infecting [AK17]. infection [Özt10a, Yon10]. infections [Mar11]. Infectious [Gan15]. Infestation [Özt13, Ürk13a]. Influence [BDdF19, PP16, Yt09]. Influenced [ACSHvdH17]. Influencing [ACSHvdH17, RGB^{+}16]. information [Kle13]. infralittoral [Kar07]. Inhabiting [Ero09, ÖL08, Ugu07, YYBP07]. Inhibitory [Naz12a]. Initial [Gök09]. insect [Top09].

Insecta [DB08, Top09]. Inspired [CPE^{+}15]. Instantaneous [YOH^{+}16]. Integrated [BDB20, Yab09b]. Interaction [Utk08]. Intestine [BM17]. Intra [MG16]. Intra-community [MG16]. invader [Dev11]. Invalidated [Par13]. Invasive [FA14]. Investigation [ATG^{+}14, Bil07, DADL17, Kay14a, Rad13, ŞHG^{+}08, ÜA08, Yon14, Yük12]. Investigations [AAP^{+}17, Dın10, Özt13]. Invigorate [SKK16]. Involvement [Deg16]. Iodine [Ayd11b]. ionizing [Erk14]. Ipomea [SOO15]. Iran [Has14, KHK^{+}13]. Iraq [HSAE16]. ischemic [Mar09]. Isıkların [Tür09]. Isıklı [Yac09]. Island [MUK17, Sid13, CGZ21]. Isolated [DTT^{+}16, NLM16, Kut10]. Isolates [LL08]. Isolation [BS18, dMSD^{+}16]. Isparta [BAU^{+}12, ÇKAE10, BAU^{+}12]. Isporuz [Ayd14a]. Issue [Özk14b, Özk14c, vGdV16]. issues [Evr09]. Istanbul [GD13, Saç14, Doğ10, Erk07, Mol11a]. Istranca [Özu11]. Italian [BA16, CFFM16]. IUU [ML16]. Izmir [Taş10b, Aky11b, Dın10, Gür12, Kut10, Şen11, Tan12]. Iznik [Doğ09b, Tar08].

Java [MUK17]. jelly [Dev11]. Jenys [ÜA08]. Journal [Ano20b, Ano20a, Jon20]. Juvenile [Ayd11a, BKB15, BDB20, Köp12, Law14, Lin14, SAA^{+}17, Yt09, ZPW^{+}17, Akb10]. Juveniles [Mat14, Yt12, DED^{+}10].

Kafadanbacaklıların [Şen13b]. Kahraman [Yon10]. Kahraman-maras [Yon10]. Kahverengi [Pol12]. Kainji [OAO^{+}15]. Kalitesine [Alp13]. Kalamar [Şen14]. Kalite [KÖÇ12, Yeş13]. kanagurta [AB15]. karabalık’ta [Tur11]. Karaburun [Yab09a]. Karacaören [Öİ08, Öİ08]. Karachi [AB15]. Karadenizin [Özd11]. Karadeniz [Gön13, PiYB09, ŞHG^{+}08, Dev11]. Karadeniz’de [Aks11a]. Karadeniz’deki [Atı12]. Karakaya [Ayd12, Ero09]. Karatas [Gel09, BDAK10, ÖL08]. Karatas/Adana [Gel09]. Karbosulfanin [Çap11]. karides [BG10]. Karsilastirilmasi [Çob11, Özd11]. Karyomorphology [Ani14]. Karyotype [GUA14]. Kashmir [NBBR17]. Kaynagi [BAU^{+}12]. Keban [Dar10, AST^{+}12, Ayd12, Çob11, PÇD^{+}15a, PÇD^{+}15b, Yük11, Yük12]. kefali [Açı13]. Kelaart [Özv13b]. Kemiksi [YPY07]. Kerevit [Yük11]. Kilis [Dağ08, Dal08, Dağ08]. Killed [YOH^{+}16]. Kimyasal [Duy12]. Kınacık [Dal08]. Kırklareli [Erd12, Uzm08]. Kırmızı [Pol12, Bau13]. Kirsehir [Gür11]. Kiyilarinda [AKA11a]. Kıyısalında [Pol12]. Kıyısındaki [Kes12a]. Klinoptilolit’in [Yiğ11b]. klunzingeri [ÖL08]. Knowledge [Tad09]. known [Bos09]. Köftelerinin [KÖÇ12]. Kompozisyonlarinin [Özd11]. Kompozisyonu [ÇKB^{+}13, Duy12]. kompozisyonundaki [Usl12a]. Kompozisyonunun [ŞHG^{+}08]. Kong [Tiw19, Tiw19]. Korea [ML16]. Körfezi’nde [Yeş11]. Körfezi [Pol12, Taş11, Tay11]. Körfezi’nde [Aka11b]. Kosullarında [Duy12, Şen13b]. Kosullarinin [Oğu11]. kottelati [Baş11b, BSİU12]. Køycegiz [Gül09]. Köyündeki [YÖ10]. kullanilan [BG10, Bay12]. Kullanılarak [Kes12a]. Kullanım [Aka13]. Kullanimi [Yiğ11a, Erg12, Gün11]. Kültür [Şen13b]. Kültürü [Çan11]. Kuluckahanede [Ayd11d]. Kuluckahanelerinde [YÖ10]. Kum [Çol14]. Kürk [Sal11]. Kuzeydogu [BG10, Pol12]. KY808492 [AKV^{+}17].

L. [Açı13, Alp13, Ayd09, Ayd14b, BDAK10, Çol14, Gür11, İÖÇK10, Kay08, Kay14a, Opi14, Öza09, Özt13, Öİ08, Pol08a, PiYB09, RBT^{+}17, Sağ11, Şen09, Şen13a, Sib08, Sim14, Tur07, Tür09, Ünv12, Yac09, Yt12].

L.1758 [DÖD^{+}13]. labeo [MKS17, OOB16]. Laboratory [GAN17, OO13]. labrax [Akb12, Alp13, Baş11a, Jal14, Kor08]. Lacépède [CFFM16]. Lactic [NLM16]. Lactococcus [ATG^{+}14]. Lagoon [Beh14, BDdF19, Gel09, Gül09, Kut10, Özc09, Özo10, SYG^{+}08]. lagoons [Çev08]. Lake [BSİU12, Bay12, Ber09, Boz09a, Boz09b, Boz10a, Bra19, Bra20, DB08, DAS14, Erg10, EG12, Ero09, FGF20, Gür11, NBBR17, OAO^{+}15, PÇD^{+}15a, PÇD^{+}15b, Saç14, Sal14, Sib08, Soy10, Taş07b, Ünv12, Yac09, YPY07, Doğ09b, AİGS11, Ayd12, Bal07, BÇÖ07, ÇKB^{+}13, Dar10, FGF20, JZ15, Özt13, ÖTGG07, Öİ08, RCR^{+}16, TNS17, Tar08, Taş10a, Yük12]. lakes [Top09, YYBP07]. Lakhimpur [GDAD16]. lalandi [PLS^{+}17]. Lamarck [Şen14]. Lambda [Gün09]. Lamp [KKP16]. Lanao [RCR^{+}16]. Lankester [AİGS11]. Large [FGF20]. Larvae [Ayd10b, Baş11a, ER15, FY21, Naz12a, UEA18, RGO13, ÜA08]. Larval [Abt14, Çel10]. larvaların [Tay11]. Larviculture [PLS^{+}17]. Laundry [MMdA^{+}15]. leachii [Özv14a, Tan12]. Lead [BM17, SCEA08, SRB^{+}15, Ber15]. Leaf [ATAA17, LVA15]. Lecithin [MYMZ15]. LED [KKP16]. Leeches [Sağ11]. Legislations [Yeş13]. leidyi [Dev11]. leidyi’nin [Dev11]. Leiognathus [ÖL08]. Length [Bau13, Ber09, Ero09, Gül09, Gül08, Jaw12a, eGK14, Lha14, Özc09, PiYB09, SAMN16, TWT19, YKŞ15, Zla17, ÖEAÖ10, Pol08a, Şah10]. length-frequency [Pol08a]. Length/Weight [Zla17]. Length/Width [Özc09]. Length/Width-Weight [Özc09]. Lengths [Tar08]. leptodactylus [Ber09, Yük11]. Lessepsian [Özv14a, Tan12]. Lesser [SIAF16]. Lethal [Gün09, HDS^{+}16]. Lettuce [ATAA17]. Leucaena [LVA15]. Leukocyte [Şah09]. Level [LWP^{+}16, Erk07, Kel10]. Levels [ATAA17, AB15, Esr14, GTO15, Kay14a, Köp12, SCEA08, Utk08, Kel10]. Levrek [Alp13]. life [Duy12]. Light [DB08, Har15, Öks08, Usl12b]. Light-Trapping [DB08]. lights [Tür09]. Lin [Öİ08]. Line [AAP^{+}17, KLPK16, ZPW^{+}17]. Lineage [CFFM16, PCV^{+}16]. linking [Kle13]. Linnaeus [Akt14, Aya11b, Ayd11b, Has14, Küç08, OO13, Özt12, SCEA08, YPY07, Yt09]. Linne [Akb13]. Linneaus [AKÖ09, Aka10, Ayd11a]. Lipid [Bah07, Gar20, Köp12, Dun10, Gül07, Usl12a]. Lipidiozis [DÖD^{+}13]. Lipit [Pol12]. lipolytica [BMBS17]. Lipoprotein [MYMZ15]. lipped [Akt14]. List [Beh14]. Listeria [Göz10]. Litopenaeus [MKK15, RPCC15]. Live [Baş11a, Dın10, Naz12a, Kar11]. Liver [Har13, Utk08, Ozt07]. Living [Ayd12, Erg10]. Liza [Gül07]. Lizardfish [Man08]. Locality [Özc13]. Lohit [AKD^{+}17]. Loligo [Şen14]. long [Dun10]. Longfin [FY21]. longiceps [AA13]. Longline [Ayd14b]. Looming [vGdVW16]. low [Kel10]. Lower [TTÇY08, OOB16]. Luciobarbus [Baş11b, BSİU12, BMKS20]. lucioperca [Öİ08, YPY07]. lucius [Yac09]. Lupara [Per10]. luridus [Öks10]. luscus [Ayd11d, Şah10]. luth [Par13]. Lutjanidae [Jaw12b]. Lutjanus [AAP^{+}17, Jaw12b, MMdA^{+}15]. Luzon [CGZ21]. Lymphocystis [Ber21c]. Lysine [Lin14]. Lysozyme [Bul12].

Miktari [AKA11a]. Miktarina [Oğu11]. Miktarlari [Özd11]. Mackerel [AB15, Baç12, Khi11, YOH^{+}16, Erd10]. Macro [OAA15]. macrobenthic [Taş10b]. Macrognathus [Abu13]. macropomum [DAS^{+}16, dMSD^{+}16]. macrostigma [Bau13, Bil07]. macrostomus [YG08b]. Maheshkhali [Akt14]. Mahseer [RMKM17]. Maize [ATAA17, SOO15]. Major [AK17, Ani14, MKS17, Zoe11].

Makroalg [Pol12]. Malatya [Ero09]. Male [AAP^{+}17, Yue18]. Manage [Sim14]. Management [BM17, CFFM16, JZ15, Mar11, Özg09]. Manasbal [NBBR17]. Manaus [Lem12]. Manipulating [Nik15]. Manipulator [P19]. Mannans [SBW^{+}17]. Manure [AOO16]. maras [Yon10]. Marbled [Öks10]. March [Tiw19]. marinades [Özp10]. marinated [İÖÇK10]. Marine [ATEA^{+}16, CPE^{+}15, DFDRMGT17, Jal14, Kor08, Yab09a, Çan11, Man15, NCA13]. Market [Cha20]. marketing [RAB^{+}17]. markets [Mol11a]. Marmara [Çol14, Yıl11]. Marmara’da [Çol14]. Martin [Sid13]. mascarensis [BMKS20]. Masculinization [SK13]. Mass [Jaw12b]. Mastacembelus [Ero09]. Matlab [Zla17]. Matlab-Based [Zla17]. Mature [Aya11a, Aya11b, Gel09, Özo12]. Mavi [Kes12a]. maxima [Ayd11a, Ayd11b, Ayd11c]. Meagre [Par13]. Meal [GTO15, Har13, LVA15, LRMO19, SOO15, SOA^{+}15, VBR19, Yt12]. measures [Cha14]. Meat [Akb13, Aya11a, Ber09, KÖÇ12, Öks12, Özo12]. Meats [Öza09]. Mechanical [BMK17]. Medicinal [Sağ11]. medicinalis [Sağ11]. mediterranea [Özv13a]. Mediterranean [Aya10, BDAK10, DÖD^{+}13, Duy08, Pol08b, Rad13, TMGY16, BG10, Kes12b, Öks12, ÖL08, Özc09, Özv13b, Özv13a, Özv14a]. mediterraneus [Atı12, Baç12, Erd10]. Mekong [DTT^{+}16]. Melek [Ürk13b]. Melet [TTÇY08, Esr14]. Melibe [Özv13b]. menace [Cha14]. Menemen [Tad09]. Menemen-İzmir [Tad09]. Mercury [AB15]. Meristik [Aka11b]. Merlangius [Özd11, Özt12]. merlangus [Özd11, Özt12]. Mersin [Man08, YÖ10, Aya10, Aya11a, Aya11b]. Meshes [Bra20]. Mesocosm [PLS^{+}17]. Mesohaline [MG16]. Mesopotamichthys [MYMZ15]. Mesozooplankton [İşi09]. metabisulphide [Erk07]. Metabolic [AEU18]. Metabolism [PC16, Dog08]. Metal [Akt14, Esr14, Kay14b]. Metallerin [KMK09]. Metals [GC16, Gün08a, IEN^{+}19, MK14, RSSG16]. Method [Ayd09, Pol08a, Sel14]. Methods [Alt13, Bra20, KAA^{+}17, NS17, Mar11, Tok08, Ünl09]. Methyl [P19]. Methyltestosterone [SK13]. Metolachlor [GAN17]. Metrik [Aka11b]. Metrik-Meristik [Aka11b]. Mevsimsel [ÇKAE10]. Mevzuatlar [Yeş13]. Mezgit [Özd11]. Microalgae [Usl12a]. Microbial [UEA18]. Microbiome [RBT^{+}17]. Microcapsules [Naz12b]. Micromesistius [Oeh08]. Micronuclei [PMOBF18]. microphotography [Çel10]. Midst [vGdV16]. midwater [Erd10, ÖEAÖ10]. Midwest [FGF20]. Midyesi [Çol14]. mikroalglerde [Usl12a]. Milne [Özc13]. Mineral [Çağ11, Öks12]. Mirror [Ayd09, Yt09]. Mitigation [SKB^{+}18, Cha14]. Mixture [GAN17]. Mnemiopsis [Dev11]. Model [Ayd08a, Deg16, SMJO15, vGdVW16]. Modeling [ANM19]. modifiable [Mar09]. Modifications [Dem08b]. Molds [DTT^{+}16]. Molecular [BL17, CFFM16]. molitrix [HDS^{+}16]. Mollusca [Tan12]. Monitoring [Yab09a]. monocytogenes [Göz10]. monofilament [AKÖ09, Aka10, Ayd08b, Ors10b, Ors10a]. Monthly [Yel11]. Morphologic [Şen14]. Morphological [OAO^{+}15, Ter08, Yük12]. Morphometric [Aya10, Çob09, DAS^{+}16, Rad13, SPC^{+}17]. Morphometry [DNMS13]. Mortality [HDS^{+}16, Öİ08, SIAF16, TMGY16]. mortem [YOH^{+}16]. moschata [Şen11]. Mosquitofish [Gün10]. mossulensis [YG08a]. most [YPY07]. Motion [BJSN19]. mountain [Top09]. Mounting [Nik15]. Movement [KDCS20]. mrigala [BL17, SSG13]. Mud [NPR16]. Mugil [Bah07, OO13]. Mullet [Ayd14a, Bah07, OO13, Aks11b, Gül07]. Mullus [Aks11b, Aks11a, Ayd14a, Özd11]. multidentata [Gök14]. multifilament [Ayd08b]. multifiliis [Özt10b]. Multiple [Alp13, Gün14]. Munzur [Bau13, Bau13]. Muscat [AA13, Jaw12b]. musciformis [Sid13]. Muscle [Har13, Öks08, RSSG16, YOH^{+}16, Kab09]. Muscles [KY17]. Musky [Şen11]. Mussel [Akt14, Öza09]. Mustafa [MKK15]. Mutton [MMdA^{+}15]. mykiss [ATG^{+}14, Çap11, Did13, Göz10, Gün08a, Gün08b, Kel10, Kub09, MS07, Öz10, Özp10, Ozt07, Ser09, TYB^{+}13, UAH^{+}07, Yiğ11b, YÖ10, Yıl11, Yon10, Yon14]. Myofibrillar [KY17]. Myristica [SAA^{+}17]. Mytilus [Öza09, SOA^{+}15]. Myxozoans [AK17].

N [Özd11]. Na-Alginate [Naz12b]. Nadu [Cha14]. Nam [DTT^{+}16]. nanocomposite [Dur10]. Nanoparticles [SGH^{+}16]. Nanotechnology [Kil11]. Nanoteknolojik [Sür11]. Nardo [Özc09]. Native [OAA15]. Natural [ASTH21, Dur10, Meg14, Öz10]. Naturally [ATG^{+}14]. ’nde [Taş11]. ’ndeki [Açı13, Tay11, YPY07]. Nearby [TWT19]. Nearshore [SIAF16]. Neck [YOH^{+}16]. Neck-Breaking [YOH^{+}16]. Need [Har15]. Nehri [Çob11]. nehri’ndeki [Bau13]. Nemacheilidae [GUA14]. Neotropical [BDdF19, SPC^{+}17]. Net [Ayd08a, Ayd14a, BMK17, SIAF16]. Nets [Tos07, AKÖ09, Aka10, AS10, BG10, Ors10b, Ors10a, Özy09]. Neural [Dın10]. Ng [DDAD17]. Nigella [Alt13]. Niger [GC16, IEN^{+}19, Fıl09]. Nigeria [Ade12, AF11, GC16, IEN^{+}19, OF18, OAO^{+}15]. nigrodigitatus [AF11]. Nikolskii [Abt14]. Nil [Tür09]. Nile [LRMO19, Opi14, OAE^{+}18, RSA^{+}21, RBT^{+}17, SOA^{+}15, SBL^{+}18, Tür09, Yue18]. niloticus [OAE^{+}18, Bra20, KANO21, LRMO19, Opi14, OAE^{+}18, PCV^{+}16, RSA^{+}21, RBT^{+}17, SOA^{+}15, Tür09, Yue18]. ’nin [Açı13, Alp13, Ayd11d, ÇKAE10, Çob11, Duy12, YPY07, YYBP07]. nisin [VdFHLAA15]. Nitrifier [RPCC15]. Non [Mar09]. Non-modifiable [Mar09]. Nona [HBPH19]. nonylphenol [Kle13]. NOR [YG08b]. North [Dun10, İşi09, Kar10, TMGY16, vGdVW16]. North-East [Dun10]. North-Eastern [TMGY16]. North-Sea [vGdVW16]. Northeast [AKDD17, BG10]. Northeastern [BDAK10, DDAD17, Pol08b, Aya10, NCA13]. northern [RGO13, RGB^{+}16]. Northwest [CGZ21, Has14]. notata [Aka11b]. Note [Ano20b, Ano20a, Ber17a, Ber21b, Fro21a, Fro21b, Jon20]. nov [RPCC15]. Novel [LZ21, RCP^{+}16]. Nuclear [PMOBF18]. Nucleolar [YG08a]. Nucleotide [ÖÖK07]. Number [KBD17]. Numerical [KLPK16]. Nutmeg [SAA^{+}17]. Nutrient [Usl12a, Usl12b]. Nutritional [LVA15, LRMO19, RSA^{+}21]. Nutritive [MS07].

O [Bah07]. Oahe [FGF20]. obesus [RCV^{+}17]. Occurance [AOO16]. Occurrance [AF11]. Ocean [MKK15, RCV^{+}17]. Oceans [VdV16]. ocellaris [KTT16]. ocellatus [Mat14]. Octopodidae [Şen11]. Octopus [Öza09, Şen11, Öza09]. Odisha [PRI16]. off [ÖL08]. officinalis [Aky11b, BDAK10, Öza09, Özo12, Şen13a]. officinlis [Şen09]. Ogun [AF11]. Ohi [İS13]. Oil [AA13, Alt13, Has15, Kay08, Ade12, Der10, Mol08]. Oken [FA14]. Olan [BAU^{+}12, Çan11, Dev11]. Olanakları [Aka13]. Oligochaeta [Taş12]. Oligocheta [Taş10b]. Ölüm [Öİ08]. Oman [AA13, Alh12, AAP^{+}17, Jaw12a, Jaw12b]. Omani [Meh13]. Omega [Erk13, Dun10]. Omega-3 [Erk13, Dun10]. Ömrünün [Duy12]. Onchorhyncus [Gün08a, Gün08b].

Oncorhynchus [ATG^{+}14, Çap11, Did13, Göz10, Kel10, Kub09, MS07, Özp10, Ozt07, Ser09, TYB^{+}13, UAH^{+}07, Yiğ11b, YÖ10, Yıl11, Yon10, Yon14]. Oncorynchus [Öz10]. Önemi [Dik11, Sür11]. önerileri [MU10]. Onitsha [IEN^{+}19]. Önünde [ÖL08]. Oocyte [Koç10]. Ootrope [Ayd08a]. Operational [KKP16]. operatives [Doğ10]. Opistobranchia [Özv13b]. Optimal [ATAA17]. Oranlari [Öİ08]. Ordu [Esr14, Özt12, Taş07a, Taş10a, TTÇY08]. Oreochromis [Bra20, CCL18, Has15, Opi14, OAE^{+}18, PCV^{+}16, RSA^{+}21, RBT^{+}17, SOA^{+}15, Tür09, Yue18]. Organic [Khi11]. organisms [KMK09, Kay10]. Organizer [YG08a]. organizma [Dev11]. Organizmalar [KMK09]. Organogenesis [Abt14]. Organs [IEN^{+}19]. Origin [Ber17b]. Original [LWP^{+}16]. Originated [AF11]. Ornamental [AKDD17, Beh14, DADL17, RCP^{+}16, Kay13]. Oromia [Tek15]. Oromia/Ethiopia [Tek15]. Orontes [DO15]. Orta [Aks11a]. Ortam [Alp13]. Osmaniye [Boz09a, Boz09b, Boz10a]. Osteichthyes [YG08b]. Osteoglossum [Lem12]. Ostracoda [Özu11]. Other [PMOBF18, STFO15]. Otolit [Atı12, Bau13]. Otolith [Ayd10b, Baç12, Ero09, Jaw12a, Jaw12b, Bau13]. Otoliths [Ayd09]. Our [VdV16]. Outdoor [OKA16]. Ovaprim [KG16]. Overview [Alh12, Ber21a]. Ovulin [KG16]. owner [Doğ09b]. Oxidant [Kay14a]. Oxidation [Bah07, Gül07]. Oxidative [Kay14b]. oxidized [Ozt07]. Oxygen [Akb12, Kel10]. Oxynoemacheilus [GUA14]. Özelliklerinin [Çob11]. özellikleri [Ayd11d, Bay12, BG10, Aka11b, Atı12, Baş11b, BSİU12, Taş11]. Özelliklerinin [Çol14]. Oziothelphusa [P19].

Pâté [ÜBİG07]. packagings [Kil11]. Pain [YV07]. painkiller [YV07]. pair [Erd10]. pairly [ÖEAÖ10]. Palaemonetes [Özc13].

Palaemonidae [Özc13]. paleatus [ÜA08]. Pallas [Akb12, Ayd11d, Baş11a, Şah10]. Pandemic [MM21]. Pangasianodon [DTT^{+}16]. pannosa [Sid13]. parameter [Kel10]. Parameters [ATG^{+}14, Ayd12, Baç12, Bra20, DAS^{+}16, DAS14, GTO15, KÖÇ12, Kay14b, Man08, MAC19, MYMZ15, Oeh08, Öks08, SSG13, SCEA08, Yon14, Aks11b]. Parametreleri [Atı12, Aks11a]. Parametrelerinin [KÖÇ12]. Parasite [SMG^{+}18]. Parasites [Kay13]. Parasitic [RSA^{+}21]. Parasitism [Mat14]. Parasitological [Lem12]. Parent [Pep21]. Parent-Based [Pep21]. Parluciosoma [SAMN16]. Part [TTÇY08]. Pasiphaea [Gök14]. Passive [BDB20]. Past [NBBR17]. Patella [Aya10]. pathogen [Çan11]. Pathogenic [BS18, DVSS12]. Patojeni [Çan11]. Pattern [AA13]. patterns [RGO13]. Pay [JZ15]. PCR [LL07, NS17, ZPW^{+}17]. PCR-based [NS17]. PDGS120915 [JK16]. Peel [SOO15]. Peeled [BS18]. Peels [Law14]. pelagic [ÖEAÖ10]. pelagicus [Aya11b]. Penaeidae [Gan15]. Peninsula [Yab09a]. Pens [BMK17]. People [KAGD16, vGdV16]. peppered [ÜA08]. Perca [Lin14, Par13, RGB^{+}16, VBR19]. Perch [Lin14, RGB^{+}16, VBR19]. Perciformes [Par13, PMOBF18]. Performance [Akb13, AAP18, Ayd11a, CCL18, FY21, GTO15, KKP16, KST16, Law13, MKS17, MYMZ15, OKA16, Opi14, SAA^{+}17, STFO15, VHKB20, Akb10]. Performances [UAH^{+}07]. Period [Gök09]. Perna [Akt14]. Persian [Has14, RGO13]. Perspective [Ber21c]. Perspectives [RCP^{+}16]. Pertek [Dar10]. Peruvian [Mat14]. Peters [Kar11]. Phage [BS18]. Phenotype [YG08b]. Philippines [RCR^{+}16, SRB^{+}15]. Phosphate [SOA^{+}15]. Phosphorus [Law13, vGdVW16]. Photoperiods [Yt09]. phycocyanin [Usl11]. Phylogenetic [CFFM16]. Physical [Tar10]. physico [Taş10a]. physico-chemical [Taş10a]. Physiological [AAP18]. Phytoplankton [CGZ21, Çev08]. Pigment [KTT16]. Pigments [Usl11]. Pike [Yac09]. pikeperch [Öİ08, YPY07]. pilchardus [ÖÖK07]. Pink [Gök14]. pinosylvin [Göz10]. Pisces [SIAF16, YG08a]. pisi [Ayd11d]. Pistia [ATAA17]. PIT [BDB20]. Place [Kuş13]. Planktonic [Taş07b]. Planning [BA16]. Plant [Doğ09a]. Plantation [vGdVW16]. plants [AAÖ11]. Platanista [Ber21b]. platensis [Oğu11, Usl11, Yıl10]. Platichthys [Ayd11d, BKB15, SBW^{+}17]. Platichtys [Şah10]. Platycephalus [Has14]. Plecoglossus [Ayd10b]. pluvialis [Usl12b]. Poecilia [Kar11, YV07]. Policies [ML16]. Pollution [ATEA^{+}16, OF18]. Polyculture [Gök09]. polycyclic [Ade12]. Polynemidae [SIAF16]. polyunsaturated [Dun10]. Ponds [AOO16, MUK17]. ponticus [Aks11b, Aks11a, Özd11]. Populasyon [Atı12, Kuş13, Yük11, Aks11a]. Populasyonları [Kes12a]. Populasyonlarinin [Çob11, Çol14]. Populasyonunun [Baş11b, BSİU12, Öİ08]. Population [ANM19, Baç12, BSİU12, Ber09, Kuş13, Man08, Meh13, Yac09, Aks11b, Öİ08, Kar10]. Populations [CFFM16, Çol14, Kes12b, RGB^{+}16, Bal07, BÇÖ07]. porcus [Aka11b]. Porifera [CPE^{+}15]. Port [ML16]. Portunus [Aya11b]. Post [Gün14, YOH^{+}16]. Post-Hoc [Gün14]. Post-mortem [YOH^{+}16]. Potamon [AO19]. Potato [SOO15]. Potencial [MMdA^{+}15]. Potential [BL17, Naz12a, SBW^{+}17, Ayd10a]. Poultry [AOO16]. poutassou [Oeh08]. Powder [Dum12]. power [AAÖ11]. Pradesh [AKDD17, AKD^{+}17, DDAD17, KAGD16]. prashadi [Meh13]. Precedence [Par13]. Preference [Şen11, Şen13a]. Preferences [BKB15]. prejuvenile [Çel10].

Preliminary [Aky11b, Naz12b, Özu11, Rad13, ŞHG^{+}08, Şen11, ÖL08]. prepared [Özp10]. Preprocessing [RMKM17]. Presence [FA14]. Present [KHK^{+}13, NBBR17, Ram17, Tos13]. preservation [Erk14]. pressure [Erk14]. Prevalence [SMG^{+}18]. Previda [RBT^{+}17]. Principally [Bra20]. Probiotic [BL17, Did13, KDED11, MAC19, PLS^{+}17, SSG13]. Probiotics [DVSS12, RPCC15, Ram17]. Probiyotik [KDED11]. Problems [TNS17, MU10]. Procambarus [ANM19]. Procedure [PP16]. Process [Alp13, MS07]. processing [Erk14, MU10]. Produced [Dum12, Kay08]. Produces [MK14]. producing [Ade12]. Production [Ber17b, Law14, NPR16, RCP^{+}16, SKK16, SBL^{+}18, UAH^{+}07, Kab09, Meg14]. productivity [Dar10]. Products [MK14, NS17, G16, Uzm08, VdFHLAA15]. Profile [Gar20, MKS17, Sid13, NCA13, Tur07]. Profiles [Öks10]. Profilinin [Erk13]. Profitability [Yıl11]. program [ATEA^{+}16]. Projection [Akb09]. Proper [ML16]. Properties [GUA14, Öza09, Sib08, Yac09, Erg10, Gür12, Çol14]. Prospect [Ram17]. Protease [BMBS17, Naz12a]. Proteases [MMdA^{+}15]. Protection [Sağ11, Meg14]. Protein [Bah07, Dog08, Köp12, Pol12, Utk08, VBR19, Yt12, Dur09, Tok08]. Proteins [KY17, Kay10, Ünl09]. Province [Gül08, Ugu07, RGO13]. Proximate [Öks10, Sid13, Tur07]. Prussian [YYBP07]. Psetta [Ayd11a, Ayd11b, Ayd11c]. psychotherapy [Zoe11]. Pterois [FA14]. Pterophyllum [Ürk13a, Ürk13b]. public [Mar11]. Published [FA14]. Pugilina [SRB^{+}15]. Puntius [DNMS13, RCR^{+}16]. purification [Ünl09]. Purpose [AASASG^{+}15]. Purse [KKP16, ŞHG^{+}08].

Quail [STFO15]. Quality

[ASTH21, Alp13, Dum12, Erk08, KÖÇ12, MUK17, Oeh08, OAE^{+}18, Tar10, Yeş13, AST^{+}12, Mol11a, Tok08]. Quantification [LL07]. Quantitative [ZPW^{+}17, Mar11]. Quantity [Usl12b]. Quartering [BJSN19].

radiata [Koç10]. radiation [Erk14]. Raf [Duy12]. Ragged [Özv14a, Tan12]. Rainbow [Alt13, ATG^{+}14, BM17, BDB20, Bul12, Did13, Gün08a, Gün08b, Har13, KBD17, KDCS20, Kub09, MS07, Ser09, SKK16, TYB^{+}13, UAH^{+}07, VHKB20, Yıl11, Yon14, Göz10, HB07, Kel10, Özp10, Ozt07, YÖ10, Yon10]. Raised [AOO16]. Raja [Tur07]. ramada [Gül07]. Range [FA14, Top09]. Rapid [LL07, LL08, Sel14]. Rapture [Pep21]. Rastrelliger [AB15]. Rate [Akb12, Akb13, Ayd11a, Ayd11b, Baş11a, HDS^{+}16, DED^{+}10]. Rates [Dum12, Gün10, Yıl10]. Rathbun [Aya11a, Gel09, Gül09]. ratio [Ors10b, Ors10a]. Ratios [Bra19, Özp10]. Raw [AOO16, UAH^{+}07]. ray [Tur07]. Real [LL07, ZPW^{+}17]. Real-Time [ZPW^{+}17, LL07]. Reared [DÖD^{+}13, Opi14, Ayd11d]. Rearing [AKDD17, VHKB20]. Recently [Tos07]. Receptor [Yue18]. Recirculating [LZ21]. Record [DDAD17, Fer14, FA14, Gök14]. Records [ÖTGG07, Saç14]. Recreational [KBD17]. Red [ANM19, Ayd14a, Bau13, CCL18, MP17, Sid13, VGP^{+}17, Aks11b, ER15]. Red-Spotted [Bau13]. Redescription [AK17]. Redox [PC16]. Reef [ER15]. Reference [NPR16, RSA^{+}21, ZPW^{+}17]. refrigerated [Gül07]. refrigeration [Duy12]. Regan [Özv14b]. Region [Alh12, Aya10, Dar10, DÖD^{+}13, Gön10, HSAE16, TWT19, Yıl11, Akb09, Dağ08, Uzm08]. Regions [Şah09, SBW^{+}17, YG08a]. regium [Erg10]. regius [Par13]. Regulations [Özg09]. Rejimi [YYBP07]. Relation [MG16]. Relations [Bau13]. Relationship [Ero09, Jaw12a, Lha14, SAMN16, Tar08, TWT19, Zla17, ÖEAÖ10]. Relationships [Ber09, Gül09, Gül08, Özc09, PiYB09, RCV^{+}17, YKŞ15, eGK14]. Relative [Bra20]. relatives [Evr09]. Relevance [NPR16]. Reliable [ZPW^{+}17, YPY07]. remission [Zoe11]. rendalli [PMOBF18]. Renklerdeki [Tür09]. Repeatability [Ser09]. Replacement [ATAA17, Har13, SOO15]. Report [Özv14b, Tan12]. represent [Bos09]. Reproduction [Aky11b, Fıl09, Gür11, Tan12, Yac09, ÜA08]. Reproductive [AAP^{+}17, AAP18, RCR^{+}16, Ünv12]. Requirement [Lin14]. rerio [Kay14b, Koç10, SMJO15]. Research [BJSN19, SMJO15]. Researchers [Tiw19]. Reservoir [AF11, FGF20, eGK14, Tar10, Taş10b, AST^{+}12, KDCS20]. Reservoirs [Tek15]. Resident [Sim14]. Residual [LWP^{+}16]. residues [Sel14]. Resource [G16]. Resources [OF18, Aky11a]. Response [BJSN19, Küç08]. responses [KMK09]. Results [Aky11b, BA16]. retail [Mol11a]. Retention [BDB20, GMG^{+}20]. reticulata [Kar11, YV07]. reversal [Tur11, Par13]. Review [CGZ21, Gan15, KANO21, MCAA19, MAC19, OAA15, Yab09b, Ber15, Çan11]. Rhodeus [Lha14]. ring [Bos08, Bos09]. Ripe [Law14]. Risk [AB15, AASASG^{+}15, Bha16, Mar09]. Riskleri [AC08]. Risso [Gür12, Öks12, Taş11]. River [AKD^{+}17, Boz10b, DDAD17, DADL17, Esr14, IEN^{+}19, KAGD16, OF18, OOB16, Özc13, PRI16, Şah09, AAÖ11, Bau13, Ber21b, Dem09, OAE^{+}18, TTÇY08, TWT19]. Rivers [TM16]. rivulatus [Öks10]. Roach [Sib08]. rohita [MKS17]. Rohu [MKS17]. Ropes [Bra19]. Rotifera [Boz09b, Boz10a, Erd12, HSAE16]. Rotten [MP17]. Roundup [PMOBF18]. Routine [Akb12]. Routines [VHKB20]. rufa [Bha16]. rulina [Usl11]. Ruppell [AAP^{+}17, Öks10, OOB16, Jaw12a]. Rural [MM21]. ruthenus [Akb13]. Rutilus [Sib08, Tar08].

S [Erd10]. S. [Ünv12]. Sac [FY21]. Safety [Yeş13]. Saglıgı [Erg12]. Sahalarinda [Özd11]. Sakaraya [ÖTGG07]. Salalah [AAP^{+}17]. Saleem [MKK15]. salina [RPCC15]. Saline [AAP18]. Salinity [OO13, Yıl10]. Salmo [Akb12, Baş11a, Bau13, Bil07, Kar10, Sim14]. Salmon [FGF20]. Salmonella [BS18]. Salmonids [Pep21]. Salted [HBPH19, Göz10]. Salvelinus [ACSHvdH17]. Same [Bra19]. Sampled [AF11]. Samples [Zla17]. Samsun [PiYB09, Soy10, Taş07b, YPY07, Ugu07, YPY07, YYBP07]. Sand [ELKESA19, Doğ09b, Gür12]. Sander [Öİ08, YPY07]. Sanliurfa [DAS14]. Sapanca [ÖTGG07, AİGS11]. sapidus [Aya11a, Gel09, Gül09, Kes12a, Kes12b, Özo10, TMGY16]. Saponin [RBT^{+}17]. Sarda [Öks08]. Sardina [ÖÖK07]. Sardine [AA13, ÖÖK07]. Sardinella [AA13]. sargus [Ayd14b]. Sarıkum [SYG^{+}08]. Sarotherodon [Bra20]. Sartlarinda [Alp13]. sativa [Alt13]. sauce [İÖÇK10]. saurus [Man08]. Sausage [NLM16]. Save [Cha21]. scalare [Ürk13a, Ürk13b]. Scale [SBL^{+}18]. Scaly [KANO21]. schall [AF11]. Schizothorax [Abt14]. Schlegel [Ayd10b]. Schneider [Abu13]. Sciaenidae [Par13]. Science [Nik15]. Sciences [Gün14, Sed14]. Scomber [Khi11]. Scombridae [AB15]. scombrus [Khi11]. Scorpaena [Aka11b]. Scorpaenidae [FA14]. Scorpaeniformes [FA14]. Screening [CPE^{+}15, Did13]. Sea [Akb09, Akb12, Aky11a, Aky11b, AA13, Alp13, Ayd11a, Ayd11b, Ayd11c, Ayd14a, Baç12, Baş11a, Dun10, Fıl09, Gön10, Gür12, Jal14, Jaw12b, KY17, Kor08, Özc09, Özt12, Özv14a, PiYB09, Tan12, Taş07a, Yab09a, YKŞ15, Yt12, Zla17, Akb10, Aks11b, Aya10, Bat07, BG10, CGZ21, Çol14, Dev11, ER15, Erd10, FA14, İşi09, Jaw12a, Lev11, Meh13, Pol08a, ŞHG^{+}08, Şah10, Şen11, Tos07, Tur07, TMGY16, UAH^{+}07, vGdVW16]. Seabream [Ayd14b]. Seabreams [Ber17b]. Seafood [BA16, Dor12, Yeş13, Dur09, Dur10, Erk14, MU10]. Seafoods [Mol11b, Kil11]. Seahorses [Kuş13]. Searching [RCV^{+}17]. Seas [BJSN19]. Season [UAH^{+}07]. Seasonal [Çev08, ÇKAE10, Şah10, AST^{+}12, Ors10b, Pol08b, Soy10]. Seaweed [MP17, Yab09b, vGdV16, vGdVW16]. Seaweed-Plantation [vGdVW16]. Seaweeds [Sid13]. Seciciliginin [Ayd14a]. sector [MU10]. sectors [Dem08a]. Sediment [GC16, Yab09a]. Sediments [RSSG16]. seen [KMK09]. Seine [KKP16, ŞHG^{+}08]. Sektöründe [Sür11]. sektörünün [MU10]. Selected [GC16, IEN^{+}19, NLM16]. Selection [ZPW^{+}17]. Selectivity [AS10, Ayd14a, Ayd14b, Bra20, AKÖ09, Aka10]. Selı [Sal09]. Semen [dMSD^{+}16]. Semi [Taş12]. Semi-Aquatic [Taş12]. senegalensis [OOB16]. senex [P19, P19]. Senior [Par13]. Sensitivity [AF11, Str14]. Sensory [Dum12, ÜBİG07, İÖÇK10, Özp10]. Sepia [Aky11b, BDAK10, Meh13, Öza09, Özo12, Şen09, Şen13a]. Septicemia [Ürk13a]. Septisemi [Ürk13b]. Sequences [Kes12b]. Sera [SCEA08]. Serban [Özt13, Açı13]. Seriola [Öks12, PLS^{+}17]. Serratia [JK16]. Serum [Kay14a]. Sex [Deg16, Şah10, Tur11, Şen14]. Sexual [Ayd11c]. Sexually [Aya11a, Aya11b, Özo12]. Seyhan [Erg10, EG12, Özc13, Sib08]. Shabbout [DAS14]. Shabut [PÇD^{+}15a, PÇD^{+}15b]. Shad [GMG^{+}20, LZ21, Duy12]. shark [Dun10]. sharpeyi [MYMZ15]. Shatt [HSAE16]. Shelf [IEN^{+}19, Duy12]. Shell [MKK15, SOA^{+}15]. Shewanella [LL10]. Shi [Baş09]. shock [Kay10]. shops [RAB^{+}17]. shoreline [RGO13]. Shores [Özv13b, Özv13a]. Short [Gan15, GMG^{+}20]. Shrimp [BG10, BS18, Dem08b, Gök14, MKK15, NS17, Özc13, Ram17, Cha14, Meg14, Özy09, Sel14]. Shrimps [Gan15, Erk07]. Sialic [Kay14a]. Siang [DDAD17]. Sickness [BJSN19]. Siganus [Öks10]. Signal [LWP^{+}16]. Silifke [Man08]. Silifke-Mersin [Man08]. Silver [HDS^{+}16, Rad13]. similarity [Bos07]. Simulated [MKK15]. Single [CFFM16]. Sinkin [KAGD16]. Sinocyclocheilus [ZPW^{+}17]. Sinop [Gön10, SYG^{+}08, Gön13, Kar07, Kay08, Tur07]. Sinop-İnceburun [Gön10]. sinop-inceburun [Gön13]. Sırlanı [Çol14]. Sitometrinin [Gün11]. situ [CPE^{+}15]. Sivas [Ünv12]. Size [Ayd14a, Ayd14b, Ero09, Tar08]. Size-total [Ero09]. Skeletal [Ber17b, KY17, Ber15]. Skeleton [RSSG16]. Skin [AF11]. Small [KBD17, SBL^{+}18]. Smelt [ELKESA19, FY21, Doğ09b, Gür12]. Smoked [KÖÇ12, ÜBİG07, Bil07]. Smoking [MCAA19]. Snail [Çağ11]. Snapper [AAP^{+}17, MMdA^{+}15]. Socio [Alh12, Doğ09b, Doğ10, KAGD16]. Socio-Economic [Alh12, Doğ09b, KAGD16]. Socio-economical [Doğ10]. sodium [Erk07]. Software [PKMB17]. Soil [MUK17]. Solander [Ero09]. Sold [MK14, Erk07]. Soluble [PÇD^{+}15b, PÇD^{+}15a]. Some [Baç12, BDAK10, Dın10, DAS14, IEN^{+}19, KÖÇ12, KY17, Küç08, Öks08, OOB16, Öza09, RCR^{+}16, Sağ11, Şah09, Tar10, Taş10a, Yon14, Yük12, Aks11b, Kar11, KMK09, Kel10, ÖEAÖ10, Özt10b, Pol08a, SKB^{+}18, Top09, UACÖ13, Yon10, BSİU12]. Soovi [SIAF16]. sophore [DNMS13]. sorunlari [MU10]. sound [LWP^{+}16]. Source [BAU^{+}12, LWP^{+}16, VBR19, Yt12]. Sources [KTT16, Law13, Ugu07]. Sous [Mol11b]. Sous-Vide [Mol11b]. South [FGF20, HSAE16, TM16, AS10, CGZ21, LRMO19, ML16, RCV^{+}17, Şah10]. South-Eastern [Şah10]. Southeast [Özt12]. Southeastern [ŞHG^{+}08]. Southern [DB08, Taş07a, YKŞ15]. sowerbyi [AİGS11]. Soya [Har13]. Soybean [VBR19]. sp [CCL18, Has15, JK16, Ürk13a, Ürk13b, Yon10]. Sparus [AKÖ09, Aka10, Yt12]. Spatial [Bal07, BÇÖ07, ER15, RGO13]. Spawning [AA13, Baş09, KG16, LZ21, Şen09, Şen13a, Ser09]. Spearfishing [Özg09]. Special [RSA^{+}21]. Species [AK17, Aya10, Ayd08a, ÇKAE10, ÇKB^{+}13, CPE^{+}15, GD13, Gök09, Gül08, KY17, NS17, NPR16, OAA15, RCR^{+}16, Sağ11, TM16, VGP^{+}17, YKŞ15, AKA11a, Çan11, Dun10, Kay13, eGK14, Ter08, Ugu07, Yel11]. specifications [BG10]. Spectral [LWP^{+}16]. Speed [PRI16]. Spermatological [DAS14, Ayd11d]. spermatolojik [Ayd11d]. Spi [Usl11]. Spi-rulina [Usl11]. spilurus [DDAD17]. Spinefoot [Öks10]. spinosissimus [Özv14b]. Spiny [Abu13, Ero09, Özv14b]. Spirulina [Oğu11, Yıl10]. Sponges [CPE^{+}15]. Spotted [Bau13, YOH^{+}16]. spp [Çel10, Kuş13, MP17, YÖ10]. spp. [YÖ10]. sprat [Pol08a]. Sprattus [Pol08a]. Squalius [Açı13, Özt13]. Squid [Şen14]. St. [Sid13]. Stackhouse [CŞA10]. stage [Çel10]. Stages [ZPW^{+}17]. Standard [Bra19]. Starry [SBW^{+}17]. Starter [STFO15, VdFHLAA15]. State [Ade12, AF11, IEN^{+}19, Tos13]. States [FE15, ML16, SKB^{+}18]. Status [Kuş13, RSA^{+}21, Ram17, RCP^{+}16, Cha14, KHK^{+}13]. Steindachner [Atı12, Baç12, GUA14, ÖL08]. stellatus [SBW^{+}17]. Sterlet [Akb13]. Stock [Has14, Öİ08, Sim14]. Stocked [KBD17]. Stocking [AO19, FGF20, Opi14, Sim14]. Stok [Öİ08, Gön13]. Storage [ASTH21, ELKESA19, ÖÖK07, Gül07].

Stored [ÖÖK07, Bil07, Göz10]. Strains [NLM16]. Strategies [ANM19, SKB^{+}18]. stratiotes [ATAA17]. stratosphericus [RPCC15]. Stream [AKDD17, ÇKAE10, Dal08, GD13, İSU14, İS12, İS13, Sal09, Sim14, Kar10, Uzm08]. Streams [ACSHvdH17, Özu11, Tar10]. Stres [KMK09]. Stress [EN17, KMK09, Kay14b, Küç08, PC16]. Stresses [SMG^{+}18]. Stressing [Küç08]. String [Şen13a]. Striped [Çol14, DTT^{+}16]. Strips [Özo10]. stroke [Mar09]. Structure [Dem09, İS13, YPY07]. structures [Bos07, Bos08, Bos09, Pol08a]. Study [AASASG^{+}15, BKB15, BMKS20, Dem08b, Jaw12b, KY17, Koç10, Kub09, MUK17, NBBR17, Naz12b, OAA15, OKA16, Özu11, ŞHG^{+}08, Şen11, YV07, Yıl11, Ayd11d, Erd10, KANO21, ÖL08, Yon10, Gar20]. Subjective [Fro21c]. substances [UACÖ13]. Substituted [SOA^{+}15]. Substrates [MG16]. subtilis [AKV^{+}17]. Success [FGF20]. Successful [LZ21]. succession [Boz10b]. Sucul [KMK09]. Sudak [Öİ08, YPY07]. Suggested [vGdVW16]. suggestions [MU10]. Suitability [LVA15]. Suitable [NS17]. sulfate [SGH^{+}16]. Sultanate [AA13, AAP^{+}17]. Summer [İşi09]. Summit [Cha20, Tiw19]. Sundarban [NPR16]. super [RAB^{+}17]. super-shops [RAB^{+}17]. Supplemented [RBT^{+}17]. Supplements [PLS^{+}17]. Supply [HBPH19, HBP^{+}18]. Support [BA16]. Surimi [Dum12, Kab09]. Survey [BA16, RSA^{+}21]. Survival [AO19, ACSHvdH17, GMG^{+}20, KAA^{+}17, OKA16, Opi14, Yt09, DED^{+}10]. Susceptibilities [TYB^{+}13]. Sustainability [Sağ11]. Sustainable [KANO21, Meg14]. Suyuna [BAU^{+}12]. Swamp [ANM19]. Sweet [SOO15]. Swimmer [Aya11a, Aya11b]. Symphysodon [Çel10]. symptomatology [Zoe11]. Synodontis [AF11]. Synodus [Man08]. Synonym [Par13]. System [Alt13, Bah07, CPE^{+}15, KDCS20, LZ21, OKA16, SBL^{+}18, vGdVW16, RAB^{+}17]. Systems [BMKS20, PLS^{+}17, Ade12, Erk08].

T [Kle13]. T-box6 [Kle13]. T1693crge [Ünv12]. Tackle [vGdVW16]. Tag [GMG^{+}20]. Tagging [Pep21]. Tags [BDB20]. Tahtali [Taş10b]. taken [Zla17]. Tambaqui [DAS^{+}16, dMSD^{+}16]. Tamil [Cha14]. tanaica [Duy12]. tanks [Akb10]. taraklı [Dev11]. tatlı [Açı13]. Taurus [Top09]. Tayini [PiYB09, YPY07, Şen14]. Tebuconazole [Kay14a]. Technical [Bay12, Bra19, KKP16, Tos07, BG10]. Technique [Per10]. Technological [Öza09]. technologies [Erk14]. Technology [Mol11b, Kab09]. Tekirdag [Erd12]. teknik [Bay12, BG10]. Tekvando [Gün09]. Teleost [SPC^{+}17]. Teleostei [BMKS20, DDAD17, GUA14]. Temminck [Ayd10b]. temparatures [Göz10]. Tempe [TNS17]. Temperate [SBW^{+}17]. Temperature [FY21, Usl12b, Kut10]. temperatures [Bil07]. Temporal [ER15, RGO13]. Ten [vGdV16]. Tench [Gür11]. Tenualosa [HBPH19, HBP^{+}18]. Teorisinin [Aka13]. Term [GMG^{+}20]. Tespiti [ÇKB^{+}13]. Test [AASASG^{+}15, Özo10]. Testosterone [Yue18]. Tests [Gün14]. testudineus [RCR^{+}16]. Tethyidae [Özv13b]. Tetraodontidae [Özv14b]. Tezu [AKD^{+}17]. thallus [MP17]. Thawing [Alp13, Tok08]. Their [Doğ09a, Kuş13, NLM16, OAA15, Dor12, RSA^{+}21]. Theory [Sed14]. therapeutants [Özt10b]. there [BA16]. Thermal [PC16]. thermotolerance [Kay10]. Thevolta [Bra20]. thinlip [Gül07]. thornback [Tur07]. Thrace [Erd12, Taş12]. Threadfin [SIAF16]. Threatened [FY21]. threats [Cha14]. Three [Ani14, Bra20, IEN^{+}19, KKP16, Özt12, RGB^{+}16]. Throughput [Pep21]. Thunnus [DÖD^{+}13, Gar20, RCV^{+}17]. thynnus [DÖD^{+}13]. Ticari [ÇKB^{+}13]. Tilapia [CCL18, LL08, Opi14, PCV^{+}16, RSA^{+}21, RBT^{+}17, SOA^{+}15, SK13, Tür09, VGP^{+}17, Yue18, LL10, OAE^{+}18, SBL^{+}18, PMOBF18]. Tilapyası’nın [Tür09]. Time [Baş11a, Ser09, ZPW^{+}17, Ayd08b, LL07]. Tinca [Gür11]. Tirsi [Duy12]. Tissue [RCV^{+}17]. Tissues [Gün08a]. Tödürge [Ünv12]. Tolerance [OO13]. tombroides [RMKM17]. Top [Bra19]. Tor [RMKM17]. Total [Kay14a, LL07, Utk08, Dun10, Erk07, Ero09]. Toxic [OF18]. Toxicity [AASASG^{+}15, Gün08b, Has15]. Trabzon [AKA11a, AAÖ11]. Trace [Akt14, MK14, KMK09]. Traceability [BA16]. Trachurus [Atı12, Baç12, Erd10]. Tracking [RSSG16]. Traded [Lem12]. Traditional [Gök09]. Traits [DAS^{+}16]. trammel [AKÖ09, Aka10, AS10, Özy09]. transcriptional [Kle13]. Transgenic [RCP^{+}16]. Transgenik [AC08]. Transmitters [GMG^{+}20]. Transponders [BDB20]. Transportation [Ade09, Öz10]. Trap [BDdF19]. Trapping [DB08]. Trawl [Ayd08a, Ayd14a, Dem08b, Gön10, Tos07, Erd10, ÖEAÖ10, Yel11]. Treated [Kay14a, Yon14]. Treatment [Ayd11b]. Trends [Ram17]. Tributaries [GD13]. Tributary [KAGD16]. Trichoptera [DB08]. Trol [Ayd14a, Özd11]. Trolling [KLPK16]. Trolü [Yeş11]. trolüile [Gön13]. Trout [Akb12, ACSHvdH17, Alt13, ATG^{+}14, Baş11a, Bau13, BM17, BDB20, Bul12, BAU^{+}12, Did13, Gün08a, Gün08b, Har13, KBD17, KDCS20, Kub09, MS07, Ser09, Sim14, SKK16, TYB^{+}13, UAH^{+}07, VHKB20, Yıl11, Yon14, Göz10, HB07, Kar10, Öz10, Özp10, Ozt07, YÖ10, Yon10]. trouts [Kel10, UACÖ13]. True [Bra20, Bos09]. trunculus [Çol14]. trutta [Akb12, Baş11a, Bau13, Bil07, Ayd12, Kar10, Sim14]. Tüketilen [Erk13]. tumba [RCR^{+}16]. Tuna [DÖD^{+}13, Gar20, RCV^{+}17]. Tunas [Per10]. Tuncelli [Sal14]. Tunicates [Dın10]. Tür [AKA11a, ÇKB^{+}13]. Turan [Baş11b, BSİU12]. Turbot [Ayd11a, Ayd11b, Ayd11c]. Türkiye [Oğu11, Sal11]. Turkey [Ayd10a, Ber09, BDAK10, Boz09a, Boz09b, Boz10a, Boz10b, Çev08, ÇKAE10, Dağ08, Dal08, Dem08b, Doğ09b, Duy08, Ero09, GD13, Gel09, İS13, Özc13, Özo10, ÖTGG07, Rad13, Şah09, Sal09, Sal14, SYG^{+}08, Tad09, Tan12, Taş07a, Taş07b, Taş10a, Top09, TTÇY08, Usl11, Yab09a, Yac09, YPY07, AKÖ09, Aka10, AAÖ11, Dem08a, Dem09, DO15, DB08, DÖD^{+}13, Dın10, Fer14, Fıl09, Gök09, Gök14, Gül09, Gül08, Gür11, İS12, Jal14, Kay08, Kay13, Kes12b, eGK14, Kor08, Lha14, Man08, MK14, MU10, Mol11a, Özc09, Özu11, Özv13b, Özv13a, Özv14a, Özv14b, PÇD^{+}15a, PÇD^{+}15b, TYB^{+}13, YKŞ15, YÖ10, Yıl11]. Turkish [AKA11a, Alp13, Ayd11d, Ayd14a, BSİU12, Bau13, Bay12, BG10, BAU^{+}12, Çan11, ÇKAE10, ÇKB^{+}13, Çol14, Dağ08, Dev11, Duy12, Erd12, Erg12, İSU14, İS12, İşi09, KÖÇ12, KMK09, KDED11, Kuş13, MU10, ÖL08, Öİ08, PiYB09, ŞHG^{+}08, Şen13b, Şen14, Taş12, Tos07, Tur11, Tür09, Ünv12, Usl12a, Yeş13, YÖ10, YPY07, YYBP07]. Türkiye [İpe12, İS12, Yeş13, YPY07]. Türkiye’de [Erk13, MU10]. Türkiye’nin [Kes12a]. Türlerin [Çan11]. Türlerinin [Yeş11]. Türlere [ÇKB^{+}13]. türlerine [Tay11]. Türlerinin [Pol12]. Tursiops [LWP^{+}16]. Türü [Dağ08]. Türünün [Aka11b]. Tuzla [Çev08]. Two [AK17, Dağ08, NLM16, OF18, ÖTGG07, RCR^{+}16, Sid13, BMKS20, HB07]. Tylerius [Özv14b]. Type [Bah07, Gel09, VdFHLAA15, AAÖ11]. Types [Bra19, Şah09]. Typing [LL08].

Ultrasonography [RMKM17].

Ultrastructural [SPC^{+}17]. Ultrastructure [Koç10]. Uluabat [ÇKB^{+}13, ÇKB^{+}13]. Ulugöl [Taş10a]. umbla [Çob11]. Umbrina [Baş09]. Underneath [Ayd08a]. Understanding [Ber17b]. Unintended [BA16]. Union [ML16]. Units [CFFM16, Mar11]. Unregulated [ML16]. Unreported [ML16]. Unripe [Law14]. upper [Kar07, Abu13, DAM18, SAMN16]. Upregulation [EN17]. Urbanization [PRI16]. üreme [Ünv12]. Ürünleri [Dik11, Sür11]. ürünleri [MU10, KDED11]. Ürünlerinde [Aka13]. Ürünlerinin [Erk13, Yeş13]. Usage [Gün14, Sed14]. Use [AASASG^{+}15, BMK17, DVSS12, Der10, Erg12, Khi11, VdFHLAA15, Dur09, Mar11]. Used [Ayd14b, Bay12, Tos07, AKÖ09, Aka10, AS10]. Usefulness [Özo10]. Using [Bra20, Dum12, Gün12, KTT16, Kes12b, LL07, LWP^{+}16, Öz10, Ter08, Ünl09]. Usual [Bra19]. UTF [Gül07]. Utilization [Akb13, Baş11a, Köp12, Yt12]. Uygulamalar [Sür11]. Uygulamaları [KDED11]. Uygulanan [Alp13, Yeş13]. Üzerine [Çan11, Yiğ11b]. Üzerinde [KMK09]. Üzerindeki [BAU^{+}12]. üzerine [Tur11, ÖL08, ŞHG^{+}08]. Uzuncayir [Sal14].

V [Öza09]. Valenciennes [AA13, HDS^{+}16, OOB16]. Valorised [G16]. Value [LRMO19, MS07]. Values [RSA^{+}21]. Vandelli [Lem12]. vannamei [MKK15, RPCC15]. Variability [RGB^{+}16]. Variable [MUK17]. Variation [Kes12b, OAO^{+}15, Şah10, Soy10]. variations [Çev08, Ors10b]. Varied [GTO15]. Varieties [RCP^{+}16]. Various [TNS17, Ors10a]. Varlıgı [YÖ10]. Ve [ÇKB^{+}13, AKA11a, Aka11b, AC08, Atı12, Ayd14a, Bay12, BG10, ÇKAE10, Çob11, Çol14, Dev11, Duy12, Erg12, Gön13, İS12, İpe12, KMK09, Kuş13, MU10, Özd11, Öİ08, PiYB09, Pol12, Sür11, Tay11, Tur11, Ürk13b, Yeş13]. Venus [Çol14]. verbana [Sağ11]. verdigi [Dev11]. Vertical [AST^{+}12, Boz09b, Boz10a]. Vibrio [Çan11, Did13, LL08]. Vibrios [Jal14, Kor08]. Vicinity [Yab09a]. Vide [Mol11b]. Viet [DTT^{+}16]. View [Şen14, Dem08a]. village [YÖ10]. Villupuram [Cha14]. Viransehir [Aya10]. viridis [Akt14, Özv13b]. Viruses [Gan15]. Vitamin [Çağ11]. Vitamins [PÇD^{+}15a, PÇD^{+}15b]. Vitro [Özt10b]. Volema [SRB^{+}15]. Volta [Bra19]. Vona [Taş07a]. vulgaris [Öza09, Şen14].

W [Kel10, Öz10]. Walbaum

[ATG^{+}14, Did13, Göz10, Gün08a, Kub09, Ozt07, Par13, TYB^{+}13, YÖ10]. Warrants [PC16]. Wasted [Tek15]. Wastes [G16]. Water [ATAA17, AOO16, BAU^{+}12, DTT^{+}16, Fro21a, GC16, OF18, OAE^{+}18, SMG^{+}18, Tar10, Yab09a, AST^{+}12, MUK17]. Waterborne [SGH^{+}16]. Watermite [Fer14]. Waters [İşi09, Meh13, SIAF16, RGO13]. Wavelet [LWP^{+}16]. Way [YOH^{+}16]. Ways [SKK16]. Wedge [Çol14]. Weight [Ber09, DAS^{+}16, Gül09, Gül08, eGK14, Lha14, Özc09, PCV^{+}16, PiYB09, SAMN16, TWT19, YKŞ15, Zla17, ÖEAÖ10, Şah10]. West [Çol14]. Western [BMKS20, DÖD^{+}13]. Wet [UAH^{+}07]. wetland [GDAD16]. Wetlands [TWT19]. which [Ayd14b, Bos09]. whistle [LWP^{+}16]. White [Ayd14b, MKK15]. Whiting [Oeh08, Özt12, Akb10]. Width-Weight [Özc09]. Wild [Öks12]. Will [VdV16, Cha21]. Willingness [JZ15]. Winckworth [Meh13]. Wind [BDdF19]. Winter [CGZ21, UAH^{+}07]. Within [Doğ09a, CFFM16, Ugu07]. without [VdFHLAA15]. Wollo [TM16]. World [Har15].

X [CCL18]. X-xylanase [CCL18].

Xenobiotic [EN17]. Xia [MP17]. xylanase [CCL18].

Yag [Erk13, Pol12]. Yakalan [Dağ08]. Yanıtlar [KMK09]. Yapilan [Çan11]. Yapinin [YPY07]. Yarrowia [BMBS17]. Yarttıgı [KMK09]. Yas [PiYB09, YPY07]. Yas-boy [PiYB09]. Yasayan [YYBP07, ÖL08, PiYB09]. Yavrularinin [Yiğ11b]. Yellow [Rad13, RGB^{+}16, VBR19]. Yemlerinde [Yiğ11a]. Yene [Uzm08]. Yengec [Kes12a]. Yeri [Kuş13]. Yetistiricik [Kuş13]. Yetistiriciliginde [Dik11]. Yetistiriciliginde [KDED11]. yetistirilen [Ayd11d]. Yield [Akb13, Ber09, Öks12, Özo12]. YlTun15 [BMBS17]. Yogunluk [AKA11a]. Yolk [FY21]. Yolk-Sac [FY21]. Yöresinde [Dağ08]. Young [Tiw19].

zararlar [Dev11]. Zargana [KÖÇ12, PiYB09]. zarudnyi [Abt14]. Zebrafish [AASASG^{+}15, Kay14b, Koç10, SMJO15, Kle13]. zeolite [Öz10]. Zeway [JZ15]. Zhang [MP17]. Zinc [Gün08a, Gün08b, SGH^{+}16]. Zona [Koç10]. zone [Kar07]. Zooplankton [Boz10b, GD13, İS13, Sal14, İS12]. Zooplanktonu [İS12, İpe12, Sal11].

References

Al-Anbouri:2013:SPI

[AA13] Ibrahim Said Al-Anbouri.

Spawning pattern of Indian oil sardine, Sardinella longiceps Valenciennes, 1847 of Oman Sea, Muscat, Sultanate of Oman. Journal of FisheriesSciences.com, 7 (1):??, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/

fisheries-aqua/spawning-
pattern-of-indian-oil-
sardine-sardinella-longiceps-
valenciennes-1847-of-oman-
sea-muscat-sultanate-of-
oman.php?aid=544.

Aksungur:2011:EAE

[AAÖ11] Muharrem Aksungur, Orhan Ak, and Atilla Özdemir. The effect on aquatic ecosystems of river type hydroelectric power plants: the case of Trabzon–Turkey. Journal of FisheriesSciences.com, 5 (1):??, ???? 2011. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
effect-on-aquatic-ecosystems-
of-river-type-hydroelectric-
power-plants-the-case-
of-trabzonturkey.php?aid=
803.

Almamari:2017:RIM

[AAP^{+}17] Dawood Almamari, Laith A. A., Rumeaida Mat Piah, Abdulaziz Al-Marzouqi, Mikhail Chesalin, and Saeed Rabee. Reproductive investigations of male and female blue line snapper, Lutjanus coeruleolineatus (Ruppell, 1838) from Salalah Coast, Sultanate of Oman. Journal of FisheriesSciences.com, 11 (1):??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/reproductive-

investigations-of-male-
and-female-blue-line-snapperlutjanus-
coeruleolineatus-ruppell-
1838-from-salalah-coastsultanate.
php?aid=17518.

Amachree:2018:EDP

[AAP18] Dokuboba Amachree, Akinrotimi O. A., and Dabor Po. Effects of different physiological saline concentrations on the reproductive performance of Clarias gariepinus (Burchell, 1822). Journal of FisheriesSciences.com, 12 (4):??, ???? 2018. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/effects-
of-different-physiological-
saline-concentrations-
on-the-reproductive-performance-
of-clarias-gariepinus-
burchell-1822.php?aid=
23668.

Arellano-Aguilar:2015:UZE

[AASASG^{+}15] Omar Arellano-Aguilar, Soledad Solis-Angeles, Luis Serrano-García, Eduardo Morales-Sierra, Alejandra Mendez-Serrano, and Regina Montero-Montoya. Use of the zebrafish embryo toxicity test for risk assessment purpose: Case study. Journal of FisheriesSciences.com, 9 (4):??, ???? 2015. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/use-
of-the-zebrafish-embryo-

                                        toxicity-test-for-risk-
                                        assessment-purpose-case-
                                        study.php?aid=8192.

Ahmed:2015:MHL

[AB15] Quratulan Ahmed and Levent Bat. Mercury (Hg) levels in Indian mackerel Rastrelliger kanagurta (Scombridae) from Karachi fish harbour and its risk assessment. Journal of FisheriesSciences.com, 9(3):??, ???? 2015. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/mercury-
                                        hg-levels-in-indian-mackerel-
                                        rastrelliger-kanagurtascombridae-
                                        from-karachi-fish-harbour-
                                        and-its-risk-assessment.
                                        php?aid=5822.

Abtahi:2014:LOS

[Abt14]  Behrooz Abtahi.  Larval

organogenesis of Schizothorax zarudnyi (Nikolskii, 1897) (Cyprinidae): Histological aspects. Journal of FisheriesSciences.com, 8 (2):??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/larval-
                                        organogenesis-of-schizothorax-
                                        zarudnyi-nikolskii-1897-
                                        cyprinidae-histological-
                                        aspects.php?aid=261.

Abujam:2013:FFH

[Abu13] Santoshkumar Singh Abujam. Food and feeding habit of spiny eel Macrognathus aral (Bloch and Schneider) from Upper Assam. Journal of FisheriesSciences.com, 7(4):??, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/food-and-
feeding-habit-of-spiny-
eel-macrognathus-aral-
bloch-and-schneider-from-
upper-assam.php?aid=425.

Akhan:2008:TBF

[AC08] Süleyman Akhan and Mehmet Ali Canyurt. Transgenik baliklar: Fayda ve riskleri. Journal of FisheriesSciences.com, 2(3):??, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/transgenik-
baliklar-fayda-ve-riskleri.
php?aid=18431.

Acikel:2013:SBG

[Açı13] Melike Açıkel. Serban baraj gölü (Afyonkarahisar)’ndeki tatlı su kefali (Squalius cephalus, L. 1758)’nin Argulus foliaceus (Crustacea, Branchiura) enfestasyonunun araştırılması. Journal of FisheriesSciences.com, 7 (??):??, 2013. CODEN JFOIAQ. ISSN 1307-234X.

Alberto:2017:FIB

[ACSHvdH17] Ashley Alberto, Simon C.

Courtenay, Andre St-Hilaire, and Michael R. van den Heuvel. Factors influencing brook trout (Salvelinus frontinalis) egg survival and development in streams influenced by agriculture. Journal of FisheriesSciences.com, 11(2):??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/factors-
influencing-brook-trout-
salvelinus-frontinalis-
egg-survival-and-development-
in-streams-influenced-
by-agriculture.php?aid=
18377.

Adeyemo:2009:EHT

[Ade09] Olanike K. Adeyemo. Effect of handling and transportation on haematology of African catfish (Clarias gariepinus). Journal of FisheriesSciences.com, 3 (4):??, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/effect-
of-handling-and-transportation-
on-haematology-of-african-
catfish-clarias-gariepinus.
php?aid=1158.

Adeyemo:2012:EAP

[Ade12] Olanike K. Adeyemo. Ecotoxicological assessment for polycyclic aromatic hydrocarbon in aquatic systems of oil producing communities in Delta State, Nigeria. Journal of FisheriesSciences.com, 6(1):??, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/ecotoxicological-
assessment-for-polycyclic-
aromatic-hydrocarbon-in-
aquatic-systems-of-oil-
producing-communities-
in-delta-state-nigeria.
php?aid=681.

Akinrotimi:2018:EAM

[AEU18] O. A. Akinrotimi, O. M.

Edun, and Oik Ukwe. Effects of anaesthetics on metabolic enzyme activities in African catfish, Clarias gariepinus (Burchell, 1822). Journal of FisheriesSciences.com, 12 (1):??, ???? 2018. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/effects-
of-anaesthetics-on-metabolic-
enzyme-activities-in-african-
catfish-clarias-gariepinus-
burchell-1822.php?aid=
22021.

Akinyemi:2011:ASO

[AF11] Adeolu Akanji Akinyemi and Olufemi Paul Fajana. Antibiotics sensitivity and occurrance of bacteria originated from skin, gill and buccal cavity of Chrysicthys nigrodigitatus, Distichodus engycephalus and Synodontis schall sampled in Arakanga Reservoir, Abeokuta, Ogun State, Nigeria. Journal of FisheriesSciences.com, 5 (2):164–171, ???? 2011.

CODEN JFOIAQ. ISSN 1307-234X. URL https:

                                        //www.fisheriessciences.
                                        com/fisheries-aqua/.

Akcaalan:2011:NCB

[AİGS11] Reyhan Akçaalan, Melek İşinibilir, Cenk Gürevin, and Adnan Sümer. A new contribution of biodiversity of Sapanca Lake: Craspedacusta sowerbyi Lankester, 1880 (Cnidaria: Hydrozoa). Journal of FisheriesSciences.com, 5(1):??, ???? 2011. CODEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/a-new-contribution-
                                        of-biodiversity-of-sapanca-
                                        lake-craspedacusta-sowerbyi-
                                        lankester-1880-cnidaria-
                                        hydrozoa.php?aid=817.

Ahmad:2017:RHT

[AK17] Ishtiyaq Ahmad and Harpreet Kaur. Redescription and histopathology of two species of myxozoans infecting gills of fingerlings of Indian major carps. Journal of FisheriesSciences.com, 11 (1):??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/redescription-
                                        and-histopathology-of-
                                        two-species-of-myxozoans-
                                        infecting-gills-of-fingerlings-
                                        of-indian-major-carps.
                                        php?aid=17324.

Akamca:2010:SMT

[Aka10] Erhan Akamca. The selectivity of monofilament trammel nets used in catch of gilt head bream (Sparus aurata, Linneaus, 1758) in İskenderun Bay–Turkey. Journal of FisheriesSciences.com, 4(1): ??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
selectivity-of-monofilament-
trammel-nets-used-in-catch-
of-gilt-head-bream-sparus-
aurata-linneaus-1758-in-
skenderun-bayturkey.php?
aid=1090.

Ak:2011:TKD

[AKA11a] Orhan Ak, Sebahattin Kutlu, and Ilhan Aydin. Trabzon kiyilarinda demersal tür dağilimi ve ekonomik baliklarin yoğunluk miktari. (Turkish) [Distribution of demersal species and density of the economic fishes in the coastal Trabzon]. Journal of FisheriesSciences.com, 5(??):99–106, 2011. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/distribition-
of-demersal-species-and-
density-of-the-economic-
fishes-in-the-coastal-
trabzon.php?aid=776.

Akalin:2011:IKI

[Aka11b]  Sencer Akalın.  izmir

körfezi’nde iki iskorpit türünün (scorpaena notata and scorpaena porcus) boy-ağırlık ilişkisi ve metrik-meristik Özellikleri. Journal of FisheriesSciences.com, 5(4): ??, ???? 2011. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/lengthweight-
                                        relationship-and-metricmeristic-
                                        characteristics-of-two-
                                        scorpion-fishes-scorpaena-
                                        notata-and-scorpaena-porcus-
                                        in-304zmir-bay.php?aid=
                                        720.

Akar:2013:BTU

[Aka13]  Mustafa Akar.  Bayes

teorisinin su Ürünlerinde kullanım olanakları. Journal of FisheriesSciences.com, 7 (??):??, 2013. CODEN JFOIAQ. ISSN 1307-234X.

Akbulut:2009:HDF

[Akb09] Bilal Akbulut. Historical development and future projection of aquaculture in the Black Sea Region. Journal of FisheriesSciences.com, 3 (2):??, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/historical-
                                        development-and-future-
                                        projection-of-aquaculture-
                                        in-the-black-sea-region.
                                        php?aid=1489.

Akbulut:2010:GPJ

[Akb10] Bilal Akbulut. Growth performance of juvenile beluga (Huso huso), captured from Black Sea, fed with whiting in tanks. Journal of FisheriesSciences.com, 4 (2):??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/growth-
performance-of-juvenile-
beluga-huso-huso-captured-
from-black-sea-fed-with-
whiting-in-tanks.php?aid=
1031.

Akbulut:2012:ROC

[Akb12] Bilal Akbulut. Routine oxygen consumption rate of the black sea trout (Salmo trutta labrax Pallas, 1811). Journal of FisheriesSciences.com, 6(2):??, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/routine-
oxygen-consumption-rate-
of-the-black-sea-troutsalmo-
trutta-labrax-pallas-1811.
php?aid=657.

Akbulut:2013:EFR

[Akb13] Bilal Akbulut. Effect of feeding rate on growth performance, food utilization and meat yield of sterlet (Acipenser ruthenus Linne, 1758). Journal of FisheriesSciences.com, 7(4):??, ???? 2013. CODEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/

                                        fisheries-aqua/effect-
                                        of-feeding-rate-on-growth-
                                        performance-food-utilization-
                                        and-meat-yield-of-sterlet-
                                        acipenser-ruthenus-linne-
                                        1758.php?aid=454.

Abujam:2017:EFA

[AKD^{+}17] Santoshkumar Abujam, Ram Kumar, Achom Darshan, Budhin Gogoi, and Debangshu Narayan Das. Enumeration of fish and alimentation of fisherman at Tezu River in Lohit District of Arunachal Pradesh, India. Journal of FisheriesSciences.com, 11 (3):??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/enumeration-
                                        of-fish-and-alimentation-
                                        of-fisherman-at-tezu-river-
                                        in-lohit-district-of-arunachal-
                                        pradesh-india.php?aid=
                                        20092.

Abujam:2017:CRH

[AKDD17] Santosh Kumar Abujam, Ram Kumar, Achom Darshan, and Debangshu Narayan Das. Captive rearing of hill stream ornamental fishes of Arunachal Pradesh, Northeast India. Journal of FisheriesSciences.com, 11 (1):??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/captive-
                                        rearing-of-hill-stream-
                                        ornamental-fishes-of-arunachal-

pradesh-northeast-india.
php?aid=17964.

Akamca:2009:SMT

[AKÖ09] Erhan Akamca, Volkan Barış Kiyağa, and Caner Enver Özyurt. The selectivity of monofilament trammel nets used in catch of gilt head bream (Sparus aurata, Linneaus, 1758) in İskenderun Bay–Turkey. Journal of FisheriesSciences.com, 3(4): ??, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
selectivity-of-monofilament-
trammel-nets-used-in--
catch-of-gilt-head-bream-
sparus-aurata-linneaus-
1758-in-skenderun-bayturkey.
php?aid=1142.

Aksu:2011:OKA

[Aks11a] Hakan Aksu. Orta karadeniz’de avlanan barbunya (Mullus barbatus ponticus, Essipov, 1927) balıklarının bazı populasyon parametreleri. Journal of Fisheries-Sciences.com, 5(??):??, 2011. CODEN JFOIAQ. ISSN 1307-234X.

Aksu:2011:ESP

[Aks11b] Harun Aksu. Estimation of some population parameters of red mullet (Mullus barbatus ponticus, Essipov, 1927) caught in the Black Sea. Journal of FisheriesSciences.com, 5(4): ??, ???? 2011. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/estimation-
of-some-population-parameters-
of-red-mullet-mullus-barbatus-
ponticus-essipov-1927-
caught-in-the-black-sea.
php?aid=740.

Aktar:2014:TMC

[Akt14]  Mahbuba Aktar.  Trace

metal concentrations in the green-lipped mussel Perna viridis (Linnaeus, 1758) collected from Maheshkhali Channel, Cox’s Bazar, Bangladesh. Journal of FisheriesSciences.com, 8(1):??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/trace-metal-
concentrations-in-the-
greenlipped-mussel-perna-
viridis-linnaeus-1758-
collected-from-maheshkhali-
channel-coxs-bazar-ban.
php?aid=397.

A:2017:EBB

[AKV^{+}17] Ashwitha A., Thamizharasan K., Vithya V., Karthik R., and Vijaya Bharathi S. Effectiveness of bacteriocin from Bacillus subtilis (KY808492) and its application in biopreservation. Journal of FisheriesSciences.com, 11(3):??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/effectiveness-
of-bacteriocin-from-bacillus-
subtilis-ky808492-and-
its-application-in-biopreservation.
php?aid=19575.

Akyol:2011:CFF

[Aky11a] Okan Akyol. Coastal fisheries and fishing resources of Bozcaada (Aegean Sea). Journal of FisheriesSciences.com, 5(1):??, ???? 2011. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/coastal-
fisheries-and-fishing-
resources-of-bozcaada-
aegean-sea.php?aid=807.

Akyol:2011:PRC

[Aky11b]  Okan Akyol.  Prelimi-

nary results on the cuttlefish, Sepia officinalis, reproduction in Izmir Bay (Aegean Sea). Journal of FisheriesSciences.com, 5(2): ??, ???? 2011. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/preliminary-
results-on-the-cuttlefish-
sepia-officinalis-reproduction-
in-izmir-bay-aegean-sea.
php?aid=771.

Alhabsi:2012:FCA

[Alh12] Manaa Saif Alhabsi. The fisheries community of Albatinah Region in Oman: a socio-economic overview. Journal of FisheriesSciences.com, 6(3):??, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
fisheries-community-of-
albatinah-region-in-oman-
a-socioeconomic-overview.
php?aid=602.

Alparslan:2013:OSC

[Alp13] Yunus Alparslan. Ortam şartlarinda (23 ± 4C) birden fazla uygulanan çözündürme işleminin levrek baliği (Dicentrarchus labrax, L., 1758)’nin et kalitesine etkisi. (Turkish) [Effect of multiple thawing process in ambient conditions (23 ± 4C) on quality of sea bass (Dicentrarchus labrax, L., 1758)]. Journal of FisheriesSciences.com, 7(1):??, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/effect-
of-multiple-thawing-process-
in-ambient-conditions-
234c-on-quality-of-sea-
bass-dicentrarchus-labrax-
l-1758.php?aid=557.

Altinterim:2013:ENS

[Alt13] Başar Altınterim. The effects of Nigella sativa oil on the immune system of rainbow trout with different application methods. Journal of FisheriesSciences.com, 7 (3):??, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
effects-of-nigella-sativa-
oil-on-the-immune-system-
of-rainbow-trout-with-
different-application-
methods.php?aid=439.

Anita:2014:KTI

[Ani14] Bhatnagar Anita. Karyomorphology of three Indian major carps from Haryana, India. Journal of FisheriesSciences.com, 8 (2):??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/karyomorphology-
of-three-indian-major-
carps-from-haryana-india.
php?aid=252.

Alcorlo:2019:MPD

[ANM19] Paloma Alcorlo, Victor Noguerales, and Salvador Mollá. Modeling the population dynamics of the red swamp crayfish (Procambarus clarkii) in Doñana: Application to the harvesting strategies. Journal of FisheriesSciences.com, 13 (2):1–11, ???? 2019. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/modeling-
the-population-dynamics-

                                        of-the-red-swamp-crayfish-
                                        procambarus-clarkii-in-
                                        doana-application-to-the-
                                        harvesting-strategies.
                                        php?aid=24370.

Anonymous:2020:JFC

[Ano20a]  Anonymous.  Journal of

FisheriesSciences.com: Editor note. Journal of FisheriesSciences.com, 14 (3):??, ???? 2020. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/journal-
                                        of-fisheriessciencescom-
                                        editor-note.pdf.

Anonymous:2020:ENJ

[Ano20b] Anonymous. Editor note: Journal of FisheriesSciences.com. Journal of FisheriesSciences.com, 14(3):??, ???? 2020. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/editor-
                                        note-journal-of-fisheriessciencescom.
                                        pdf.

Akpaniteaku:2019:ESD

[AO19] Rupert C. Akpaniteaku and Michael O. Oguayo. Effect of stocking density on survival and growth of Potamon ebonyicum (freshwater crab). Journal of FisheriesSciences.com, 13 (2):??, ???? 2019. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/

fisheries-aqua/effect-
of-stocking-density-on-
survival-and-growth-of-
potamon-ebonyicum-freshwater-
crab.php?aid=24441.

Ajani:2016:CBF

[AOO16] E. K. Ajani, O. Orisasona, and B. O. Omitoyin. Comparision of bacterial flora and frequency of occurance in water and Clarias gariepinus raised in ponds fertilized with raw with poultry manure. Journal of FisheriesSciences.com, 10 (1):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/comparision-
of-bacterial-flora-and-
frequency-of-occurance-
in-water-and-clarias-gariepinus-
raised-in-ponds-fertilized-
with-raw-with.php?aid=
8211.

Aydin:2010:STN

[AS10] Mehmet Aydın and Çetin Sümer. Selectivity of trammel nets used for common dentex (Dentex dentex) fishery in the South Aegean. Journal of FisheriesSciences.com, 4(4):??, ???? 2010. CODEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/selectivity-
of-trammel-nets-used-for-
common-dentex-dentex-dentex-

                                        fishery-in-the-south-aegean.
                                        php?aid=841.

Alpaslan:2012:VSC

[AST^{+}12] Kenan Alpaslan, Ahmet Sesli, Rıdvan Tepe, Nurten Özbey, Nevin Birici, Tünay Şeker, and Mehmet Ali Turan Koçer. Vertical and seasonal changes of water quality in Keban Dam Reservoir. Journal of FisheriesSciences.com, 6(3):252– 262, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/vertical-
                                        and-seasonal-changes-of-
                                        water-quality-in-keban-
                                        dam-reservoir.php?aid=
                                        598.

Abbas:2021:ENA

[ASTH21] Mahmoud Mahrous M. Abbas, Sabry M. Shehata, Abdelrahman S. Talab, and Mohamed H. Effect of natural additives on the quality attributes of cultivated common carp fillets during frozen storage. Journal of FisheriesSciences.com, 15 (2):3–13, ???? 2021. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/effect-
                                        of-natural-additives-on-
                                        the-quality-attributes-
                                        of-cultivated-common-carp-
                                        fillets-during-frozen-
                                        storage.php?aid=36171.

Adedokun:2017:AOR

[ATAA17] M. A. Adedokun, H. M.

Tairu, O. Adeosun, and O. Ajibola. Assessment of the optimal replacement levels of maize with water lettuce leaf (Pistia stratiotes) based diets for Clarias gariepinus. Journal of FisheriesSciences.com, 11 (2):??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/assessment-
of-the-optimal-replacement-
levels-of-maize-with-water-
lettuce-leaf-pistia-stratiotes-
based-diets-for-clarias-
gariepinus.php?aid=18455.

Abo-Taleb:2016:DME

[ATEA^{+}16] Hamdy A. Abo-Taleb, Sawsan M. Aboul Ezz, Nagwa E. Abdel Aziz, Mohamed M. Abou Zaid, and El Raey M. Detecting marine environmental pollution by biological beacons and GIS program. Journal of FisheriesSciences.com, 10(4):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/fisheries-
assessment-of-chwaka-bay-
zanzibarfollowing-a-holistic-
approach.php?aid=16503.

Avsever:2014:ICB

[ATG^{+}14] Meriç Lütfi Avsever, T. Tansel Tanrıkul, Derya Güroy, Seçil Metin, Hasan Akşit, and Serra Tunalıgil. Investigation of certain blood parameters in rainbow trout (Oncorhynchus mykiss Walbaum, 1792) naturally infected with Lactococcus garvieae. Journal of FisheriesSciences.com, 8(2):114– 120, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/investigation-
of-certain-blood-parameters-
in-rainbow-trout-oncorhynchus-
mykiss-walbaum-1792-naturally-
infected-with-lactococcus-
garvieae.php?aid=257.

Atilgan:2012:DKI

[Atı12] Elvan Atılgan. Doğu karadeniz’deki istavrit, Trachurus mediterraneus (Steindachner, 1868)’in otolit Özellikleri ve bazı populasyon parametreleri. Journal of FisheriesSciences.com, 6(??):??, 2012. CODEN JFOIAQ. ISSN 1307-234X.

Ayas:2010:DMC

[Aya10] Deniz Ayas. Distribution and morphometric characteristics of patella species (Archaeogastropoda) in Mersin– Viranşehir region of the Northeastern Mediterranean Sea. Journal of FisheriesSciences.com, 4(2):??, ???? 2010. CODEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/

fisheries-aqua/distribution-
and-morphometric-characteristics-
of-patella-species-archaeogastropoda-
in-mersinviranehir-region-
of-the-northeastern-m.
php?aid=1036.

Ayas:2011:CCC

[Aya11a] Deniz Ayas. The chemical composition of carapace meat of sexually mature swimmer blue crab (Callinectes sapidus, Rathbun 1896) in the Mersin Bay. Journal of FisheriesSciences.com, 5(4):308– 316, ???? 2011. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
chemical-composition-of-
sexually-mature-blue-swimmer-
crab-portunus-pelagicus-
linnaeus-1758-in-the-mersin-
bay.php?aid=714.

Ayas:2011:CCS

[Aya11b] Deniz Ayas. The chemical composition of sexually mature blue swimmer crab (Portunus pelagicus, Linnaeus 1758) in the Mersin Bay. Journal of FisheriesSciences.com, 5(??):??, 2011. CODEN JFOIAQ. ISSN 1307-234X.

Aydin:2008:DSE

[Ayd08a]  Celalettin Aydın.  De-

termined species escaped from codend, underneath the ootrope and above the headline model in a bottom trawl net. Journal of FisheriesSciences.com, 2 (2):??, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/determined-
species-escaped-from-codend-
underneath-the--ootrope-
and-above-the-headline-
model-in-a-bottom-trawl-
net.php?aid=1642.

Aydin:2008:ETD

[Ayd08b] Ilker Aydın. Effects of time of the day on catch composition in monofilament and multifilament gillnets fishing. Journal of FisheriesSciences.com, 2(3):??, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/effects-
of-time-of-the-day-on-
catch-composition-in-monofilament-
and-multifilament-gillnets-
fishing.php?aid=1607.

Aydin:2009:ADM

[Ayd09]  Rahmi Aydın.  Age de-

termination of mirror carp (Cyprinus carpio L.,1758) on otoliths with broken and burnt method. Journal of FisheriesSciences.com, 3 (1):??, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/age-
determination-of-mirror-

carp-cyprinus-carpio-l1758-
on-otoliths-with-broken-
and-burnt-method.php?aid=
1502.

Aydin:2010:APG

[Ayd10a] Hamdi Aydın. Aquaculture potential in Gümüşhane (Turkey) and its assesment. Journal of FisheriesSciences.com, 4(2):??, ???? 2010. CODEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/aquaculture-
potential-in-gmhane-turkey-
and-its-assesment.php?
aid=1056.

Aydin:2010:OEF

[Ayd10b] İlhan Aydın. Otolith extraction in fish larvae and daily increments: Example of ayu, (Plecoglossus altivelis Temminck & Schlegel, 1846). Journal of FisheriesSciences.com, 4(2):??, ???? 2010. CODEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/otolith-
extraction-in-fish-larvae-
and-daily-increments-example-
of-ayu-plecoglossus-altivelis-
temminck--schlegel-1846.
php?aid=1043.

Aydin:2011:EFF

[Ayd11a] İlhan Aydın. The effect of feeding frequency and feeding rate on growth performance of juvenile black sea turbot (Psetta maxima, Linneaus, 1758). Journal of FisheriesSciences.com, 5 (1):??, ???? 2011. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
effect-of-feeding-frequency-
and-feeding-rate-on-growth-
performance-of-juvenile-
black-sea-turbot-psetta-
maxima-linneaus-1758.php?
aid=820.

Aydin:2011:EIT

[Ayd11b] İlhan Aydın. Effect of iodine treatment on the hatching rate of black sea turbot, (Psetta maxima Linnaeus, 1758), eggs. Journal of FisheriesSciences.com, 5(1):??, ???? 2011. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/effect-
of-iodine-treatment-on-
the-hatching-rate-of-black-
sea-turbot-psetta-maxima-
linnaeus-1758-eggs.php?
aid=805.

Aydin:2011:ESD

[Ayd11c] İlhan Aydın. The effect of sexual dimorphism on growth of the black sea turbot, Psetta maxima. Journal of FisheriesSciences.com, 5(1):??, ???? 2011. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-

effect-of-sexual-dimorphism-
on-growth-of-the-black-
sea-turbot-psetta-maxima.
php?aid=815.

Aydin:2011:KYP

[Ayd11d] İlhan Aydın. Kuluçkahanede yetiştirilen pisi balığı (Platichthys flesus luscus Pallas, 1814)’nin spermatolojik özellikleri. (Turkish) [A study on the spermatological characteristics of hatchery-reared flounder (Platichthys flesus luscus Pallas, 1814)]. Journal of FisheriesSciences.com, 5(4):??, ???? 2011. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/a-study-
on-the-spermatological-
characteristics-of-hatcheryreared-
flounder-platichthys-flesus-
luscus-pallas-1814.php?
aid=725.

Aydin:2012:CGP

[Ayd12] Rahmi Aydın. Comparison of the growth parameters of Capoeta trutta (Heckel, 1843) living in Keban and Karakaya Dam Lake. Journal of FisheriesSciences.com, 6(4):??, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/comparison-
of-the-growth-parameters-
of-capoeta-trutta-heckel-
1843-living-in-keban-and-
karakaya-dam-lake.php?
aid=572.

Aydin:2014:DTB

[Ayd14a]  Celalettin Aydın.  Dip

trol balıkçılığında barbunya (Mullus barbatus) ve isporuz (Diplodus annularis) boy seçiciliğinin geliştirilmesi. (Turkish) [Improving size selectivity of red mullet (Mullus barbatus) and annular sea bream (Diplodus annularis) in bottom trawl net]. Journal of FisheriesSciences.com, 8(1):??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/improving-
                                        size-selectivity-of-red-
                                        mullet-mullus-barbatus-
                                        and-annular-sea-bream-
                                        diplodus-annularis-in-
                                        bottom-trawl-net.php?aid=
                                        404.

Aydin:2014:SDH

[Ayd14b] Cokun Menderes Aydın. The selectivity of different hook size in the bottom longline which used for catching white seabream (Diplodus sargus L., 1758) in Finike Bay. Journal of FisheriesSciences.com, 8(3): ??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

                                        //www.fisheriessciences.
                                        com/fisheries-aqua/the-
                                        selectivity-of-different-
                                        hook-size-in-the-bottom-
                                        longline-which-used-for-
                                        catching-white-seabream-

diplodus-sargus-l-1758-
in-fin.php?aid=65.

Boecker:2016:CTU

[BA16] Andreas Boecker and Daniele Asioli. Could there be unintended effects of government support for seafood traceability implementation on business planning? Results of a survey among Italian fishery businesses. Journal of FisheriesSciences.com, 10(3):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/could-there-
be-unintended-effects-
of-government-support-
for-seafood-traceability-
implementation-on-business-
planning-results-of-a.
php?aid=11142.

Bacinar:2012:OCS

[Baç12] Nimet Selda Baçinar. Otolith characteristics and some population parameters of eastern black sea horse mackerel, Trachurus mediterraneus (Steindachner, 1868). Journal of FisheriesSciences.com, 6(2):??, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/otolith-
characteristics-and-some-
population-parameters-
of-eastern-black-sea-horse-
mackerel-trachurus-mediterraneus-

                                        steindachner-1868.php?
                                        aid=653.

Bahar:2007:EFT

[Bah07]  Tokur Bahar.  The ef-

fects of Fenton type (Fe^{+2} /H_{2}O_{2}) oxidation system on lipid and protein oxidation of grey mullet (Mugil cephalus). Journal of FisheriesSciences.com, 1(1):41– 47, February 2007. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

                                        //www.fisheriessciences.
                                        com/fisheries-aqua/the-
                                        effects-of-fenton-type-
                                        fe2h2o2-oxidation-system-
                                        on-lipid-and-protein-oxidation-
                                        of-grey-mullet-mugil-cephalus.
                                        php?aid=1761.

Balik:2007:SDEa

[Bal07] Ismet Balik. Spatial distributions of economic fish populations in Lake Egirdir. Journal of FisheriesSciences.com, 1(2):88–96, April 2007. CO-DEN JFOIAQ. ISSN 1307-234X.

Basaran:2009:SBS

[Baş09]  Fatih Başaran.  Spawn-

ing behaviour of shi drum (Umbrina cirrosa) after hormone administration. Journal of FisheriesSciences.com, 3(2):??, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/spawning-
                                        behaviour-of-shi-drum-
                                        umbrina-cirrosa-after-

hormone-administration.
php?aid=1481.

Bascinar:2011:DGE

[Baş11a] Nimet Selda Başçınar. Determination of gastric evacuation time and rate of black sea trout (Salmo trutta labrax Pallas, 1811) larvae through live feed utilization. Journal of FisheriesSciences.com, 5(4):??, ???? 2011. CODEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/determination-
of-gastric-evacuation-
time-and-rate-of-black-
sea-trout-salmo-trutta-
labrax-pallas-1811-larvae-
through-live-feed-utilization.
php?aid=717.

Basiacik:2011:ABG

[Baş11b] Seda Başıaçık. Adıgüzel baraj gölü (denizli) bıyıklı balık (Luciobarbus kottelati Turan, Ekmekçi, İlhan & Engin, 2008) populasyonunun bazı büyüme Özellikleri. Journal of FisheriesSciences.com, 5(??):??, 2011. CODEN JFOIAQ. ISSN 1307-234X.

Bat:2007:CEB

[Bat07] Levent Bat. The changed ecosystem of the Black Sea and its impact on anchovy fisheries. Journal of FisheriesSciences.com, 1(4):191– 226, 2007. CODEN JFOIAQ. ISSN 1307-234X.

Bulut:2012:ICG

[BAU^{+}12] Cafer Bulut, Ufuk Akçimen, Kazim Uysal, Şakir Çinar, Ramazan Küçükkara, and Soner Savaser. Isparta çandir göksu kaynaği Üzerindeki alabalik işletmelerinin dere suyuna olan etkileri. (Turkish) [The effects of trout aquaculture facility on the source of Isparta Çandır Göksu to the creek water]. Journal of FisheriesSciences.com, 6(4):331– 340, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
effects-of-trout-aquaculture-
facility-on-the-source-
of-isparta-and305r-gksu-
to-the-creek-water.php?
aid=570.

Bausta:2013:MNK

[Bau13]  Asiye Bausta.  Munzur

nehri’ndeki kırmızı benekli alabalığın (Salmo trutta macrostigma (Dummeril., 1858)) otolit boyutları-balık boyu arasındaki ilişki. (Turkish) [Otolith dimensionsfish length relations between of red-spotted trout (Salmo trutta macrostigma (Dummeril, 1858)) in Munzur River]. Journal of FisheriesSciences.com, 7(1): ??, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/

fisheries-aqua/otolith-
dimensionsfish-length-
relations-between-of-redspotted-
trout-salmo-trutta-macrostigma-
dummeril-1858-in-munzur-
river.php?aid=553.

Bayhan:2012:ABG

[Bay12] Yusuf Kenan Bayhan. Atatürk baraj gölü (adıyaman) balıkçılığı ve kullanılan av araçlarının teknik özellikleri. (Turkish) [The fisheries in Ataturk Dam Lake (Adiyaman) and technical features of the fishing gears being used]. Journal of FisheriesSciences.com, 6(3):??, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
fisheries-in-ataturk-dam-
lake-adiyaman-and-technical-
features-of-the-fishing-
gears-being-used.php?aid=
600.

Balik:2007:SDEb

[BÇÖ07] İsmet Balık, Hıdır Çubuk, and Remziye Özkök. Spatial distributions of economic fish populations in Lake Eğirdir. Journal of FisheriesSciences.com, 1(2): ??, ???? 2007. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/spatial-
distributions-of-economic-
fish-populations-in-lake-
e287irdir.php?aid=1755.

Bilen:2010:SBA

[BDAK10] Canan Türeli Bilen, Önder Duysak, Erhan Akamca, and Volkan Kiyağa. Some biological aspects of cuttlebone (Sepia officinalis L., 1758) of the coastal area of Karataş (Northeastern Mediterranean–Turkey). Journal of FisheriesSciences.com, 4(4):??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

                                        //www.fisheriessciences.
                                        com/fisheries-aqua/some-
                                        biological-aspects-of-
                                        cuttlebone-sepia-officinalis-
                                        l-1758-of-the-coastal-
                                        area-of-karata-northeastern-
                                        mediterraneanturkey.php?
                                        aid=846.

Borchert:2020:RPI

[BDB20] Amanda Borchert, Jacob L. Davis, and Michael E. Barnes. Retention of passive integrated transponders (PIT) tags in juvenile rainbow trout and brown trout. Journal of FisheriesSciences.com, 14 (1):??, ???? 2020. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/retention-
                                        of-passive-integrated-
                                        transponders-pit-tags-
                                        in-juvenile-rainbow-trout-
                                        and-brown-trout.pdf.

Biehl:2019:ICT

[BDdF19] Thayna M. Biehl, David V.

Dantas, and Eduardo Guilherme Gentil de Farias. Influence of crab trap geometry and wind direction over the catch fluctuations in a neotropical coastal lagoon. Journal of FisheriesSciences.com, 13 (1):??, ???? 2019. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/influence-
of-crab-trap-geometry-
and-wind-direction-over-
thecatch-fluctuations-
in-a-neotropical-coastal-
lagoon.php?aid=24209.

Behera:2014:CLD

[Beh14] Durga Prasad Behera. A check list on distribution of ornamental fishes in Chilika Lagoon, East Coast of India. Journal of FisheriesSciences.com, 8(1):??, ???? 2014. CODEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/a-check-
list-on-distribution-of-
ornamental-fishes-in-chilika-
lagoon-east-coast-of-india.
php?aid=399.

Berber:2009:LWR

[Ber09] Selçuk Berber. The length– weight relationships, and meat yield of crayfish (Astacus leptodactylus Eschcholtz, 1823) population in Apolyont Lake (Bursa, Turkey). Journal of FisheriesSciences.com, 3(2):??, ???? 2009. CODEN JFOIAQ. ISSN 1307-234X. URL https:/

                                        /www.fisheriessciences.
                                        com/fisheries-aqua/the-
                                        lenghtweight-relationships-
                                        and-meat-yield-of-crayfish-
                                        astacus-leptodactylus-
                                        eschcholtz-1823-population-
                                        in-apolyont-lake-bursa.
                                        php?aid=1487.

Berillis:2015:FCL

[Ber15] Panagiotis Berillis. Factors that can lead to the development of skeletal deformities in fishes: a review. Journal of FisheriesSciences.com, 9(3):??, ???? 2015. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/factors-
                                        that-can-lead-to-the-development-
                                        of-skeletaldeformities-
                                        in-fishes-a-review.php?
                                        aid=6678.

Berillis:2017:ENF

[Ber17a]  Panagiotis Berillis.  Ed-

itor note — FisheriesSciences. Journal of FisheriesSciences.com, 11(3): ??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/Editor-
                                        note-fisheriessciences.
                                        php?aid=21555.

Berillis:2017:SDS

[Ber17b] Panagiotis Berillis. Skeletal deformities in seabreams. understanding the genetic origin can improve production? Journal of FisheriesSciences.com, 11 (2):??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/skeletal-
deformities-in-seabreams-
understanding-the-genetic-
origin-can-improve-production.
php?aid=18654.

Berillis:2021:EOE

[Ber21a]  Panagiotis Berillis.  Edi-

torial on overview of an Escox. Journal of FisheriesSciences.com, 15(1):1, ???? 2021. CODEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/editorial-
on-overview-of-an-escox.
php?aid=34743.

Berillis:2021:NGR

[Ber21b] Panagiotis Berillis. A note on Ganges River dolphin (Platanista gangetica). Journal of FisheriesSciences.com, 15 (2):1–2, ???? 2021. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/a-note-
on-ganges-river-dolphin-
platanista-gangetica.php?
aid=36170.

Berillis:2021:PFL

[Ber21c]  Panagiotis Berillis.  Per-

spective for frameworks of lymphocystis disease in fish. Journal of FisheriesSciences.com, 15(1):1, ???? 2021. CODEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/perspective-
                                        for-frameworks-of-lymphocystis-
                                        disease-in-fish.php?aid=
                                        34744.

Beyhan:2010:KAK

[BG10] Yusuf Kenan Beyhan and Gökhan Gökçe. Kuzeydoğu akdeniz’de karides balikçiliği ve kullanilan ağlarin teknik özellikleri. (Turkish) [shrimp fishery in northeast Mediterranean Sea and technical specifications of the shrimp nets]. Journal of FisheriesSciences.com, 4(2):129– 135, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/shrimp-
                                        fishery-in-northeast-mediterranean-
                                        sea-and-technical-specifications-
                                        of-the-shrimp-nets.php?
                                        aid=1046.

Bhattacharya:2016:DFG

[Bha16] Shreya Bhattacharya. Doctor fish Garra rufa: Health and risk. Journal of FisheriesSciences.com, 10(1): ??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/doctor-
fish-garra-rufa-health-
and-risk.php?aid=8208.

Bilgin:2007:ICC

[Bil07] Šengül Bilgin. Investigation on changes in the chemical composition of hot smoked Salmo trutta macrostigma, Dumeril 1858, stored different temperatures. Journal of FisheriesSciences.com, 5 (4):68–80, April 2007. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/investigation-
on-changes-in-the-chemical-
composition-of-hot-smoked-
salmo-trutta-macrostigma-
dumeril-1858-stored-different-
temperatures.php?aid=746.

Bao-Ji:2019:RMR

[BJSN19] Zhang Bao-Ji and Ma Song-Nan. Research on motion response and sickness incidence of the fishing boat in heading and quartering seas. Journal of FisheriesSciences.com, 13(2):??, ???? 2019. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/research-
on-motion-response-and-
sickness-incidence-of-
the-fishing-boat-in-heading-
and-quartering-seas.php?
aid=24458.

Baki:2015:SFP

[BKB15] Birol Baki, Ali Karaçuha, and Hakan Baki. A study on feeding preferences of juvenile flounder (Platichthys flesus). Journal of FisheriesSciences.com, 9(2):??, ???? 2015. CODEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/a-study-
                                        on-feeding-preferences-
                                        of-juvenileflounder-platichthys-
                                        flesus.php?aid=6448.

Bhatnagar:2017:MCD

[BL17] Anita Bhatnagar and Ritu Lamba. Molecular characterization and dosage application of autochthonous potential probiotic bacteria in Cirrhinus mrigala. Journal of FisheriesSciences.com, 11 (2):??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/molecular-
                                        characterization-and-dosage-
                                        application-ofautochthonous-
                                        potential-probiotic-bacteria-
                                        in-cirrhinus-mrigala.php?
                                        aid=18630.

Berillis:2017:HGC

[BM17] Panagiotis Berillis and Eleni Mente. Histology of goblet cells in the intestine of the rainbow trout can lead to improvement of the feeding management. Journal of FisheriesSciences.com, 11 (4):??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/histology-
of-goblet-cells-in-the-
intestine-of-the-rainbow-
trout-can-lead-to-improvement-
of-the-feeding-management.
php?aid=21556.

Bessadok:2017:CCA

[BMBS17] Boutheina Bessadok, Mahmoud Masri, Thomas Breuck, and Saloua Sadok. Characterization of the crude alkaline extracellular protease of Yarrowia lipolytica YlTun15. Journal of FisheriesSciences.com, 11 (4):??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/characterization-
of-the-crude-alkaline-
extracellular-protease-
of-yarrowia-lipolytica-
yltun15.php?aid=20978.

Berillis:2017:UCA

[BMK17] Panagiotis Berillis, Elena Mente, and Konstantinos A. Kormas. The use of copper alloy in aquaculture fish net pens: Mechanical, economic and environmental advantages. Journal of FisheriesSciences.com, 11 (4):??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.

                                        com/fisheries-aqua/the-
                                        use-of-copper-alloy-in-
                                        aquaculture-fish-net-pens-
                                        mechanical-economic-and-
                                        environmental-advantages.
                                        php?aid=20468.

Benali:2020:CSG

[BMKS20] Sid Ahmed Benali, Ghoumari Sidi Mohamed, Larbi Doukara Kamel, and Gaouar Semir Bechir Suheil. Contribution to the study of growth of Luciobarbus mascarensis (Teleostei: Cyprinidae) in two freshwater systems in Western Algeria. Journal of FisheriesSciences.com, 14 (4):??, ???? 2020. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/contribution-
                                        to-the-study-of-growth-
                                        of-luciobarbus-mascarensis-
                                        teleostei-cyprinidae-in-
                                        two-freshwater-systems-
                                        in-westernalgeria.pdf.

Bostanci:2007:ESD

[Bos07] Derya Bostanci. The evaluation of the similarity and difference on the ages of different bony structures of the Crucian carp (Carassius gibelio Bloch, 1782). Journal of FisheriesSciences.com, 1 (??):124–129, 2007. CODEN JFOIAQ. ISSN 1307-234X.

Bostanci:2008:RCB

[Bos08] Derya Bostanci. The ring characteristics on bony structures for age determination of fish. Journal of FisheriesSciences.com, 2(2): ??, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
ring-characteristics-on-
bony-structures-for-age-
determination-of-fish.
php?aid=1633.

Bostanci:2009:RWR

[Bos09] Derya Bostanci. The ring which is not represent the true age on bony structures of known age fishes. Journal of FisheriesSciences.com, 3(2):??, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/the-ring-
which-is-not-represent-
the-true-age-on-bony-structures-
of-known-age-fishes.php?
aid=1482.

Bozkurt:2009:CCF

[Boz09a]  Ahmet Bozkurt.  Clado-

cera and Copepoda fauna of Aslantaş Dam Lake (Osmaniye–Turkey). Journal of FisheriesSciences.com, 3(4):??, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/cladocera-
and-copepoda-fauna-of-
aslanta-dam-lake-osmaniyeturkey.
php?aid=1165.

Bozkurt:2009:CVD

[Boz09b] Ahmet Bozkurt. Composition and vertical distribution of rotifera in Aslantas Dam Lake (Osmaniye– Turkey). Journal of FisheriesSciences.com, 3(4):??, ???? 2009. CODEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/composition-
                                        and-vertical-distribution-
                                        of--rotifera-in-aslantas-
                                        dam-lake-osmaniyeturkey.
                                        php?aid=1141.

Bozkurt:2010:CVD

[Boz10a] Ahmet Bozkurt. Composition and vertical distribution of rotifera in Aslantas Dam Lake (Osmaniye– Turkey). Journal of FisheriesSciences.com, 4(1):??, ???? 2010. CODEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/composition-
                                        and-vertical-distribution-
                                        of-rotifera-in-aslantas-
                                        dam-lake-osmaniyeturkey.
                                        php?aid=1089.

Bozkurt:2010:ZSA

[Boz10b] Ahmet Bozkurt. Zooplankton succession of the Asi River (Hatay–Turkey). Journal of FisheriesSciences.com, 4(4):??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/

fisheries-aqua/zooplankton-
succession-of-the-asi-
river-hatayturkey.php?
aid=857.

Braimah:2019:ATE

[Bra19] Lawrence Issah Braimah. Assessment of technical and economic efficiencies of gillnet types with different as against the usual same, hanging ratios at the top and down ropes, compared with a standard gillnet, in the Volta Lake, Ghana. Journal of FisheriesSciences.com, 13(4):??, ???? 2019. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/assessment-
of-technical-and-economic-
efficiencies-of-gillnet-
types-with-different-as-
against-the-usual-same-
hanging-ratios-at-the.
php?aid=24658.

Braimah:2020:EGS

[Bra20] Lawrence Issah Braimah. Estimate of gillnet selectivity parameters and true relative abundance of Oreochromis niloticus and Sarotherodon galileus, using four indirect methods, with gillnets principally of three different meshes in Thevolta Lake, Ghana. Journal of FisheriesSciences.com, 14 (4):??, ???? 2020. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/

                                        fisheries-aqua/estimate-
                                        of-gillnet-selectivity-
                                        parameters-and-true-relativeabundance-
                                        of-oreochromis-niloticus-
                                        and-sarotherodon-galileususing-
                                        four.pdf.

Biswas:2018:ISP

[BS18] Sutapa Biswas and M. P. Safeena. Isolation of Salmonella phage and its ability to control pathogenic Salmonella in peeled and deveined shrimp. Journal of FisheriesSciences.com, 12 (3):??, ???? 2018. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/isolation-
                                        of-salmonella-phage-and-
                                        its-ability-to-control-
                                        pathogenic-salmonella-
                                        in-peeled-and-deveined-
                                        shrimp.php?aid=23400.

Basiacik:2012:ABG

[BSİU12] Seda Başıaçık, Hasan M. Sarı, Ali İlhan, and M. Ruşen Ustaoğlu. Adigüzel baraj gölü (denizli) biyikli balik (Luciobarbus kottelati Turan, ekmekçi, ilhan & engin, 2008) populasyonunun bazi büyüme Özellikleri. (Turkish) [Some growth features of barbel fish population (Luciobarbus kottelati Turan, Ekmekçi, İlhan & engin, 2008) in Adıgüzel Dam Lake (Denizli)]. Journal of FisheriesSciences.com, 6(1):32–38, ???? 2012.

CODEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/some-
growth-features-of-barbel-
fish-population-luciobarbus-
kottelati-turan-ekmeki-
304lhan--engin-2008-in-
ad305gzel-dam-lake-denizli.
php?aid=684.

Bulut:2012:ECG

[Bul12] Cafer Bulut. The effects on cortisol, glucose and lysozyme activity in different concentration of formaldehyde in rainbow trout. Journal of FisheriesSciences.com, 6(4):??, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
effects-on-cortisol-glucose-
and-lysozyme-activity-
in-different-concentration-
of-formaldehyde-in-rainbow-
trout.php?aid=571.

Cagiltay:2011:AAF

[Çağ11]  Ferhat Çağıltay.  Amino

acid, fatty acid, vitamin and mineral contents of the edible garden snail (Helix aspersa). Journal of FisheriesSciences.com, 5(4): ??, ???? 2011. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/amino-acid-
fatty-acid-vitamin-and-
mineral-contents-of-the-

                                        edible-garden-snail-helix-
                                        aspersa.php?aid=737.

Canak:2011:KYD

[Çan11] Özgür Çanak. Kültürü yapilan denizel türlerin patojeni olan Vibrio harveyi Üzerine bir derleme. (Turkish) [a review on the pathogen of cultured marine species: Vibrio harveyi]. Journal of FisheriesSciences.com, 5(2): ??, ???? 2011. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/a-review-
                                        on-the-pathogen-of-cultured-
                                        marine-species-vibrio-
                                        harveyi.php?aid=770.

Capkin:2011:KGA

[Çap11] Erol Çapkın. Karbosulfanin gökkuşaği alabaliklarin (Oncorhynchus mykiss) eritrosit asetilkolinesteraz (Ache) enzim aktivitesine etkisi. Journal of FisheriesSciences.com, 5(??):??, 2011. CODEN JFOIAQ. ISSN 1307-234X.

Cheng:2018:EXX

[CCL18] Ann-Chang Cheng, Chia-Chun Chi, and Chun-Hung Liu. Effects of X-xylanase incorporated in the diet on the growth performance of red tilapia (Oreochromis sp.). Journal of FisheriesSciences.com, 12(1): ??, ???? 2018. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.

fisheriessciences.com/
fisheries-aqua/effects-
of-xxylanase-incorporated-
in-the-diet-on-the-growth-
performance-of-red-tilapia-
oreochromis-sp.php?aid=
22209.

Celik:2010:CML

[Çel10]  İhsan Çelık.  Character-

ization by microphotography of larval and prejuvenile stage of discus fish (Symphysodon spp.). Journal of FisheriesSciences.com, 4 (1):??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/characterization-
by-microphotography-of-
larval-and-prejuvenile-
stage-of-discus-fish-symphysodon-
spp.php?aid=1079.

Cevik:2008:SVP

[Çev08] Fatma Çevik. Seasonal variations of phytoplankton in the Akyatan and Tuzla lagoons (Adana, Turkey). Journal of FisheriesSciences.com, 2 (1):??, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/seasonal-
variations-of-phytoplankton-
in-the-akyatan-and-tuzla-
lagoons-adana-turkey.php?
aid=1669.

Chiesa:2016:MGC

[CFFM16] Stefania Chiesa, Laura Filonzi, Claudio Ferrari, and Francesco Nonnis Marzano. Molecular genetics confirms the existence of different management units belonging to a single phylogenetic lineage within Italian populations of Alosa fallax (Lacépède, 1803). Journal of FisheriesSciences.com, 10 (4):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/molecular-
genetics-confirms-the-
existence-of-differentmanagement-
units-belonging-to-a-single-
phylogenetic-lineagewithin-
italian-po.php?aid=11247.

Chen:2021:WPB

[CGZ21] Ying Chen, Hui Gao, and Hui Zhao. Winter phytoplankton blooms northwest of the Luzon Island in the South China Sea: a review. Journal of FisheriesSciences.com, 15(2):16– 21, ???? 2021. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/winter-
phytoplankton-blooms-northwest-
of-the-luzon-island-in-
the-south-china-sea-a-
review.php?aid=36175.

Chandramohan:2014:BMS

[Cha14] Bathrachalam Chandramohan. Bird menace: status, threats and mitigation measures in shrimp farms of Cuddalore and Villupuram District of Tamil Nadu, India. Journal of FisheriesSciences.com, 8(4): ??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/bird-
menace-status-threats-
and-mitigation-measures-
in-shrimp-farms-of-cuddalore-
and-villupuram-district-
of-tamil-nadu-india.php?
aid=47.

Chavez:2020:MAE

[Cha20] Ernesto A. Chavez. Market analysis of European summit on aquaculture, fisheries and horticulture. Journal of FisheriesSciences.com, 14 (3):??, ???? 2020. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/market-
analysis-of-european-summit-
on-aquaculture-fisheries-
and-horticulture.pdf.

Chavez:2021:SDA

[Cha21] Ernesto A. Chavez. Save the date: Aquaculture 2021 will be arriving! Journal of FisheriesSciences.com, 15(1):1, ???? 2021. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/save-the-
date-aquaculture-2021-
will-be-arriving.php?aid=
34904.

Cicek:2010:DDI

[ÇKAE10] N. Lerzan Çiçek, Hasan Kalyoncu, Cengiz Akköz, and Ö. Osman Ertan. Dariören deresive isparta çayi (isparta)’nin epilitik algleri ve mevsimsel dağılımları. (Turkish) [seasonal distribution and species of epilithic algae in Darıören and Isparta Stream (Isparta– Turkey)]. Journal of FisheriesSciences.com, 4(1):78– 90, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/seasonal-
                                        distribution-and-species-
                                        of-epilithic-algae-in-
                                        darren-and-isparta-stream-
                                        ispartaturkey.php?aid=
                                        1084.

Cinar:2013:UAG

[ÇKB^{+}13] Šakir Çinar, Ramazan Küçükkara, İsmet Balık, Hıdır Çubuk, Mustafa Ceylan, Kamile Gonca Erol, Vedat Yeğen, and Cafer Bulut. Uluabat (apolyont) gölü’ndeki balik faunasinin tespiti, tür kompozisyonu ve ticari avciliğin türlere göre dağilimi. (Turkish) [Detection of fish fauna, species composition and distribution of commercial fishing according to species in Lake Uluabat]. Journal of FisheriesSciences.com, 7(4):309– 316, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.

fisheriessciences.com/
fisheries-aqua/detection-
of-fish-fauna-species-
composition-and-distribution-
of-commercial-fishing-
according-to-species-in-
lake-uluabat.php?aid=437.

Coban:2009:MAG

[Çob09] Deniz Çoban. Morphometric and allometric growth of cultured common dentex (Dentex dentex). Journal of FisheriesSciences.com, 3 (1):??, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/morphometric-
and-allometric-growth-
of-cultured-common-dentex-
dentex-dentex.php?aid=
1504.

Coban:2011:CUH

[Çob11] Mehmet Zülfü Çoban. Capoeta umbla (Heckel, 1843)’nin hazar gölü (dicle nehri) ve Keban Baraj gölü (firat nehri) populasyonlarinin büyüme Özelliklerinin karşilaştirilmasi. Journal of FisheriesSciences.com, 5 (??):??, 2011. CODEN JFOIAQ. ISSN 1307-234X.

Colakoglu:2014:BMK

[Çol14]  Serhat Çolakoğlu.  Batı

Marmara’da kum midyesi (Chamelea gallina L., 1758) ve kum şırlanı (Donax trunculus L., 1758) populasyonlarının büyüme Özelliklerinin incelenmesi. (Turkish) [Properties growth of populations the striped venus (Chamelea gallina L., 1758) and the wedge clam (Donax trunculus L., 1758) in the West Marmara Sea]. Journal of FisheriesSciences.com, 8 (1):??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/properties-
growth-of-populations-
the-striped-venus-chamelea-
gallina-l-1758-and-the-
wedge-clam-donax-trunculus-
l-1758-in-the-west-marmara-
sea.php?aid=396.

CunhaWanick:2015:SCS

[CPE^{+}15] Rodrigo Cunha Wanick, Ricardo Pilz Vieira, Ricardo Erthal Santelli, Rodolfo Pinheiro Da Rocha Paranhos, and Cristiano Carvalho Coutinho. Screening for cultivable species of marine sponges (Porifera) in an ecologically inspired ex situ system. Journal of FisheriesSciences.com, 9 (4):??, ???? 2015. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/screening-
for-cultivable-species-
of-marine-sponges-porifera-
in-an-ecologically-inspired-
ex-situ-system.php?aid=
8184.

Cirik:2010:BAC

[CŞA10]  Şükran Cirik, Eylem Şen,

and İlknur Ak. Brown algae Cystoseira barbata (Stackhouse) C. Agardh culture and changes in its chemical composition. Journal of FisheriesSciences.com, 4 (4):??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/brown-algae-
cystoseira-barbata-stackhouse-
c-agardh-culture-and-changes-
in-its-chemical-composition.
php?aid=855.

Deori:2017:IFG

[DADL17] Dibya Jyoti Deori, Santoshkumar Abujam, Sanjay Dakua, and Shantikumar Laishram. Investigation on feeding, growth and condition factor of certain ornamental fish from Brahmaputra River at Dibrugarh, Assam, India. Journal of FisheriesSciences.com, 11(3):??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/investigation-
on-feeding-growth-and-
condition-factor-of-certain-
ornamental-fish-from-brahmaputra-
river-at-dibrugarh-assam-
india.php?aid=19868.

Dagli:2008:KYY

[Dağ08]  Mahmut Dağli.  Kilis

yöresinde yakalan iki anormal cyprinid türü. (Turkish) [The two abnormal cyprinid fish caught[ed] in Kilis (Turkey) Region]. Journal of FisheriesSciences.com, 2(4):662–665, ???? 2008. CODEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
two-abnormal-cyprinid-
fish-caughted-in-kilis-
turkey-region.php?aid=
1554.

Dali:2008:FFK

[Dal08] Mahmut Dali. The fish fauna of Kınacık Brook and Afrin Stream (Kilis, Turkey). Journal of FisheriesSciences.com, 2(3):??, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/the-fish-
fauna-of-knack--brookand-
-afrin--stream--kilis--
turkey.php?aid=1600.

Dakua:2018:FHI

[DAM18] Sanjay Dakua, Santoshkumar Abujam, and Islam Mofidul. Feeding habits of Indian flying barb Esomus danrica (Hamilton, 1822) from Upper Assam, India. Journal of FisheriesSciences.com, 12(3): ??, ???? 2018. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/feeding-
habits-of-indian-flying-

barb-esomus-danricahamilton-
1822-from-upper-assam-
india.php?aid=23376.

Dartay:2010:FPG

[Dar10] Müride Dartay. The fishing productivity with gillnets in Pertek Region of Keban Dam Lake. Journal of FisheriesSciences.com, 4 (4):??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
fishing-productivity-with-
gillnets-in-pertek-region-
of-keban-dam-lake.php?
aid=849.

Dogu:2014:DSS

[DAS14] Zafer Doğu, Faruk Aral, and Erdinç Sahinöz. The determination of some spermatological and hematological parameters of shabbout (Barbus grypus, H; 1843) in Atatürk Dam Lake, Şanliurfa. Journal of FisheriesSciences.com, 8(4):265– 277, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
determination-of-some-
spermatological-and-hematological-
parameters-of-shabbout-
barbus-grypus-h-1843-in-
atatrk-dam-lake-anliurfa.
php?aid=2.

DeMello:2016:EGP

[DAS^{+}16] Fernanda De Mello, Carlos Al Oliveira, Danilo Streit, Jr., Emiko K. de Resende, Sheila N. de Oliveira, Darci C. Fornari, Rogério V. Barreto, Jayme A. Povh, and Ricardo P. Ribeiro. Estimation of genetic parameters for body weight and morphometric traits to tambaqui Colossoma macropomum. Journal of FisheriesSciences.com, 10(2):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/estimation-
of-genetic-parameters-
for-body-weight-and-morphometric-
traits-to-tambaqui-colossoma-
macropomum.php?aid=9292.

Diken:2008:LTC

[DB08] Gürkan Diken and Yunus Ömer Boyacı. Light-trapping of caddisflies (Insecta: Trichoptera) from Eğirdir Lake in the Southern Turkey. Journal of FisheriesSciences.com, 2(3):653–661, ???? 2008. CODEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/lighttrapping-
of-caddisflies-insecta-
trichoptera-from-e286irdir-
lake-in-the-southern-turkey.
php?aid=1610.

Darshan:2017:FRB

[DDAD17] Achom Darshan, Rashmi Dutta, Santoshkumar Abujam, and Debangshu Narayan Das. First record of Batasio spilurus Ng from the Siang River of Arunachal Pradesh, Northeastern India (Teleostei: Bagridae). Journal of FisheriesSciences.com, 11(2):??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/first-record-
of-batasio-spilurus-ng-
from-the-siang-river-of-
arunachal-pradesh-northeastern-
india-teleostei-bagridae.
php?aid=18454.

Didinen:2010:FAG

[DED^{+}10] Behire Isil Didinen, Seçil Ekici, Öznur Diler, Seval Bahadir Koca, and Arife Dulluç. The factors affecting growth and survival rate of juveniles in culture of Astacidae family freshwater crayfish. Journal of FisheriesSciences.com, 4 (3):??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
factors-affecting-growth-
and-survival-rate-of-juveniles-
in-culture-of-astacidae-
family-freshwater-crayfish.
php?aid=1025.

Degani:2016:MHE

[Deg16] Gad Degani. A model of hormonal and environmental involvement in growth and sex differentiation in European eel Anguilla anguilla. Journal of FisheriesSciences.com, 10(3):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/a-model-
                                        of-hormonal-and-environmental-
                                        involvement-in-growth-
                                        and-sex-differentiation-
                                        in-european-eel-anguilla-
                                        anguilla.php?aid=10578.

Demir:2008:AFF

[Dem08a]  Orhan Demır.  Aquacul-

ture and fish feed sectors of Turkey general view. Journal of FisheriesSciences.com, 2(4):??, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/aquaculture-
                                        and-fish-feed-sectors-
                                        of-turkey-general-view.
                                        php?aid=1540.

Demrc:2008:SST

[Dem08b] Aydn Demrc. A study on shrimp trawl designs and modifications in İskenderun Bay (Turkey). Journal of FisheriesSciences.com, 2(4): ??, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/a-study-
                                        on-shrimp-trawl-designs-
                                        and-modifications-in-iskenderun-
                                        bay-turkey.php?aid=1552.

Demirci:2009:FSR

[Dem09]  Sevil Demırcı.  Fisheries

structure in River Asi, Hatay–Turkey. Journal of FisheriesSciences.com, 3(2): ??, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/fisheries-
structure-in-river-asi-
hatayturkey.php?aid=1474.

Dernekbasi:2010:UCO

[Der10] Seval Dernekbaşı. Use of canola oil in fish feeds. Journal of FisheriesSciences.com, 4(4):??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/use-
of-canola-oil-in-fish-
feeds.php?aid=838.

Develi:2011:ITO

[Dev11] Elif Eker Develi. işgalci bir taraklı organizma olan Mnemiopsis leidyi’nin karadeniz, hazar denizi ve baltık denizi’ne verdiği zararlar. (Turkish) [Damages to the Black Sea, Caspian Sea and Baltic Sea by the invader comb jelly Mnemiopsis leidyi]. Journal of FisheriesSciences.com, 5(4): ??, ???? 2011. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/damages-
to-the-black-sea-caspian-
sea-and-baltic-sea-by-
the-invader-comb-jelly-

                                        mnemiopsis-leidyi.php?
                                        aid=723.

Diaz-Frances:2017:ACC

[DFDRMGT17] Inmaculada Díaz-Francés, José Díaz-Ruiz, Guillermo Manjón, and Rafael García-Tenorio. ^{210}Po activity concentrations in cooked marine food. Journal of FisheriesSciences.com, 11 (1):??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/210po-activity-
                                        concentrations-in-cooked-
                                        marine-food.php?aid=18057.

Didinen:2013:SCP

[Did13] Behire Işıl Didinen. Screening for candidate probiotic bacteria for the control of Vibrio anguillarum in rainbow trout (Oncorhynchus mykiss, Walbaum). Journal of FisheriesSciences.com, 7(4):??, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/screening-
                                        for-candidate-probiotic-
                                        bacteria-for-the-control-
                                        of-vibrio-anguillarum-
                                        in-rainbow-trout-oncorhynchus-
                                        mykiss-walbaum.php?aid=
                                        433.

Dikel:2011:UYB

[Dik11] Suat Dikel. Su Ürünleri yetiştiriciliğinde boylamanin Önemi. Journal of Fisheries-Sciences.com, 5(??):??, 2011.

CODEN JFOIAQ. ISSN 1307-234X.

Dincaslan:2010:HIN

[Dın10] Yunus Emre Dınçaslan. Histological investigations on the neural complexes in some tunicates (Ascidiacea) live in Izmir Bay from Turkey. Journal of FisheriesSciences.com, 4(1):??, ???? 2010. CODEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/histological-
investigations-on-the-
neural-complexes-in-some-
tunicates-ascidiacea-live-
in-izmir-bay-from-turkey.
php?aid=1082.

deMello:2016:DIF

[dMSD^{+}16] Fernanda de Mello, Juliana Saraiva Garcia, Alexandra Depince, Catherine Labbé, and Danilo Pedro Streit, Jr. DNA isolation from fresh and cryopreserved semen of tambaqui (Colossoma macropomum). Journal of FisheriesSciences.com, 10(2):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/dna-
isolation-from-fresh-and-
cryopreserved-semen-of-
tambaqui-colossoma-macropomum.
php?aid=9182.

Das:2013:NCF

[DNMS13] Saon Das, Sudarshana Nandi, Sandip Majumder, and Surjya Kumar Saikia. New characterization of feeding habits of Puntius sophore (Hamilton, 1822) through morphometry. Journal of FisheriesSciences.com, 7 (3):225–231, ???? 2013. CODEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/new-
characterization-of-feeding-
habits-of-puntius-sophore-
hamilton-1822-through-
morphometry.php?aid=453.

Demirci:2015:FHC

[DO15] Sevil Demirci and Ukran Yalcin Ozdilek. Feeding habits of Capoeta barroisi (Cyprinidae) in the Asi Basin (Orontes) Turkey. Journal of FisheriesSciences.com, 9 (3):??, ???? 2015. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/feeding-
habits-of-capoeta-barroisi-
cyprinidae-in-the-asi-
basinorontes-turkey.php?
aid=6891.

Diler:2013:LCR

[DÖD^{+}13] Öznur Diler, Mehmet Rüstü Özen, Behire şıl Didinen, Nezire Lerzan Çiçek, and Yasar Özvarol. Lipidiozis in captive reared bluefin tuna (Thunnus thynnus L.1758) from the Western Mediterranean Region in Turkey. Journal of FisheriesSciences.com, 7(3):287– 291, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/lipidiozis-
in-captive-reared-bluefin-
tuna-thunnus-thynnus-l1758-
from-the-western-mediterranean-
region-in-turkey.php?aid=
440.

Dogan:2008:PMF

[Dog08]  Gaye Dogan.  Protein

metabolism in fishes. Journal of FisheriesSciences.com, 2(1):??, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/protein-
metabolism-in-fishes.php?
aid=1666.

Dogan:2009:AFW

[Doğ09a]  Gaye Doğan.  Antinutri-

tional factors within plant foodstuffs and their effects on fishes. Journal of FisheriesSciences.com, 3 (4):??, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/antinutritional-
factors-within-plant-foodstuffs-
and-their-effects-on-fishes.
php?aid=1160.

Dogan:2009:SEA

[Doğ09b]  Kadir Doğan.  Socio-

economic analysis of sand smelt fishermen fishing boat owner of Iznik lake (Bursa/ Turkey). Journal of FisheriesSciences.com, 3(1):??, ???? 2009. CODEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/socioeconomic-
                                        analysis-of-sand-smelt-
                                        fishermen-fishing-boat-
                                        owner-of-iznik-lake-bursaturkey.
                                        php?aid=1500.

Dogan:2010:SEA

[Doğ10]  Kadir Doğan.  Socio-

economical analysis of fishery co-operatives and associates in Istanbul. Journal of FisheriesSciences.com, 4 (4):??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/socioeconomical-
                                        analysis-of-fishery-cooperatives-
                                        and-associates-in-istanbul.
                                        php?aid=861.

Doruyol:2012:ECC

[Dor12] Hande Doruyol. The effect of climate change on seafood and their consumption. Journal of FisheriesSciences.com, 6(4):??, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

                                        //www.fisheriessciences.
                                        com/fisheries-aqua/the-
                                        effect-of-climate-change-
                                        on-seafood-and-their-consumption.
                                        php?aid=568.

Duc:2016:WMI

[DTT^{+}16] Pham Minh Duc, Dang Thuy Mai Thy, Ngo Thi Mong Trinh, Tran Ngoc Tuan, and Kishio Hatai. Water molds isolated from eggs and fry of striped catfish (Pangasianodon hypophthalmus) in the Mekong Delta of Viet Nam. Journal of FisheriesSciences.com, 10 (1):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/water-molds-
isolated-from-eggs-and-
fry-of-striped-catfish-
pangasianodon-hypophthalmus-
in-the-mekong-delta-of-
viet-nam.php?aid=8214.

Duman:2012:CPU

[Dum12] Muhsine Duman. The chips produced using surimi powder different rates of chemical composition and sensory quality determination. Journal of FisheriesSciences.com, 6(4):??, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
chips-produced-using-surimi-
powder-different-rates-
of-chemical-composition-
and-sensory-quality-determination.
php?aid=580.

Dunne:2010:DTL

[Dun10] Peter G. Dunne. Determination of the total lipid and the long chain omega-3 polyunsaturated fatty acids, EPA and DHA, in deep-sea fish and shark species from the North-East Atlantic. Journal of FisheriesSciences.com, 4(3):??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/determination-
of-the-total-lipid-and-
the-long-chain-omega3-
polyunsaturated-fatty-
acids-epa-and-dha-in-deepsea-
fish-and-shark-speci.php?
aid=1014.

Dursun:2009:UEP

[Dur09]  Seda Dursun.  The use

of edible protein films in seafood. Journal of FisheriesSciences.com, 3(4):??, ???? 2009. CODEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/the-use-
of-edible-protein-films-
in-seafood.php?aid=1144.

Dursun:2010:ANB

[Dur10]  Seda Dursun.  Applica-

tion of natural biopolymer based nanocomposite films in seafood. Journal of FisheriesSciences.com, 4 (1):??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/application-
of-natural-biopolymer-

based-nanocomposite-films-
in-seafood.php?aid=1088.

Duysak:2008:CDI

[Duy08] Önder Duysak. Cephalopod distribution in İskenderun Bay (Eastern Mediterranean– Turkey). Journal of FisheriesSciences.com, 2(2):??, ???? 2008. CODEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/cephalopod-
distribution-in-iskenderun-
bay-eastern-mediterraneanturkey.
php?aid=1631.

Duyar:2012:TAT

[Duy12] Hünkar Duyar. Tirsi (Alosa tanaica Grimm, 1901)’nin kimyasal kompozisyonu ve buzdolabı koşullarında raf Ömrünün belirlenmesi. (Turkish) [Determination of chemical composition and shelf life of shad (Alosa tanaica Grimm, 1901) in refrigeration conditions]. Journal of FisheriesSciences.com, 6 (1):??, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/determination-
of-chemical-composition-
and-shelf-life-of-shad-
alosa-tanaica-grimm-1901-
in-refrigeration-conditions.
php?aid=692.

Dahiya:2012:EPB

[DVSS12] Tejpal Dahiya, Ravi Kant Verma, Gajender Singh, and Ramchander Sihag. Elimination of pathogenic bacterium, Aeromonas hydrophila by the use of probiotics. Journal of FisheriesSciences.com, 6(3): 209–214, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/elimination-
of-pathogenic-bacterium-
aeromonas-hydrophila-by-
the-use-of-probiotics.
php?aid=604.

Erguden:2012:FFS

[EG12] Sibel Alagöz Ergüden and M. Ziya Lugal Göksu. The fish fauna of the Seyhan Dam Lake (Adana). Journal of FisheriesSciences.com, 6(1): 39–52, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/the-fish-
fauna-of-the-seyhan-dam-
lake-adana.php?aid=683.

Kirankaya:2014:CLW

[eGK14] erife Gülsün Kirankaya. Condition, length–weight and length–length relationships for five fish species from Hirfanli Reservoir, Turkey. Journal of FisheriesSciences.com, 8(3):??, ???? 2014. CODEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/condition-

lengthweight-and-lengthlength-
relationships-for-five-
fish-species-from-hirfanli-
reservoir-turkey.php?aid=
69.

El-Lahamym:2019:EFS

[ELKESA19] A. A. El-Lahamym, K. I.

Khalil, S. A. El-Sherif, and Mahmud A. A. Effect of frozen storage on the chemical composition of sand smelt (Atherina hepsetus) fish burger and finger. Journal of FisheriesSciences.com, 13(1):??, ???? 2019. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/effect-
of-frozen-storage-on-the-
chemical-composition-of-
sandsmelt-atherina-hepsetus-
fish-burger-and-finger.
php?aid=24196.

Ekambaram:2017:UHE

[EN17] Padmini Ekambaram and Meenakshi Narayanan. Upregulation of HSP70 extends cytoprotection to fish brain under xenobiotic stress. Journal of FisheriesSciences.com, 11 (1):??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/upregulation-
of-hsp70-extends-cytoprotection-
to-fish-brain-under-xenobiotic-
stress.php?aid=17325.

El-Regal:2015:STD

[ER15] Mohamed Abu El-Regal. Spatial and temporal distribution of larvae of coral reef fishes in the Red Sea. Journal of FisheriesSciences.com, 9(4):??, ???? 2015. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/spatial-
                                        and-temporal-distribution-
                                        of-larvae-of-coral-reef-
                                        fishes-in-the-red-sea.
                                        php?aid=8187.

Erdem:2010:SHM

[Erd10] Ercan Erdem. A study on horse mackerel (Trachurus mediterraneus, S.) fishing by pair midwater trawl in Black Sea. Journal of FisheriesSciences.com, 4(4): ??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/a-study-
                                        on-horse-mackerel-trachurus-
                                        mediterraneus-s-fishing-
                                        by-pair-midwater-trawl-
                                        in-black-sea.php?aid=845.

Erdogan:2012:RFT

[Erd12] Sevil Erdoğan. The rotifera fauna of Turkish Thrace (Edirne, Tekirdağ, Kırklareli). Journal of FisheriesSciences.com, 6(2): ??, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

                                        //www.fisheriessciences.

com/fisheries-aqua/the-
rotifera-fauna-of-turkish-
thrace-edirne-tekirda-
kirklareli.php?aid=651.

Erguden:2010:GPC

[Erg10] Sibel Alagöz Ergüden. Growth properties of Chondrostoma regium (Heckel, 1843) living in Seyhan Dam Lake (Adana). Journal of FisheriesSciences.com, 4(4):??, ???? 2010. CODEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/growth-
properties-of-chondrostoma-
regium-heckel-1843-living-
in-seyhan-dam-lake-adana.
php?aid=847.

Ergonul:2012:BSV

[Erg12] Mehmet Borga Ergönül.

Balık sağlığı ve immunostimulanların kullanımı. (Turkish) [Fish health and the use of immunostimulants]. Journal of FisheriesSciences.com, 6(3):??, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/fish-
health-and-the-use-of-
immunostimulants.php?aid=
607.

Erkan:2007:DTS

[Erk07] Nuray Erkan. Determination of the total sodium metabisulphide level of shrimps sold in Istanbul. Journal of FisheriesSciences.com, 1(1): 26–33, February 2007. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/determination-
                                        of-the-total-sodium-metabisulphide-
                                        level-of-shrimps-sold-
                                        in-istanbul.php?aid=1759.

Erkan:2008:QAS

[Erk08] Nuray Erkan. Quality assurance systems in food industry. Journal of FisheriesSciences.com, 2(1):??, ???? 2008. CODEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/quality-
                                        assurance-systems-in-food-
                                        industry.php?aid=1676.

Erkan:2013:TTU

[Erk13] Nuray Erkan. Türkiye’de tüketilen su Ürünlerinin omega-3 (-3) yağ asidi profilinin değerlendirilmesi. Journal of FisheriesSciences.com, 7(2):??, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/evaluation-
                                        of-omega-3-9693-fatty-
                                        acid-profile-of-seafood-
                                        products-consumed-in-turkey.
                                        php?aid=518.

Erkan:2014:ASP

[Erk14]  Nuray Erkan.  Alterna-

tive seafood preservation technologies: ionizing radiation and high pressure processing. Journal of FisheriesSciences.com, 8(3): ??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/alternative-
seafood-preservation-technologies-
ionizing-radiation-and-
high-pressure-processing.
php?aid=77.

Eroglu:2009:OST

[Ero09] Mücahit Eroğlu. Otolith size-total length relationship in spiny eel, Mastacembelus mastacembelus (Banks & Solander, 1794) inhabiting in Karakaya Dam Lake (Malatya, Turkey). Journal of FisheriesSciences.com, 3(4):??, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/otolith-
sizetotal-length-relationship-
in--spiny-eel-mastacembelus-
mastacembelus-banks--solander-
1794-inhabiting-inkarakaya-
dam-lake-malatya-turkey.
php?aid=1150.

Esra:2014:HML

[Esr14] Candan Deniz Esra. Heavy metal levels in Cladophora crispata in Melet River (Ordu). Journal of FisheriesSciences.com, 8(2):??, ???? 2014. CODEN JFOIAQ. ISSN 1307-

234X. URL https://www.

                                        fisheriessciences.com/
                                        fisheries-aqua/heavy-metal-
                                        levels-in-cladophora-crispata-
                                        in-melet-river-ordu.php?
                                        aid=254.

Evridiki:2009:GIC

[Evr09] Papastavrou Evridiki. Gender issues in caring for demented relatives. Journal of FisheriesSciences.com, 3 (1):??, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/gender-
                                        issues-in-caring-for-demented-
                                        relatives.php?aid=1516.

Fortunato:2014:PGP

[FA14] Roberta Callicó Fortunato and Esteban Avigliano. Presence of genus Pterois (Oken, 1817) (Scorpaeniformes, Scorpaenidae): Extension of invasive range in Caribbean Sea and first published record for Los Frailes Archipelago. Journal of FisheriesSciences.com, 8(2):??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/presence-
                                        of-genus-pterois-oken-
                                        1817-scorpaeniformes-scorpaenidae-
                                        extension-of-invasive-
                                        range-in-caribbean-sea-
                                        and-first-published.php?
                                        aid=247.

Failler:2015:AGS

[FE15] Pierre Failler and Hachim El Ayoubi. The ACP group of states and the challenge of developing aquaculture. Journal of FisheriesSciences.com, 9(4):??, ???? 2015. CODEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/the-acp-
group-of-states-and-the-
challenge-of-developingaquaculture.
php?aid=8183.

Ferruh:2014:FRW

[Fer14] Aç Ferruh. The first record for watermite Arrenurus berolinensis from Turkey, (Acari: Hydrachnidia). Journal of FisheriesSciences.com, 8(2):??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
first-record-for-watermite-
arrenurus-berolinensis-
from-turkey-acari-hydrachnidia.
php?aid=249.

Fincel:2020:LED

[FGF20] Mark J. Fincel, Cameron W.

Goble, and Eli A. Felts. Lake elevation drives stocking success of Chinook salmon in Lake Oahe, South Dakota, a large midwest reservoir. Journal of FisheriesSciences.com, 14(4): ??, ???? 2020. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.

                                        com/fisheries-aqua/lake-
                                        elevation-drives-stocking-
                                        success-of-chinook-salmon-
                                        in-lake-oahe-south-dakota-
                                        a-large-midwest-reservoir.
                                        pdf.

Filiz:2009:AGR

[Fıl09]  Halit Fılız.  Age and

growth, reproduction and diet of the black goby, (Gobius niger) from Aegean Sea, Turkey. Journal of FisheriesSciences.com, 3(3): ??, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

                                        //www.fisheriessciences.
                                        com/fisheries-aqua/age-
                                        and-growth-reproduction-
                                        and-diet-of-the-black-
                                        goby-gobius-niger-from-
                                        aegean-sea-turkey.php?
                                        aid=1450.

Frosty:2021:ENCa

[Fro21a] Rebeaca Frosty. Editorial note on cold water fisheries. Journal of FisheriesSciences.com, 15(3): ??, ???? 2021. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/editorial-
                                        note-on-cold-water-fisheries.
                                        pdf.

Frosty:2021:ENCb

[Fro21b]  Rebeaca Frosty.  Edito-

rial note on common carp …. Journal of FisheriesSciences.com, 15(3):

??, ????  2021.  CO-

DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/editorial-
note-on-common-carp.pdf.

Frosty:2021:SF

[Fro21c] Rebeaca Frosty. Subjective of fisheries. Journal of FisheriesSciences.com, 15 (3):??, ???? 2021. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/subjective-
of-fisheries.pdf.

Fangue:2021:ETH

[FY21] Nann A. Fangue and Yuzo R.

Yanagitsuru. Effects of temperature on hatching and growth performance of embryos and yolk-sac larvae of a threatened estuarine fish: Longfin smelt. Journal of FisheriesSciences.com, 15(3):??, ???? 2021. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/effects-
of-temperature-on-hatching-
and-growth-performance-
of-embryos-and-yolksac-
larvae-of-a-threatened-
estuarine-fish-longfin-
sme.pdf.

G:2016:FWP

[G16] Caruso G. Fishery wastes and by-products: a resource to be valorised. Journal of FisheriesSciences.com, 10 (1):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/fishery-
wastes-and-byproducts-a-
resource-to-be-valorised.
php?aid=8210.

Ganjoor:2015:SRI

[Gan15] Mohammedsaeed Ganjoor. A short review on infectious viruses in cultural shrimps (Penaeidae family). Journal of FisheriesSciences.com, 9(3):??, ???? 2015. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/a-short-
review-on-infectious-viruses-
in-cultural-shrimps-penaeidae-
family.pdf.

George:2017:HCA

[GAN17] A. D. I. George, O. A. Akinrotimi, and U. K. Nwokoma. Haematological changes in African catfish (Clarias gariepinus) exposed to mixture of atrazine and metolachlor in the laboratory. Journal of FisheriesSciences.com, 11(3): ??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/haematological-
changes-in-african-catfish-
clarias-gariepinusexposed-
to-mixture-of-atrazine-

and-metolachlor-in-the-
laboratory.php?aid=19982.

Garcia:2020:CCT

[Gar20] Breixo Ventoso García. Can canned tuna improve cardiovascular health? Study of the effect of consumption of a functional food developed from Thunnus albacares on the lipid profile. Journal of FisheriesSciences.com, 14(4):??, ???? 2020. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/can-
canned-tuna-improve-cardiovascular-
health-study-of-the-effect-
of-consumption-of-a-functional-
food-developed-fromthunnus-
albaca.pdf.

Godwin:2016:BSH

[GC16] Abu Onisokyetu Monica Godwin and Nwokoma Gift Chinenye. Bioaccumulation of selected heavy metals in water, sediment and blue crab (Callinectes amnicola) from Bodo Creek, Niger Delta, Nigeria. Journal of FisheriesSciences.com, 10(3):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/bioaccumulation-
of-selected-heavy-metals-
in-water-sediment-and-
blue-crab-callinectes-
amnicola-from-bodo-creek-

                                        niger-delta-nigeria.php?
                                        aid=11132.

Gaygusuz:2013:SCD

[GD13] Özcan Gaygusuz and Zeynep Dorak. Species composition and diversity of the zooplankton fauna of Darlik Stream (Istanbul–Turkey) and its tributaries. Journal of FisheriesSciences.com, 7 (4):329–343, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/species-
                                        composition-and-diversity-
                                        of-the-zooplankton-fauna-
                                        of-darlik-stream-304stanbulturkey-
                                        and-its-tributaries.php?
                                        aid=429.

Gogoi:2016:EFD

[GDAD16] Budhin Gogoi, Ramen Das, Santosh Kumar Abujam, and Debangshu Narayan Das. Enumeration of fish from Dulakhojiya Beel (wetland) of Lakhimpur District, Assam, India. Journal of FisheriesSciences.com, 10 (4):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/enumeration-
                                        of-fish-from-dulakhojiya-
                                        beel-wetland-of-lakhimpur-
                                        district-assam-india.php?
                                        aid=11333.

Gelibolu:2009:DHA

[Gel09] Serap Gelibolu. Determination of haemocytes amount and haemocytes type in mature blue crab (Callinectes sapidus, Rathbun, 1896) captured in Akyatan Lagoon (Karataş/Adana– Turkey). Journal of FisheriesSciences.com, 3(3):??, ???? 2009. CODEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/determination-
of-haemocytes-amount-and-
haemocytes-type-in-mature-
blue-crab-callinectes-
sapidus-rathbun-1896-captured-
in-akyatan-la.php?aid=
1464.

Gravenhof:2020:STS

[GMG^{+}20] Dylan A. Gravenhof, Hilary A. Morey, Cameron W. Goble, Mark J. Fincel, and Jacob L. Davis. Short term survival and tag retention of gizzard shad implanted with dummy transmitters. Journal of FisheriesSciences.com, 14(2):??, ???? 2020. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/short-term-
survival-and-tag-retention-
of-gizzard-shad-implanted-
with-dummy-transmitters.
pdf.

Gokcek:2009:AST

[Gök09] C. Kaya Gökçek. An alternative species for traditional carp polyculture in Turkey: Initial growing period. Journal of FisheriesSciences.com, 3(1):??, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

                                        //www.fisheriessciences.
                                        com/fisheries-aqua/an-
                                        alternative-species-for-
                                        traditional-carp-polyculture-
                                        in-turkey-initial-growing-
                                        period.php?aid=1513.

Gokoglu:2014:FRP

[Gök14]  Mehmet Gökoğlu.  The

first record of pink glass shrimp (Pasiphaea multidentata) from in the Gulf of Antalya/Turkey. Journal of FisheriesSciences.com, 8 (1):??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

                                        //www.fisheriessciences.
                                        com/fisheries-aqua/the-
                                        first-record-of-pink-glass-
                                        shrimp-pasiphaea-multidentata-
                                        from-in-the-gulf-of-antalyaturkey.
                                        php?aid=405.

Gonener:2010:DBC

[Gön10] Sedat Gönener. The distribution and biomass of catchable fish caught by commercial bottom trawl in the Black Sea (Sinop-İnceburun Region). Journal of FisheriesSciences.com, 4(3):??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

                                        //www.fisheriessciences.
                                        com/fisheries-aqua/the-
                                        distribution-and-biomass-
                                        of-catchable-fish-caught-
                                        by-commercial-bottom-trawl-

in-the-black-sea-sinop304nceburun-
region.php?aid=1028.

Gonener:2013:KSI

[Gön13]  Sedat Gönener.  Karad-

eniz de (sinop-inceburun) dip trolüile avlanabilir balıkların stok büyüklükleri ve biyokütle değişimleri. Journal of FisheriesSciences.com, 7(2): ??, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
biomass-and-density-changes-
of-catchable-fish-caught-
by-bottom-trawl-in-the-
black-sea-sinop304nceburun-
region.php?aid=498.

Gozu:2010:IPD

[Göz10] Buket Bura Gözü. The impact of pinosylvin on the development of Listeria monocytogenes in the salted rainbow trout (Oncorhynchus mykiss, Walbaum, 1792) stored at different temparatures. Journal of FisheriesSciences.com, 4(4):??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
impact-of-pinosylvin-on-
the-development-of-listeria-
monocytogenes-in-the-salted-
rainbow-trout-oncorhynchus-
mykiss-walbaum-1792.php?
aid=844.

G:2015:GPH

[GTO15] Solomon S. G., Okomoda V. T., and Anyebe P. O. Growth performance and hematological parameters of Clarias gariepinus fed varied levels of colanitida meal. Journal of FisheriesSciences.com, 9(2):??, ???? 2015. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/growth-
                                        performance-and-hematological-
                                        parameters-of-clarias-
                                        gariepinus-fed-varied-
                                        levels-of-cola-nitida-
                                        meal.php?aid=5823.

Gaffaroglu:2014:KPO

[GUA14] Muhammet Gaffaroğlu, Sevgi Unal, and Muradiye Karasu Ayata. Karyotype properties of Oxynoemacheilus angorae (Steindachner,1897) (Teleostei, Nemacheilidae) from Anatolia. Journal of FisheriesSciences.com, 8 (4):??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/karyotype-
                                        properties-of-oxynoemacheilus-
                                        angorae-steindachner1897-
                                        teleostei-nemacheilidae-
                                        from-anatolia.php?aid=
                                        8748.

Gulsun:2007:FAC

[Gül07]  Özyurt Gülsün.  Fatty

acid composition and lipid oxidation during refrigerated storage (4 [BAD-UTF-8]C) of thinlip grey mullet (Liza ramada). Journal of FisheriesSciences.com, 1(4):??, ???? 2007. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/fatty-acid-
composition-and-lipid-
oxidation-during-refrigerated-
storage-4-c-of-thinlip-
grey-mullet-liza-ramada.
php?aid=1726.

Gulnaz:2008:LWR

[Gül08]  Özcan Gülnaz.  Length–

weight relationships of five freshwater fish species from the Hatay Province, Turkey. Journal of FisheriesSciences.com, 2(1):??, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/lengthweight-
relationships-of-five-
freshwater-fish-species-
from-the-hatay-province-
turkey.php?aid=1662.

Gulahn:2009:LWR

[Gül09] Anl Gülahn. Length–weight relationships in blue crab, Callinectes sapidus (Rathbun, 1896) in Køyçegiz Dalyan Lagoon Area–Turkey. Journal of FisheriesSciences.com, 3(1):??, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/lengthweight-

                                        relationships-in-blue-
                                        crab-callinectes-sapidus-
                                        rathbun-1896-in-kycegiz-
                                        dalyan-lagoon-areaturkey.
                                        php?aid=1507.

Gundogdu:2008:AHM

[Gün08a] Ayşe Gündoğdu. The accumulation of the heavy metals (copper and zinc) in the tissues of rainbow trout (Onchorhyncus mykiss, Walbaum, 1792). Journal of FisheriesSciences.com, 2 (1):??, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

                                        //www.fisheriessciences.
                                        com/fisheries-aqua/the-
                                        accumulation-of-the-heavy-
                                        metals-copper-and-zinc-
                                        in-the-tissues-of-rainbow-
                                        trout-onchorhyncus-mykiss-
                                        walbaum-1792.php?aid=1664.

Gundogdu:2008:ATZ

[Gün08b]  Ayşe Gündoğdu.  Acute

toxicity of zinc and copper for rainbow trout (Onchorhyncus mykiss). Journal of FisheriesSciences.com, 2(4):??, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/acute-toxicity-
                                        of-zinc-and-copper-for-
                                        rainbow-trout-onchorhyncus-
                                        mykiss.php?aid=1537.

Guner:2009:DLC

[Gün09] Utku Güner. Determination of lambda cyhalotrin (Tekvando 5ec) 96 hour lethal concentration 50 at Gambusia affinis (Baird & Girard, 1853). Journal of FisheriesSciences.com, 3(3): ??, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/determination-
of-lambda-cyhalotrin-tekvando-
5ec-96-hour-lethal-concentration-
50-at-gambusia-affinis-
baird--girard-1853.php?
aid=1457.

Guner:2010:BCM

[Gün10] Utku Güner. Bioaccumulation of cobalt in mosquitofish (Gambusia affinis Baird & Girars, 1853) at different flow rates and concentrations. Journal of FisheriesSciences.com, 4(1): ??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/bioaccumulation-
of-cobalt-in-mosquitofish-
gambusia-affinis-bairdgirars-
1853-at-different-flow-
rates-and-concentrations.
php?aid=1092.

Guner:2011:FSH

[Gün11] Utku Güner. Flow sitometrinin hidrobiyolojide kullanımı. Journal of Fisheries-Sciences.com, 5(??):??, 2011. CODEN JFOIAQ. ISSN 1307-234X.

Guner:2012:UFC

[Gün12] Utku Güner. Using flow cytometry in hydrobiology. Journal of FisheriesSciences.com, 6(1):??, ???? 2012. CODEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/using-flow-
                                        cytometry-in-hydrobiology.
                                        php?aid=691.

Gundodu:2014:UCM

[Gün14] Sedat Gündodu. The usage of common multiple comparison tests (post-hoc) in fisheries sciences. Journal of FisheriesSciences.com, 8 (4):??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

                                        //www.fisheriessciences.
                                        com/fisheries-aqua/the-
                                        usage-of-common-multiple-
                                        comparison-tests-posthoc-
                                        in-fisheries-sciences.
                                        php?aid=48.

Gurbuz:2011:ARF

[Gür11] Özlem Ablak Gürbüz. Age and reproduction features of tench (Tinca tinca (L., 1758)) from Hirfanli Dam Lake, Kirşehir, Turkey. Journal of FisheriesSciences.com, 5(2):??, ???? 2011. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/age-and-
                                        reproduction-features-

of-tench-tinca-tinca-l-
1758-from-hirfanli-dam-
lake-kir350ehir-turkey.
php?aid=765.

Gurkan:2012:BPS

[Gür12]  Ule Gürkan.  Biomet-

ric properties of the sand smelt Atherina boyeri Risso, 1810 from the Izmir Bay (Aegean Sea). Journal of FisheriesSciences.com, 6(1): ??, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/biometric-
properties-of-the-sand-
smelt-atherina-boyeri-
risso-1810-from-the-izmir-
bay-aegean-sea.php?aid=
689.

Harlioglu:2013:IFM

[Har13] Ayşe Gül Harlıoğlu. Impact of fish meal replacement with full-fat soya on the muscle and liver fatty acid composition in rainbow trout. Journal of FisheriesSciences.com, 7(2):??, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/impact-
of-fish-meal-replacement-
with-fullfat-soya-on-the-
muscle-and-liver-fatty-
acid-composition-in-rainbow-
trout.php?aid=487.

Hartl:2015:FEC

[Har15] Mark Gj Hartl. Fish ecotoxicology in a changing world:

Do we need new biomarker endpoints in light of climate change? Journal of FisheriesSciences.com, 9 (4):??, ???? 2015. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/fish-
ecotoxicology-in-a-changing-
world-do-we-need-new-biomarker-
endpoints-in-light-of-
climate-change.php?aid=
8191.

Hashem:2014:SAB

[Has14] Seyed Ahmadreza Hashem. Stock assessment of bartail flathead (Platycephalus indicus Linnaeus, 1758) in Northwest of Persian Gulf, Iran. Journal of FisheriesSciences.com, 8(2): ??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/stock-assessment-
of-bartail-flathead-platycephalus-
indicus-linnaeus-1758-
in-northwest-of-persian-
gulf-iran.php?aid=266.

Hassan:2015:TIH

[Has15] Marina Hassan. Toxicity induced historical changes in gills of Oreochromis sp. by exposed to clove oil. Journal of FisheriesSciences.com, 9(4):??, ???? 2015. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/toxicity-

induced-historical-changes-
in-gills-of-oreochromis-
sp-by-exposed-to-clove-
oil.php?aid=8188.

Harmantepe:2007:ETD

[HB07] Fatma Burcu Harmantepe and Sevket Büyükhatipoğlu. Effect of two different feeds on growth of rainbow trout and feed cost. Journal of FisheriesSciences.com, 1(4): 168–175, ???? 2007. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/effect-
of-two-different-feeds-
on-growth-of-rainbow-trout-
and-feed-cost.php?aid=
1723.

HossainABM:2018:SCA

[HBP^{+}18] Arman Hossain A.B.M, Shankor Bisshas, Mehedi Hasan Pramanik Md, Monjurul Hasan Md, Istiaque Haidar Md, Aovijite Bosu, Ashikur Rahman Md, and Anisur Rahman Md. Supply chain analysis of hilsa (Tenualosa ilisha) egg in Bangladesh. Journal of FisheriesSciences.com, 12 (4):??, ???? 2018. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/supply-
chain-analysis-of-hilsa-
tenualosa-ilisha-egg-in-
bangladesh.php?aid=23689.

Hossain:2019:SCA

[HBPH19] Abm. Arman Hossain, Shankor Bisshas, Md. Mehedi Hasan Pramanik, and Md. Monjurul Hasan. Supply chain analysis of salted Tenualosa ilisha (Nona ilish) in Bangladesh. Journal of FisheriesSciences.com, 13 (4):??, ???? 2019. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/supply-
                                        chain-analysis-of-salted-
                                        tenualosa-ilisha-nona-
                                        ilish-inbangladesh.php?
                                        aid=25407.

Hedayati:2016:ESC

[HDS^{+}16] Aliakbar Hedayati, Fatemeh Darabitabar, Fardin Shaluei, Zahra Ghaffari, and Alireza Kashiri. Evaluation of silver carp (Hypophthalmichthys molitrix Valenciennes, 1844) mortality rate in the face of lethal concentrations of commercial gasoline. Journal of FisheriesSciences.com, 10(4):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/evaluation-
                                        of-silver-carp-hypophthalmichthys-
                                        molitrixvalenciennes-1844-
                                        mortality-rate-in-the-
                                        face-of-lethalconcentrations-
                                        of-comme.php?aid=11238.

Hammadi:2016:DRS

[HSAE16] Naeem Shanad Hammadi, Salman Dawood Salman, and Salih Abdul-Qadir Al-Essa. Diversity of rotifera in the shatt Al-Arab region, south of Iraq. Journal of FisheriesSciences.com, 10(2):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/diversity-
of-rotifera-in-the-shatt-
alarab-region-south-of-
iraq.php?aid=8566.

Ibemenuga:2019:BSH

[IEN^{+}19] Keziah N. Ibemenuga, Faith A.

Ezike, Moses C. Nwosu, Lucy C. Anyaegbunam, Ebelechukwu I. Okoye, and Joseph Effiong Eyo. Bioaccumulation of some heavy metals in some organs of three selected fish of commercial importance from Niger River, Onitsha Shelf, Anambra State, Nigeria. Journal of FisheriesSciences.com, 13(3):??, ???? 2019. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/bioaccumulation-
of-some-heavy-metals-in-
some-organs-of-three-selected-
fish-of-commercial-importance-
from-niger-river-onitsha-
shelf.php?aid=24499.

Inanli:2010:CCM

[İÖÇK10] Ayşe Gürel İnanlı, Emine Özpolat, Özlem Emir Çoban, and Nermin Karaton. Chemical composition of marinated anchovy (Engraulis encrasicolus L., 1758) and sensory evaluation in different sauce. Journal of FisheriesSciences.com, 4(4):??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/chemical-
                                        composition-of-marinated-
                                        anchovy-engraulis-encrasicolus-
                                        l-1758-and-sensory-evaluation-
                                        in-different-sauce.php?
                                        aid=840.

Ipek:2012:GCVb

[İpe12] Necla İpek. Görgüşan çayı ve geban deresi (elazığtürkiye) zooplanktonu. Journal of FisheriesSciences.com, 6(??):??, 2012. CODEN JFOIAQ. ISSN 1307-234X.

Ipek:2012:GCVa

[İS12] Necla İpek and Serap Saler. Görgüşan çayi ve Geban deresi (Elaziğ-Türkiye) zooplanktonu. (Turkish) [Zooplankton of Gorgusan stream and Geban stream (Elazıg– Turkey)]. Journal of FisheriesSciences.com, 6(2):155– 163, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/zooplankton-
                                        of-gorgusan-stream-and-
                                        geban-stream-elazgturkey.
                                        php?aid=649.

Ipek:2013:ZCS

[İS13] Necla İpek and Serap Saler.

Zooplankton community structure of Ohi Stream (Elazig– Turkey). Journal of FisheriesSciences.com, 7(1):83– 88, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/zooplankton-
community-structure-of-
ohi-stream-elazigturkey.
php?aid=539.

Isinibilir:2009:SMC

[İşi09] Melek İşinibilir. Summer mesozooplankton communities in the Turkish coastal waters of North Aegean Sea. Journal of FisheriesSciences.com, 3(3):237– 242, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/summer-
mesozooplankton-communities-
in-the-turkish-coastal-
waters-of-north-aegean-
sea.php?aid=1452.

Ilhan:2014:GCB

[İSU14] Ali İlhan, Hasan M. Sarı, and M. Ruşen Ustaoğlu. Gönen çayi (Balikesir) balik faunasi. (Turkish) [fish fauna of gönen stream (balıkesir)]. Journal of FisheriesSciences.com, 8(3):194–198, ???? 2014. CODEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/fish-
fauna-of-gnen-stream-bal305kesir.
php?aid=66.

Jale:2014:MVA

[Jal14] Korun Jale. Marine vibrios associated with diseased sea bass (Dicentrarchus labrax) in Turkey. Journal of FisheriesSciences.com, 8(1):??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/marine-
vibrios-associated-with-
diseased-sea-bass-dicentrarchus-
labrax-in-turkey.php?aid=
409.

Jawad:2012:RBF

[Jaw12a] Laith A. Jawad. Relationship between fish length and otolith dimensions in the carangid fish (Carangoides coeruleopinnatus (Rüppell, 1830)) collected from the Sea of Oman. Journal of FisheriesSciences.com, 6 (3):??, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/relationship-
between-fish-length-and-
otolith-dimensions-in-
the-carangid-fish-carangoides-
coeruleopinnatus-rppell-
1830-collected-fr.php?
aid=606.

Jawad:2012:SOM

[Jaw12b] Laith A. Jawad. Study on the otolith mass asymmetry in Lutjanus bengalensis (family: Lutjanidae) collected from Muscat City on the Sea of Oman. Journal of FisheriesSciences.com, 6 (1):??, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/study-on-
the-otolith-mass-asymmetry-
in-lutjanus-bengalensis-
family-lutjanidae-collected-
from-muscat-city-on-the-
sea-of-oman.php?aid=677.

Ji:2016:AES

[JK16] Keunho Ji and Young Tae Kim. Algicidal effect of Serratia sp. PDGS120915 against harmful Dinophyceae algae. Journal of FisheriesSciences.com, 10(2): ??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/algicidal-
effect-of-serratia-sp-
pdgs120915-against-harmful-
dinophyceae-algae.php?
aid=9295.

Jones:2020:ENJ

[Jon20]  April Jones.  Editorial

note: Journal of Fisheries-Sciences.com. Journal of FisheriesSciences.com, 14 (4):??, ???? 2020. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/editorial-

                                        note-journal-of-fisheriessciencescom.
                                        pdf.

Janko:2015:FWP

[JZ15] Assefa Mitike Janko and Lemma Zemedu. Fisherman’s willingness to pay for fisheries management: the case of Lake Zeway, Ethiopia. Journal of FisheriesSciences.com, 9(4):??, ???? 2015. CODEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/fisher-
                                        mans-willingness-to-pay-
                                        for-fisheries-management-
                                        the-case-of-lake-zeway-
                                        ethiopia.php?aid=8186.

Kareem:2017:EDF

[KAA^{+}17] O. K. Kareem, E. K. Ajani, M. A. Akintunde, A. N. Olanrewaju, and O. B. Oduntan. Effect of different fertilization and egg deadhesion methods on hatching and survival of Clarias gariepinus (Burchell 1822) fry. Journal of FisheriesSciences.com, 11 (1):??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/effect-
                                        of-different-fertilization-
                                        and-egg-deadhesion-methods-
                                        on-hatching-and-survival-
                                        of-clarias-gariepinus-
                                        burchell-1822-fry.php?
                                        aid=17326.

Kaba:2009:STP

[Kab09] Nilgün Kaba. Surimi technology and the production of surimi from fat and dark muscle fish. Journal of FisheriesSciences.com, 3(4): ??, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/surimi-
technology-and-the-production-
of-surimi-from-fat--and-
dark-muscle-fish.php?aid=
1169.

Kumar:2016:FFS

[KAGD16] Ram Kumar, Santosh Kumar Abujam, Arijit Ganguly, and Debangshu Narayan Das. Fish and fisheries of Sinkin Tributary with emphasis on the people’s socio-economic dependence in Dibang River Basin of Arunachal Pradesh, India. Journal of FisheriesSciences.com, 10(2): ??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/fish-
and-fisheries-of-sinkin-
tributary-with-emphasis-
on-the-peoples-socioeconomic-
dependence-in-dibang-river-
basin-of-arunachal-pr.
php?aid=9104.

Kenge:2021:RSC

[KANO21] B. N. Kenge, M. S. Abdullahi, and K. A. Njoku-Onu. A review of sustainable culture and conservation of indigenous scaly fish — case study of heterotis (Heterotis niloticus

— Cuvier 1829). Journal of FisheriesSciences.com, 15 (2):22–25, ???? 2021. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/a-review-
of-sustainable-culture-
and-conservation-of-indigenous-
scaly-fish-case-study-
of-heterotis-heterotis-
niloticuscuvier-1829.php?
aid=36177.

Karacuha:2007:AFU

[Kar07] Ali Karaçuha. Algae flora in the upper infralittoral zone of Sinop and Ayancik coastline. Journal of FisheriesSciences.com, 1(1): 1–12, February 2007. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/algae-flora-
in-the-upper-infralittoral-
zone-of-sinop-and-ayancik-
coastline.php?aid=1772.

Karabulut:2010:AST

[Kar10] Huriye Ariman Karabulut. Adaptation of stream trout (Salmo trutta) belonging to the North Firat Population to the culture condition. Journal of FisheriesSciences.com, 4 (4):??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.

fisheriessciences.com/
fisheries-aqua/adaptation-
of-stream-trout-salmo-
trutta-belonging-to-the-
north-firat-population-
to-the-culture-condition.
php?aid=868.

Karacuha:2011:EFA

[Kar11] Ali Karaçuha. Effects of feeding with aquarium fish food and some live foods on growth in fry guppies (Poecilia reticulata Peters, 1859). Journal of FisheriesSciences.com, 5(1): ??, ???? 2011. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/effects-
of-feeding-with-aquarium-
fish-food-and-some-live-
foods-on-growth-in-fry-
guppies-poecilia-reticulata-
peters-1859.php?aid=822.

Kaya:2008:FAC

[Kay08] Yalçn Kaya. Fatty acids composition of anchovy (Engraulis encrasicolus L. 1758) oil produced in Sinop– Turkey. Journal of FisheriesSciences.com, 2(4):??, ???? 2008. CODEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/fatty-acids-
composition-of-anchovy-
engraulis-encrasicolus-
l-1758-oil-produced-in-
sinopturkey.php?aid=1544.

Kayhan:2010:BIH

[Kay10] Figen Esin Kayhan. A biological importance of heat shock proteins (HSP) in aquatic organisms and thermotolerance. Journal of FisheriesSciences.com, 4(3): ??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/a-biological-
                                        importance-of-heat-shock-
                                        proteins-hsp-in-aquatic-
                                        organisms-and-thermotolerance.
                                        php?aid=1020.

Kayis:2013:PDO

[Kay13]  Şevki Kayış.  Parasites

on different ornamental fish species in Turkey. Journal of FisheriesSciences.com, 7 (2):??, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/parasites-
                                        on-different-ornamental-
                                        fish-species-in-turkey.
                                        php?aid=492.

Kaya:2014:IST

[Kay14a] Nan Kaya. Investigation of the serum total antioxidant, oxidant and sialic acid levels of Cyprinus carpio (L. 1758) treated with Tebuconazole (fungicide). Journal of FisheriesSciences.com, 8(3):??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/

fisheries-aqua/investigation-
of-the-serum-total-antioxidant-
oxidant-and-sialic-acid-
levels-of-cyprinus-carpio-
l-1758-treated-with-tebuconazole-
fungicide.php?aid=71.

Kayhan:2014:EHM

[Kay14b] Figen Esin Kayhan. Effects of heavy metal induction on oxidative stress parameters in zebrafish (Danio rerio). Journal of FisheriesSciences.com, 8(3):??, ???? 2014. CODEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/effects-
of-heavy-metal-induction-
on-oxidative-stress-parameters-
in-zebrafish-danio-rerio.
php?aid=67.

Kientz:2017:CSR

[KBD17] Jeremy L. Kientz, Michael E.

Barnes, and Dan J. Durben. Concentration of stocked rainbow trout catch and harvest by a small number of recreational anglers. Journal of FisheriesSciences.com, 11(1):??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/concentration-
of-stocked-rainbow-trout-
catch-and-harvest-by-a-
small-number-of-recreational-
anglers.php?aid=18198.

Kientz:2020:EEF

[KDCS20] Jeremy L. Kientz, Jacob L. Davis, Steven R. Chipps, and Gregory Simpson. Effect of environmental factors on the movement of rainbow trout in the Deerfield Reservoir system. Journal of FisheriesSciences.com, 14 (1):??, ???? 2020. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/effect-
                                        of-environmental-factors-
                                        on-the-movement-of-rainbow-
                                        trout-in-the-deerfield-
                                        reservoir-system.php?aid=
                                        26132.

Koca:2011:UYP

[KDED11] Seval Bahadır Koca, Behire Işıl Didinen, Seçil Ekici, and Arife Dulluç. Su Ürünleri yetiştiriciliğinde probiyotik uygulamaları. (Turkish) [Applications of probiotic in aquaculture]. Journal of FisheriesSciences.com, 5(4):326–335, ???? 2011. CODEN JFOIAQ. ISSN 1307-234X. URL https:

                                        //www.fisheriessciences.
                                        com/fisheries-aqua/applications-
                                        of-probiotic-in-aquaculture.
                                        php?aid=709.

Kelestemur:2010:ELO

[Kel10] Gülüzar Tuna Keleştemur. Effect of low oxygen level on some blood parameter levels of fry rainbow trouts (Oncorhynchus mykiss, W.

1792). Journal of FisheriesSciences.com, 4(4):??, ???? 2010. CODEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/effect-
of-low-oxygen-level-on-
some-blood-parameter-levels-
of-fry-rainbow-trouts-
oncorhynchus-mykiss-w-
1792.php?aid=864.

Keskin:2012:TAK

[Kes12a] Emre Keskin. Türkiye’nin akdeniz kıyısındaki mavi yengeç (Callinectes sapidus) populasyonları arasındaki genetik farklılığın coi gen dizileri kullanılarak değerlendirilmesi. Journal of FisheriesSciences.com, 6(??):??, 2012. CODEN JFOIAQ. ISSN 1307-234X.

Keskn:2012:DGV

[Kes12b] Emre Keskn. Determination of genetic variation among blue crab (Callinectes sapidus) populations along Mediterranean Coast of Turkey using COI sequences. Journal of FisheriesSciences.com, 6(2):??, ???? 2012. CODEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/determination-
of-genetic-variation-among-
blue-crab-callinectes-
sapidus-popuations-along-
mediterranean-coast-of-

                                        turkey-using-coi-sequences.
                                        php?aid=652.

Kenoye:2016:EEC

[KG16] Ukwe Isaac Oyekutor Kenoye and Abu Onisokyetu Monica Godwin. Evaluation of efficacy and cost effectiveness of ovulin and ovaprim hormones for spawning of African catfish (Clarias gariepinus). Journal of FisheriesSciences.com, 10 (4):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/evaluation-
                                        of-efficacy-and-cost-effectiveness-
                                        of-ovulin-and-ovaprim-
                                        hormones-for-spawning-
                                        of-african-catfish-clarias-
                                        gariepinus.php?aid=16131.

Khiari:2011:EGM

[Khi11] Zied Khiari. The extraction of gelatine from mackerel (Scomber scombrus) heads with the use of different organic acids. Journal of FisheriesSciences.com, 5(1): ??, ???? 2011. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

                                        //www.fisheriessciences.
                                        com/fisheries-aqua/the-
                                        extraction-of-gelatine-
                                        from-mackerel-scomber-
                                        scombrusheads-with-the-
                                        use-of-different-organic-
                                        acids.php?aid=810.

Karimpour:2013:PSF

[KHK^{+}13] Mohammad Karimpour, Muzaffer Mustafa Harlioğlu, Ali Asghar Khanipour, Shahram Abdolmalaki, and Önder Aksu. Present status of fisheries in Iran. Journal of FisheriesSciences.com, 7(2): 161–177, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/present-
status-of-fisheries-in-
iran.php?aid=507.

Kilinc:2011:NAF

[Kil11] Berna Kilinç. Nanotechnology applications in foodspackagings and importance for seafoods. Journal of FisheriesSciences.com, 5(4): ??, ???? 2011. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/nanotechnology-
applications-in-foodspackagings-
and-importance-for-seafoods.
php?aid=710.

Kehayias:2016:ETP

[KKP16] George Kehayias, Pavlos Koutra, and Aristotelis Pournos. Evaluation of the technical performance and operational effectiveness of three LED lamp buoys for purse seine fishery. Journal of FisheriesSciences.com, 10(4):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.

fisheriessciences.com/
fisheries-aqua/evaluation-
of-the-technical-performance-
and-operational-effectiveness-
of-three-led-lamp-buoys-
for-purse-seine-fishery.
php?aid=16493.

Klempt:2013:NIE

[Kle13]  Martin Klempt.  353-

nonylphenol induces expression of the T-box6 gene in zebrafish embryos — linking transcriptional information with deformities. Journal of FisheriesSciences.com, 7(1):??, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/353nonylphenol-
induces-expression-of-
the-tbox6-gene-in-zebrafish-
embryos--linking-transcriptional-
information-with-deformities.
php?aid=551.

Kebede:2016:ENA

[KLPK16] Gebremeskel Eshetu Kebede, Chun Woo Lee, Subong Park, and Mun Kwan Kim. Experimental and numerical analysis of trolling line for hairtail fishing. Journal of FisheriesSciences.com, 10(4):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/experimental-
and-numerical-analysis-
of-trolling-line-for-hairtail-
fishing.php?aid=11332.

Kayhan:2009:BAM

[KMK09] Figen Esin Kayhan, Mehmet Nezih Muşlu, and Nazan Deniz Koç. Bazı ağır metallerin sucul organizmalar Üzerinde yarttığı stres ve biyolojik yanıtlar. (Turkish) [stress and biological responses seen in aquatic organisms due to some trace elements]. Journal of FisheriesSciences.com, 3(2):153– 162, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/stress-
and-biological-responses-
seen-in-aquatic-organisms-
due-to-some-trace-elements.
php?aid=1476.

Koc:2010:SUZ

[Koç10]  Nazan Deniz Koç.  A

study on ultrastructure of Zona radiata during oocyte development of zebrafish (Danio rerio). Journal of FisheriesSciences.com, 4(2): ??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/a-study-
on-ultrastructure-of-zona-
radiata-during-oocyte-
development-of-zebrafish-
danio-rerio.php?aid=1040.

Kaba:2012:DZB

[KÖÇ12] Nilgün Kaba, Özgül Özer, and Bengünur Çorapci. Dumanlanmiş zargana (Belone belone euxini Günther, 1866) köftelerinin bazi kalite parametrelerinin belirlenmesi. (Turkish) [The determination of some quality parameters of smoked gar fish meat balls]. Journal of FisheriesSciences.com, 6 (4):357–367, ???? 2012. CODEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
determination-of-some-
quality-parameters-of-
smoked-gar-fish-meat-balls.
php?aid=566.

Koprucu:2012:EDP

[Köp12] Kenan Köprücü. Effects of dietary protein and lipid levels on growth, feed utilization and body composition of juvenile grass carp (Ctenopharyngodon idella). Journal of FisheriesSciences.com, 6(3):??, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/effects-
of-dietary-protein-and-
lipid-levels-on-growth-
feed-utilization-and-body-
composition-of-juvenile-
grass-carp-ctenopharyngodon-
idella.php?aid=599.

Korun:2008:MVA

[Kor08] Jale Korun. Marine vibrios associated with diseased sea bass (Dicentrarchus labrax) in Turkey. Journal of FisheriesSciences.com, 2(??):??, 2008. CODEN JFOIAQ. ISSN 1307-234X.

Kumar:2016:EFA

[KST16] B. Naga Pavan Kumar, Mahaboobi S., and Akhilesh T. Effect of feed additives on growth performance of fish. Journal of FisheriesSciences.com, 10 (3):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/effect-
of-feed-additives-on-growth-
performance-of-fish.php?
aid=11220.

K:2016:PDC

[KTT16] Nanthini Devi K., Ajith Kumar T. T., and Balasubramanian T. Pigment deficiency correction in captive clown fish, Amphiprion ocellaris using different carotenoid sources. Journal of FisheriesSciences.com, 10(1): ??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/pigment-
deficiency-correction-
in-captive-clown-fish-
amphiprion-ocellaris-using-
different-carotenoid-sources.
php?aid=8209.

Kubilay:2009:SAB

[Kub09] Ayegül Kubilay. A study on aerobic bacterial flora during incubation of rainbow trout (Oncorhynchus mykiss, Walbaum 1792) eggs in hatchery. Journal of FisheriesSciences.com, 3(1): ??, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/a-study-
on-aerobic-bacterial-flora-
during-incubation-of-rainbow-
trout-oncorhynchus-mykiss-
walbaum-1792-eggs-in-hatchery.
php?aid=1515.

Kucukgul:2008:ASR

[Küç08] Azime Küçükgül. Acute stress response in common carp (Cyprinus carpio Linnaeus, 1758) of some stressing factors. Journal of FisheriesSciences.com, 2(3): ??, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/acute-stress-
response-in-common-carp-
cyprinus-carpio-linnaeus-
1758-of-some-stressing-
factors.php?aid=1602.

Kusatan:2013:DHS

[Kuş13] Zafer Kuşatan. Denizatlarinin (Hippocampus spp.) biyolojisi, populasyon durumlari ve yetiştiricik alanindaki yeri. (Turkish) [Seahorses (Hippocampus spp.); biology, population status and their place in cultivation]. Journal of FisheriesSciences.com, 7(1):??, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/seahorses-
-hippocampus-spp-biology-
population-status-and-
their-place-in-cultivation.
php?aid=531.

Kutlu:2010:ETG

[Kut10]  Banu Kutlu.  The ef-

fect of temperature on the growth of Hantzchia amphioxys (Ehrenberg) Grunow isolated from the Homa Lagoon, Izmir. Journal of FisheriesSciences.com, 4(4): ??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
effect-of-temperature-
on-the-growth-of-hantzchia-
amphioxys-ehrenberg-grunow-
isolated-from-the-homa-
lagoon-izmir.php?aid=839.

Karaulova:2017:CSM

[KY17] E. P. Karaulova and E. V.

Yakush. The comparative study of myofibrillar proteins of skeletal muscles of some deep-sea fish species. Journal of FisheriesSciences.com, 11(2):??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
comparative-study-of-myofibrillar-
proteins-of-skeletal-muscles-

                                        of-some-deepsea-fish-species.
                                        php?aid=18343.

Lawal:2013:GEP

[Law13] Muyideen Owonire Lawal. Growth and economic performance of Clarias gariepinus fed different sources of calcium and phosphorus diets. Journal of FisheriesSciences.com, 7(2): ??, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/growth-
                                        and-economic-performance-
                                        of-clarias-gariepinus-
                                        fed-different-sources-
                                        of-calcium-and-phosphorus-
                                        diets.php?aid=517.

Lawal:2014:DER

[Law14] Muyideen Owonire Lawal. Dietary effects of ripe and unripe banana peels on the growth and economy of production of juvenile catfish (Clarias gariepinus Burchell, 1822). Journal of FisheriesSciences.com, 8 (3):??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/dietary-
                                        effects-of-ripe-and-unripe-
                                        banana-peels-on-the-growth-
                                        and-economy-of-production-
                                        of-juvenile-catfish-clarias-
                                        gariepinus-burchell-1822.
                                        php?aid=73.

Lemos:2012:PEB

[Lem12] Jefferson Lemos. Parasitological evaluation and body indices of Osteoglossum bicirrhosum (Vandelli, 1829) traded in a fair of Manaus, Amazonas, Brazil. Journal of FisheriesSciences.com, 6 (3):??, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/parasitological-
evaluation-and-body-indices-
of-osteoglossum-bicirrhosum-
vandelli-1829-traded-in-
a-fair-of-manaus-amazonas-
brazil.php?aid=597.

Levent:2011:CEB

[Lev11] Bat Levent. The changed ecosystem of the Black Sea and its impact on anchovy fisheries. Journal of FisheriesSciences.com, 5 (4):??, ???? 2011. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
changed-ecosystem-of-the-
black-sea-and-its-impact-
on-anchovy-fisheries.php?
aid=733.

Lhan:2014:LWR

[Lha14]  Ali Lhan.  The length–

weight relationship of bitterling, Rhodeus amarus (Bloch, 1782) in freshwaters of Turkey. Journal of FisheriesSciences.com, 8 (3):??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

                                        //www.fisheriessciences.
                                        com/fisheries-aqua/the-
                                        lengthweight-relationship-
                                        of-bitterling-rhodeus-
                                        amarus-bloch-1782-in-freshwaters-
                                        of-turkey.php?aid=63.

Lindberg:2014:DLR

[Lin14] Jan Erik Lindberg. Dietary lysine requirement of juvenile Eurasian perch (Perca fluviatalis). Journal of FisheriesSciences.com, 8(2): ??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/dietary-
                                        lysine-requirement-of-
                                        juvenile-eurasian-perch-
                                        perca-fluviatalis.php?
                                        aid=268.

Lee:2007:RQT

[LL07] Jung-Lim Lee and Robert E. Levin. Rapid quantification of total bacteria on cod fillets by using realtime PCR. Journal of FisheriesSciences.com, 1(2): 58–67, April 2007. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/rapid-quantification-
                                        of-total-bacteria-on-cod-
                                        fillets-by-using-realtime-
                                        pcr.php?aid=1749.

Lu:2008:DBG

[LL08] Shuxia Lu and Robert E. Levin. Dominant bacterial genera of a tilapia fish farm and rapid typing of Vibrio isolates. Journal of FisheriesSciences.com, 2(2): 183–195, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/dominant-
bacterial-genera-of-a-
tilapia-fish-farm-and-
rapd-typing-of-vibrio-
isolates.php?aid=1640.

Lu:2010:STF

[LL10] Shuxia Lu and Robert E.

Levin. Shewanella in a tilapia fish farm. Journal of FisheriesSciences.com, 4(2): 159–170, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/shewanella-
in-a-tilapia-fish-farm.
php?aid=1037.

Luthada-Raswiswi:2019:NVN

[LRMO19] R. W. Luthada-Raswiswi, S. Mukaratirwa, and G. O’Brien. Nutritional value of the Nile crocodile (Crocodylus niloticus) meal for aquaculture feeds in South Africa. Journal of FisheriesSciences.com, 13(2):??, ???? 2019. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/nutritional-
value-of-the-nile-crocodile-
crocodylus-niloticus-meal-
for-aquaculture-feeds-

in-south-africa.php?aid=
24443.

Lo:2015:NSL

[LVA15] Tiamiyu Lo, Okomoda Vt, and Agbo Ao. Nutritional suitability of leucaena leaf meal in the diet of Clarias gariepinus. Journal of FisheriesSciences.com, 9 (2):??, ???? 2015. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/nutritional-
suitability-of-leucaena-
leaf-meal-in-the-diet-
of-clarias-gariepinus.
php?aid=5819.

Lubis:2016:SAU

[LWP^{+}16] Muhammad Zainuddin Lubis, Pratiwi Dwi Wulandari, Sri Pujiyati, Totok Hestirianoto, Juliana Rodrigues Moron, and Dwi Putra Imam Mahdi. Spectral analysis using Haar wavelet (original signal, denoised signal, residual signal) and source level for whistle sound of Dolphin (Tursiops aduncus). Journal of FisheriesSciences.com, 10(3):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/spectral-
analysis-using-haar-wavelet-
original-signal-denoised-
signal-residual-signal-
and-source-level-for-whistle-
sound-of-dolphin.php?aid=
9871.

Liu:2021:CSS

[LZ21] Qinghua Liu and Yuhong Zheng. Commentary on a successful spawning of American shad in a novel recirculating aquaculture system. Journal of FisheriesSciences.com, 15(1):1– 2, ???? 2021. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/commentary-
on-a-successful-spawning-
of-american-shad-in-a-
novel-recirculating-aquaculture-
system.php?aid=34741.

Michael:2019:RPE

[MAC19] Sakyi Essien Michael, Emmanuel Delwin Abarike, and Jia Cai. A review on the probiotic effects on haematological parameters in fish. Journal of FisheriesSciences.com, 13(3):??, ???? 2019. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/a-review-
on-the-probiotic-effects-
on-haematological-parameters-
in-fish.php?aid=24562.

Manasirli:2008:PDP

[Man08]  Meltem Manaşirli.  Pop-

ulation dynamical parameters of the Atlantic lizardfish (Synodus saurus) from the Babadillimani Bight (Silifke-Mersin) in Turkey. Journal of FisheriesSciences.com, 2(3):??, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/population-
                                        dynamical-parameters-of-
                                        the-atlantic-lizardfish-
                                        synodus-saurus-from-the-
                                        babadillimani-bight-silifkemersin-
                                        in-turkey.php?aid=1618.

Mancuso:2015:EFF

[Man15]  Monique Mancuso.  Ef-

fects of fish farming on marine environment. Journal of FisheriesSciences.com, 9 (3):??, ???? 2015. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/effects-
                                        of-fish-farming-on-marine-
                                        environment.php?aid=6896.

Maria:2009:NMR

[Mar09] Polikandrioti Maria. Nonmodifiable risk factors for ischemic stroke. Journal of FisheriesSciences.com, 3 (1):??, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/nonmodifiable-
                                        risk-factors-for-ischemic-
                                        stroke.php?aid=1517.

Marios:2011:MEP

[Mar11] Tsakas Marios. Management evaluation of public hospital units expenditures on infections with the use of quantitative analysis methods. Journal of FisheriesSciences.com, 5(2): ??, ???? 2011. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/management-
evaluation-of-public-hospital-
units-expenditures-on-
infections-with-the-use-
of-quantitative-analysis-
methods.php?aid=781.

MathewsDelgado:2014:PGA

[Mat14] Patrick Mathews Delgado.

Parasitism by Gussevia asota in gills of juveniles of Astronotus ocellatus cultured in the Peruvian Amazon. Journal of FisheriesSciences.com, 8(1):??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/parasitism-
by-gussevia-asota-in-gills-
of-juveniles-of-astronotus-
ocellatus-cultured-in-
the-peruvian-amazon.php?
aid=400.

Michael:2019:FSG

[MCAA19] Sakyi Essien Michael, Jia Cai, Ampofo-Yeboah Akwasi, and Aglago Adele. Fish smoking in Ghana: a review. Journal of FisheriesSciences.com, 13(3): ??, ???? 2019. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/fish-

                                        smoking-in-ghana-a-review.
                                        php?aid=24561.

Megahed:2014:SGS

[Meg14] Muhammed E. Megahed. Sustainable growth of shrimp aquaculture and protection of natural fisheries through biofloc production as alternative to fishmeal in shrimp feeds. Journal of FisheriesSciences.com, 8(4): ??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/sustainable-
                                        growth-of-shrimp-aquaculture-
                                        and-protection-of-natural-
                                        fisheries-through-biofloc-
                                        production-as-alternative-
                                        to-fishmeal-in-shrimp-
                                        feeds.php?aid=55.

Mehanna:2013:PDH

[Meh13] Sahar F. Mehanna. Population dynamics of the hooded cuttlefish Sepia prashadi (Winckworth, 1936) from the Omani Coastal Waters of the Arabian Sea. Journal of FisheriesSciences.com, 7 (1):??, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/population-
                                        dynamics-of-the-hooded-
                                        cuttlefish-sepia-prashadi-
                                        winckworth-1936-from-the-
                                        omani-coastal-waters-of-
                                        the-arabian-sea.php?aid=
                                        536.

Mehdipour:2016:BCH

[MG16] Neda Mehdipour and Mohammad Hasan Gearmi. Benthic communities on hard substrates and intracommunity relation with environmental factors in mesohaline estuarine. Journal of FisheriesSciences.com, 10 (3):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/benthic-
communities-on-hard-substrates-
and-intracommunity-relation-
with-environmental-factors-
in-mesohaline-estuarine.
php?aid=9872.

Manthey-Karl:2014:CCC

[MK14] Monika Manthey-Karl. Chemical composition, cholesterol, trace metals and amino acid composition of different canned fish products produces and sold in Turkey. Journal of FisheriesSciences.com, 8(1):??, ???? 2014. CODEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/chemical-
composition-cholesterol-
trace-metals-and-amino-
acid-composition-of-different-
canned-fish-products-produces-
and-sold-in-tu.php?aid=
393.

Mustafa:2015:ESO

[MKK15] Saleem Mustafa, Siti Nazwa Kharudin, and Annita Yong Seok Kian. Effect of simulated ocean acidification on chitin content in the shell of white shrimp, Litopenaeus vannamei Saleem Mustafa. Journal of FisheriesSciences.com, 9(2):??, ???? 2015. CODEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/effect-
of-simulated-ocean-acidification-
on-chitin-content-in-the-
shell-of-white-shrimp-
litopenaeus-vannamei-saleem-
mustafa.php?aid=5820.

Mishra:2017:EAB

[MKS17] Kirti Mishra, Patra S. K., and Kasturi Samantaray. Effect of algae based diets on growth performance, body composition and fatty acid profile of Indian major carp, Rohu labeo rohita, Hamilton. Journal of FisheriesSciences.com, 11 (3):??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/effect-
of-algae-based-diets-on-
growth-performance-body-
composition-and-fatty-
acid-profile-of-indian-
major-carp-rohu-labeo-
rohita-h.php?aid=19156.

Midani:2016:CIU

[ML16] Amaj Rahimi Midani and Sang-Go Lee. Confronting illegal, unreported and unregulated (IUU) fishing with proper port and flagged states policies: the case of South Korea and European Union. Journal of FisheriesSciences.com, 10 (3):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/confronting-
illegal-unreported-and-
unregulated-iuu-fishing-
with-proper-port-and-flagged-
states-policies-the-case-
of-south-korea-an.php?
aid=11058.

Mandal:2021:ECP

[MM21] Sayan Mandal and Basudev Mandal. Effect of COVID-19 pandemic in fisheries and fisherman of rural area. Journal of FisheriesSciences.com, 15(3):??, ???? 2021. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/effect-
of-covid19-pandemic-in-
fisheries-and-fisherman-
of-rural-area.pdf.

Medeiros:2015:PLD

[MMdA^{+}15] Fernanda S. Medeiros, Marina Marcuschi, Caio Rodrigo Dias de Assis, Janilson Felix da Silva, Talita da Silva Esposito, and Ranilson de Souza Bezerra. Potencial laundry detergent applications of mutton snapper (Lutjanus analis) proteases. Journal of FisheriesSciences.com, 9(3):??, ???? 2015. CODEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/potencial-
                                        laundry-detergent-applications-
                                        of-mutton-snapperlutjanus-
                                        analis-proteases.php?aid=
                                        6892.

Mol:2008:CFO

[Mol08] Sühendan Mol. Consumption of fish oil and its effects on human health. Journal of FisheriesSciences.com, 2(3):??, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/consumption-
                                        of-fish-oil-and-its-effects-
                                        on-human-health.php?aid=
                                        1591.

Mol:2011:QFR

[Mol11a] Sühendan Mol. The quality of fish from retail markets in Istanbul, Turkey. Journal of FisheriesSciences.com, 5(1):??, ???? 2011. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

                                        //www.fisheriessciences.
                                        com/fisheries-aqua/the-
                                        quality-of-fish-from-retail-
                                        markets-in-istanbul-turkey.
                                        php?aid=824.

Mol:2011:SVT

[Mol11b] Sühendan Mol. Sous-vide technology and applications to the seafoods. Journal of FisheriesSciences.com, 5 (4):??, ???? 2011. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/sousvide-
technology-and-applications-
to-the-seafoods.php?aid=
730.

Martinez:2017:RTR

[MP17] Joval N. Martinez and Philip Ian P. Padilla. Rotten thallus of red seaweed, Gracilariopsis heteroclada Zhang et Xia, is associated with agardigesting Bacillus spp. Journal of FisheriesSciences.com, 11(4):??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/rotten-
thallus-of-red-seaweed-
gracilariopsis-heteroclada-
zhang-et-xia-is-associated-
with-agardigesting-bacillus-
spp.php?aid=21557.

Michalczyk:2007:EGP

[MS07] Magdalena Michalczyk and Krzysztof Surówka. The effects of gravading process on the nutritive value of rainbow trout (Oncorhynchus mykiss). Journal of FisheriesSciences.com, 1(3):130– 138, ???? 2007. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
effects-of-gravading-process-

                                        on-the-nutritive-value-
                                        of-rainbow-trout-oncorhynchus-
                                        mykiss.php?aid=1746.

Mol:2010:TUI

[MU10] Sühendan Mol and Şafak Ulusoy. Türkiye’de su ürünleri işleme sektörünün sorunlari ve çözüm önerileri. (Turkish) [the problems of seafood processing sector in Turkey and suggestions]. Journal of FisheriesSciences.com, 4(2):152– 158, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

                                        //www.fisheriessciences.
                                        com/fisheries-aqua/the-
                                        problems-of-seafood-processing-
                                        sector-in-turkey-and-suggestions.
                                        php?aid=1039.

Mustafa:2017:SDC

[MUK17] Akhmad Mustafa, Muhammad Chaidir Undu, and Kamariah. Study on determination of categories of soil quality variable concentrations in brackish water ponds of Java Island, Indonesia. Journal of FisheriesSciences.com, 11(3):??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/study-on-
                                        determination-of-categories-
                                        of-soil-quality-variable-
                                        concentrations-in-brackish-
                                        water-ponds-of-java-island-
                                        indonesia.php?aid=20093.

Mohammadi:2015:EEE

[MYMZ15] Mona Mohammadi, Vahid Yavari, Hamid Mohammadiazarm, and Mohammad Zakeri. Evaluation effect of egg lecithin on growth performance, feeding parameters and lipoprotein fractions of (Mesopotamichthys sharpeyi). Journal of FisheriesSciences.com, 9(1):??, ???? 2015. CODEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/evaluation-
effect-of-egg-lecithin-
on-growth-performance-
feeding-parameters-and-
lipoprotein-fractions-
of-mesopotamichthys-sharpeyi.
php?aid=5824.

Naz:2012:PIE

[Naz12a] Mehmet Naz. The potential inhibitory effects of the commercial diets on protease activities of Dentex dentex larvae and live foods. Journal of FisheriesSciences.com, 6(3):??, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
potential-inhibitory-effects-
of-the-commercial-diets-
on-protease-activities-
of-dentex-dentex-larvae-
and-live-foods.php?aid=

Naz:2012:PSB

[Naz12b] Mehmet Naz. A preliminary study on the biochemical compositions of Naalginate microcapsules. Journal of FisheriesSciences.com, 6(2):??, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/a-preliminary-
                                        study-on-the-biochemical-
                                        compositions-of-naalginate-
                                        microcapsules.php?aid=
                                        650.

Nesara:2019:DDE

[NB19] K. M. Nesara and C. S. Bedi. Diatomix: a diatoms enhancer. Journal of FisheriesSciences.com, 13 (2):??, ???? 2019. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/diatomix-
                                        a-diatoms-enhancer.php?
                                        aid=24398.

Naik:2017:CSP

[NBBR17] Gulzar Naik, M. H. Balkhi, F. A. Bhat, and Mudasir Rashid. Comparative study of present and past cladoceran diversity in Manasbal Lake of Kashmir (India). Journal of FisheriesSciences.com, 11(1): ??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/comparative-

study-of-present-and-past-
cladoceran-diversity-in-
manasbal-lake-of-kashmir-
india.php?aid=18197.

Nogueira:2013:CCF

[NCA13] Natacha Nogueira, Nereida Cordeiro, and Maria João Aveiro. Chemical composition, fatty acids profile and cholesterol content of commercialized marine fishes captured in Northeastern Atlantic. Journal of FisheriesSciences.com, 7(3):271– 286, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/chemical-
composition-fatty-acids-
profile-and-cholesterol-
content-of-commercialized-
marine-fishes-captured-
in-northeastern-atlantic.
php?aid=441.

Nikkhah:2015:MAC

[Nik15] Akbar Nikkhah. Manipulating aquaculture circadian biology: a mounting science. Journal of FisheriesSciences.com, 9(2): ??, ???? 2015. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/manipulating-
aquaculture-circadian-
biology-a-mounting-science.
php?aid=6449.

Nie:2016:ITS

[NLM16] Xiaohua Nie, Shengli Lin, and Xianghe Meng. Identification of two selected lactic acid bacteria strains isolated from dry-cured fish and their behaviors in fermented fish sausage. Journal of FisheriesSciences.com, 10(1):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/identification-
                                        of-two-selected-lactic-
                                        acid-bacteria-strains-
                                        isolated-from-drycured-
                                        fish-and-their-behaviors-
                                        in-fermented-fish-saus.
                                        php?aid=8219.

Nandi:2016:MCC

[NPR16] N. C. Nandi, S. K. Pramanik, and M. K. Dev Roy. Mud crab culture: Relevance of species identity in production economics with reference to Sundarban Coast. Journal of FisheriesSciences.com, 10(4):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

                                        //www.fisheriessciences.
                                        com/fisheries-aqua/mud-
                                        crab-culture-relevance-
                                        of-species-identity-in-
                                        production-economics-with-
                                        reference-to-sundarban-
                                        coast.php?aid=17623.

Nadia:2017:CDE

[NS17] Besbes Nadia and Saloua Sadok. Comparison of DNAextraction methods suitable for PCR-based applications to identify shrimp species in commercial products. Journal of FisheriesSciences.com, 11(4):??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/comparison-
of-dnaextraction-methods-
suitable-for-pcrbased-
applications-to-identify-
shrimp-species-in-commercial-
products.php?aid=21279.

Olanrewaju:2015:RES

[OAA15] Sulaiman O. Olanrewaju, Amagee A., and Kader Asa. Review and experimental study of macro algae species and their native habitat. Journal of FisheriesSciences.com, 9(3): ??, ???? 2015. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/review-
and-experimental-study-
of-macro-algae-speciesand-
their-native-habitat.php?
aid=6679.

Osman:2018:BBN

[OAE^{+}18] Alaa G. M. Osman, Khaled Y.

AbouelFadl, Abd El Baset M. Abd El Reheem, Usama M. Mahmoud, Werner Kloas, and Mohsen A. Moustafa. Blood biomarkers in Nile tilapia Oreochromis niloticus niloticus and African catfish Clarias gariepinus to evaluate water quality of the River Nile. Journal of FisheriesSciences.com, 12 (1):??, ???? 2018. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/blood-biomarkers-
                                        in-nile-tilapia-oreochromis-
                                        niloticus-niloticus-and-
                                        african-catfish-clarias-
                                        gariepinus-to-evaluate-
                                        water-quality.php?aid=
                                        21796.

Olufeagba:2015:MVC

[OAO^{+}15] S. O. Olufeagba, S. E. Aladele, V. T. Okomoda, M. O. Sifau, D. A. Ajayi, O. T. Oduoye, O. A. Bolatito, D. S. Nden, A. S. Fabunmi-tolase, and T. Hassan. Morphological variation of cichlids from Kainji Lake, Nigeria. Journal of FisheriesSciences.com, 9 (3):??, ???? 2015. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/morphological-
                                        variation-of-cichlids-
                                        from-kainji-lake-nigeria.
                                        php?aid=6893.

Ozdemir:2010:DCC

[ÖEAÖ10] Süleyman Özdemir, Ercan Erdem, Hakan Aksu, and Zekiye Birinci Özdemir. Determination of catch composition and length–weight relationship of some pelagic fishes caught by pairly midwater trawl. Journal of FisheriesSciences.com, 4(4): ??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/determination-
of-catch-composition-and-
lengthweight-relationship-
of-some-pelagic-fishes-
caught-by-pairly-midwater-
trawl.php?aid=843.

Oehlenschlager:2008:EQP

[Oeh08] Joerg Oehlenschläger. Evaluation of quality parameters of canned and frozen blue whiting (Micromesistius poutassou). Journal of FisheriesSciences.com, 2 (4):??, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/evaluation-
of-quality-parameters-
of-canned-and-frozen-blue-
whiting-micromesistius-
poutassou.php?aid=1534.

Ogundiran:2018:TEW

[OF18] M. A. Ogundiran and O. O.

Fawole. Toxic effects of water pollution on two bioindicators of aquatic resources of Asa River, Nigeria. Journal of FisheriesSciences.com, 12(2): ??, ???? 2018. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/

                                        fisheries-aqua/toxic-effects-
                                        of-water-pollution-on-
                                        two-bioindicators-of-aquatic-
                                        resources-of-asa-river-
                                        nigeria.php?aid=22708.

Oguz:2011:CAT

[Oğu11] Hande Oğuz. Çukurovabölgesi (Adana–Türkiye) iklim koşullarinin Spirulina platensis (Cyanophyta)’in c-fikosiyanin miktarina etkisi. Journal of FisheriesSciences.com, 5(??):??, 2011. CODEN JFOIAQ. ISSN 1307-234X.

Ozvarol:2008:KBG

[Öİ08] Z. Arzu Becer Özvarol and Ramazan İkiz. Karacaören i baraj gölü’ndek sudak, Sander lucioperca (Lin., 1758) populasyonunun büyüme ve Ölüm oranlari ile stok analizi. (Turkish) [Growth, mortality and stock analysis of the pikeperch, Sander lucioperca (L., 1758) population of Karacaören I Dam Lake]. Journal of FisheriesSciences.com, 2(2):134– 145, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/growth-
                                        mortality-and-stock-analysis-
                                        of-the-pikeperch--sander-
                                        lucioperca-l-1758-population-
                                        of-karacaren-i-dam-lake.
                                        php?aid=1647.

Olanrewaju:2016:CSG

[OKA16] A. N. Olanrewaju, O. K.

Kareem, and E. K. Ajani. Comparative study of growth performance and survival of African catfish (Clarias gariepinus, Burchell 1822) fry in indoor and outdoor concrete and hapa culture system. Journal of FisheriesSciences.com, 10 (1):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/comparative-
study-of-growth-performance-
and-survival-of-african-
catfish-clarias-gariepinus-
burchell-1822-fry-in-indoor-
and-outdoor.php?aid=8213.

Oksuz:2008:CSB

[Öks08]  Abdullah Öksüz.  Com-

parison of some biochemical parameters in dark and light muscle of bonito (Sarda sarda). Journal of FisheriesSciences.com, 2 (3):??, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/comparison-
of-some-biochemical-parameters-
in-dark-and-light-muscle-
of-bonito-sarda-sarda.
php?aid=1597.

Oksuz:2010:ECF

[Öks10] Abdullah Öksüz. Element compositions, fatty acid profiles, and proximate compositions of marbled spinefoot (Siganus rivulatus, Forsskål, 1775) and dusky spinefoot (Siganus luridus, Ruppell, 1878). Journal of FisheriesSciences.com, 4(2): ??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/element-
                                        compositions-fatty-acid-
                                        profiles-and-proximate-
                                        compositions-of-marbled-
                                        spinefoot-siganus-rivulatus-
                                        forsskal-1775-and-dusky.
                                        php?aid=1032.

Oksuz:2012:CMY

[Öks12] Abdullah Öksüz. Comparison of meat yield, flesh colour, fatty acid, and mineral composition of wild and cultured Mediterranean amberjack (Seriola dumerili, Risso 1810). Journal of FisheriesSciences.com, 6 (2):??, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/comparison-
                                        of-meat-yield-flesh-colour-
                                        fatty-acid-and-mineral-
                                        composition-of-wild-and-
                                        cultured-mediterranean-
                                        amberjack-seriola-dume.
                                        php?aid=648.

Ozaydin:2008:KOY

[ÖL08] Okan Özaydin and Semih Leblebici. Karataş Önünde yaşayan (doğu akdeniz) eksi baliği’nin (Leiognathus klunzingeri (Steindachner, 1898)) büyümesi Üzerine bir Ön çalişma. (Turkish) [A preliminary study on the growth of Leiognathus klunzingeri (Steindachner, 1898) inhabiting off Karataş, Eastern Mediterranean]. Journal of FisheriesSciences.com, 2(4): 672–676, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/a-preliminary-
study-on-the-growth-of-
leiognathus-klunzingeri-
steindachner-1898-inhabiting-
off-karata-eastern-mediterranean.
php?aid=1550.

Olukolajo:2013:STG

[OO13] Soyinka Olufemi Olukolajo and Lawal-Are Aderonke Omolara. Salinity tolerance of grey mullet, Mugil cephalus (Linnaeus) fry in the laboratory. Journal of FisheriesSciences.com, 7(4): 292–296, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/salinity-
tolerance-of-grey-mullet-
mugil-cephalus-linnaeus-
fry-in-the-laboratory.
php?aid=419.

Olufeagba:2016:SAB

[OOB16] S. O. Olufeagba, V. T. Okomoda, and Benny. Some aspects of the biology of Labeo coubie Ruppell, 1832 and Labeo senegalensis Valenciennes, 1842 from Lower River Benue. Journal of FisheriesSciences.com, 10 (2):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/some-
aspects-of-the-biology-
of-labeo-coubie-ruppell-
1832-and-labeo-senegalensis-
valenciennes-1842-from-
lower-river-benue.php?
aid=9183.

Ozogul:2007:NDS

[ÖÖK07] Fatih Özoğul, Yeşim Özoğul, and Esmeray Kuley. Nucleotide degradation in sardine (Sardina pilchardus) stored in different storage condition at 4C. Journal of FisheriesSciences.com, 1(1): 13–19, February 2007. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/nucleotide-
degradation-in-sardine-
sardina-pilchardus-stored-
in-different-storage-condition-
at-4c.php?aid=1769.

Opiyo:2014:CES

[Opi14] Mary A. Opiyo. Combined effects of stocking density and background colour on growth performance and survival of Nile tilapia (Oreochromis niloticus, L.) fry reared in aquaria. Journal of FisheriesSciences.com, 8 (3):??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/combined-
effects-of-stocking-density-
and-background-colour-
on-growth-performance-
and-survival-of-nile-tilapia-
oreochromis-niloticu.php?
aid=75.

Orsay:2010:CEV

[Ors10a] Bülent Orsay. Catching efficiency of various hanging ratio and colour monofilament gill nets. Journal of FisheriesSciences.com, 4 (4):??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/catching-
efficiency-of-various-
hanging-ratio-and-colour-
monofilament-gill-nets.
php?aid=852.

Orsay:2010:CCE

[Ors10b] Bülent Orsay. The comparison of catch efficiency of monofilament gill nets different colour and different hanging ratio on seasonal variations. Journal of FisheriesSciences.com, 4 (3):??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
comparison-of-catch-efficiency-
of-monofilament-gill-nets-
different-colour-and-different-
hanging-ratio-on-seasonal-
variations.php?aid=1023.

Ozulug:2007:TNR

[ÖTGG07] Müfit Özulug, Ali Serhan Tarkan, Özcan Gaygusuz, and Çigdem Gürsoy. Two new records for the fish fauna of Lake Sapanca Basin (Sakaraya, Turkey). Journal of FisheriesSciences.com, 1 (3):152–159, ???? 2007. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/two-new-
                                        records-for-the-fish-fauna-
                                        of-lake-sapanca-basin-
                                        sakaraya-turkey.php?aid=
                                        174.

Oz:2010:UNZ

[Öz10] Meryem Öz. Using of the natural zeolite clinoptilolite in transportation of fingerling trout (Oncorynchus mykiss, W., 1792). Journal of FisheriesSciences.com, 4(3):??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/using-of-
                                        the-natural-zeolite-clinoptilolite-
                                        in-transportation-of-fingerling-
                                        trout-oncorynchus-mykiss-
                                        w-1792.php?aid=1016.

Ozalp:2009:DSF

[Öza09] Berna Özalp. Determination of some functional and technological properties of octopus (Octopus vulgaris C.), Calamary (Illex coindetti V.), mussel (Mytilus galloprovincialis L.) and cuttlefish (Sepia officinalis L.) meats. Journal of FisheriesSciences.com, 3(4): ??, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/.

Ozcan:2009:LWW

[Özc09] Tahir Özcan. Length/widthweight relationships of the Mediterranean green crab Carcinus aestuarii Nardo, 1847 in the Homa Lagoon, Aegean Sea Turkey. Journal of FisheriesSciences.com, 3(1):??, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/lengthwidthweight-
relationships-of-the-mediterranean-
green-crab-carcinus-aestuarii-
nardo-1847-in-the-homa-
lagoon-aegean-sea-turkey.
php?aid=1497.

Ozcan:2013:NLG

[Özc13] Tahir Özcan. New locality for grass shrimp, Palaemonetes antennarius (H. Milne– edwards, 1837) (Decapoda: Palaemonidae) in the Seyhan River Basin (Turkey). Journal of FisheriesSciences.com, 7(3):??, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/new-
locality-for-grass-shrimp-

                                        palaemonetes-antennarius-
                                        h-milneedwards-1837-decapoda-
                                        palaemonidae-in-the-seyhan-
                                        river-basin-turkey.php?
                                        aid=443.

Ozdemir:2011:KFA

[Özd11] Süleyman Özdemir. Karadenizin farkli av sahalarinda demersal trol ile avlanan mezgit (Merlangius merlangus euxinus, N.) ve barbunya (Mullus barbatus ponticus, E.) baliklarinin av miktarlari ve boy kompozisyonlarinin karşilaştirilmasi. Journal of FisheriesSciences.com, 5(??):??, 2011. CODEN JFOIAQ. ISSN 1307-234X.

Ozgul:2009:RMS

[Özg09]  Aytaç Özgül.  Regula-

tions and management of spearfishing. Journal of FisheriesSciences.com, 3(2): ??, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/regulations-
                                        and-management-of-spearfishing.
                                        php?aid=1483.

Ozkan:2010:EB

[Özk10]  Özden Özkan.  Editorial

board. Journal of FisheriesSciences.com, 4(1):??, ???? 2010. CODEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/editorial-
                                        board.pdf.

Ozkan:2014:CEB

[Özk14a] Özden Özkan. Cover and editorial board. Journal of FisheriesSciences.com, 8 (2):??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/.

Ozkan:2014:Ia

[Özk14b] Özden Özkan. Issue 1. Journal of FisheriesSciences.com, 8(1):??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/issue-1.
pdf.

Ozkan:2014:Ib

[Özk14c] Özden Özkan. Issue 2. Journal of FisheriesSciences.com, 8(2):??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/issue-2.
pdf.

Ozogul:2010:UAT

[Özo10] Fatih Özoğul. Usefulness of API test strips for identification of bacterial flora in blue crab (Callinectes sapidus) caught from Akyatan Lagoon (Adana–Turkey). Journal of FisheriesSciences.com, 4(1):??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/

fisheries-aqua/usefulness-
of-api-test-strips-for-
identification-of-bacterial-
flora-in-blue-crab-callinectes-
sapidus-caught-from-akyatan-
lagoon-adanaturkey.php?
aid=1071.

Ozogul:2012:CCM

[Özo12] Yeşim Özoğul. The chemical composition and meat yield of sexually mature cuttlefish (Sepia officinalis). Journal of FisheriesSciences.com, 6 (2):??, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
chemical-composition-and-
meat-yield-of-sexually-
mature-cuttlefish-sepia-
officinalis.php?aid=655.

Ozpolat:2010:SFR

[Özp10] Emine Özpolat. The sensory features of rainbow trout (Oncorhynchus mykiss) marinades prepared in different asetic acid and eugenol ratios. Journal of FisheriesSciences.com, 4(4):??, ???? 2010. CODEN JFOIAQ. ISSN 1307-234X. URL https:/

/www.fisheriessciences.
com/fisheries-aqua/the-
sensory-features-of-rainbow-
trout-oncorhynchus-mykiss-
marinades-prepared-in-
different-asetic-acid-
and-eugenol-ratios.php?
aid=859.

Ozturk:2007:EOD

[Ozt07] Turkay Ozturk. Effects of oxidized diet on growth and the liver of rainbow trout, (Oncorhynchus mykiss Walbaum 1792). Journal of FisheriesSciences.com, 1(2): 81–87, April 2007. CODEN JFOIAQ. ISSN 1307-234X.

Ozturk:2010:IHI

[Özt10a] Türkay Öztürk. Ichthyophonus hoferi infection in goldfish (Carassius auratus). Journal of FisheriesSciences.com, 4(4):??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/ichthyophonus-
hoferi-infection-in-goldfish-
carassius-auratus.php?
aid=866.

Ozturk:2010:VES

[Özt10b] Türkay Öztürk. In vitro effect of some therapeutants on Ichthyophthirius multifiliis. Journal of FisheriesSciences.com, 4(3):??, ???? 2010. CODEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/in-vitro-
effect-of-some-therapeutants-
on-ichthyophthirius-multifiliis.
php?aid=1027.

Ozta:2012:CCC

[Özt12] Meryem Özta. Comparation of CPUEs for catching whiting (Merlangius merlangus (Linnaeus, 1758)) caught by gillnets from three different areas in the Southeast Black Sea (Ordu–Giresun). Journal of FisheriesSciences.com, 6(4):??, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/comparation-
of-cpues-for-catching-
whiting-merlangius-merlangus-
linnaeus-1758-caught-by-
gillnets-from-three-different-
areas-in-the.php?aid=577.

Ozturk:2013:IAF

[Özt13] M. Ouz Öztürk. Investigations on Argulus foliaceus (Crustacea, Branchiura) infestation fauna of chub (Squalius cephalus, L. 1758) from Lake Dam Serban, Afyonkarahisar. Journal of FisheriesSciences.com, 7 (4):??, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/investigations-
on-argulus-foliaceus-crustacea-
branchiura-infestation-
fauna-of-chub-squalius-
cephalus-l-1758-from-lake-
dam-serban-afyonkarahisar.
php?aid=428.

Ozulug:2011:PSO

[Özu11] Oya Özuluğ. A preliminary study on ostracoda (Crustacea) fauna of the Istranca Streams–Turkey. Journal of FisheriesSciences.com, 5 (2):??, ???? 2011. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/a-preliminary-
study-on-ostracoda-crustacea-
fauna-of-the-istranca-
streamsturkey.php?aid=
777.

Ozvarol:2013:DCM

[Özv13a] Yaşar Özvarol. The distribution of Carinaria mediterranea Blainville, 1825 (heteropods) in the Mediterranean shores of Turkey. Journal of FisheriesSciences.com, 7(3):??, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
distribution-of-carinaria-
mediterranea-blainville-
1825-heteropods-in-the-
mediterranean-shores-of-
turkey.php?aid=450.

Ozvarol:2013:DAG

[Özv13b] Yaşar Özvarol. The distribution of the alien gastropod Melibe viridis (Kelaart, 1858) (Opistobranchia: Tethyidae) in the Mediterranean Shores of Turkey. Journal of FisheriesSciences.com, 7(2):??, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
distribution-of-the-alien-
gastropod-melibe-viridis-

                                        kelaart-1858-opistobranchia-
                                        tethyidae-in-the-mediterranean-
                                        shores-of-turkey.php?aid=
                                        494.

Ozvarol:2014:DLR

[Özv14a] Yaar Özvarol. The distribution of the Lessepsian ragged sea hare, Bursatella leachii Blainville, 1817 (Aplysiidae) in the Mediterranean Coast of Turkey. Journal of FisheriesSciences.com, 8 (4):??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

                                        //www.fisheriessciences.
                                        com/fisheries-aqua/the-
                                        distribution-of-the-lessepsian-
                                        ragged-sea-harebursatella-
                                        leachii-blainville-1817-
                                        aplysiidae-in-the-mediterranean-
                                        coast-of-turk.php?aid=
                                        46.

Ozvarol:2014:FRS

[Özv14b] Yaar Özvarol. First report of the spiny blaasop Tylerius spinosissimus (Regan, 1908) (Actinopterygii: Tetraodontidae) from Gulf of Antalya, Turkey. Journal of FisheriesSciences.com, 8(3):??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/first-report-
                                        of-the-spiny-blaasop-tylerius-
                                        spinosissimus-regan-1908-
                                        actinopterygii-tetraodontidae-
                                        from-gulf-of-antalya-turkey.
                                        php?aid=62.

Ozyurt:2009:CSC

[Özy09] Caner Enver Özyurt. Characteristics of shrimp catching with trammel nets in İskenderun Bay. Journal of FisheriesSciences.com, 3 (4):??, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/characteristics-
of-shrimp-catching-with-
trammel-nets-in--iskenderun-
bay.php?aid=1162.

P:2019:MFT

[P19] Ramachandra Reddy P.

Methyl farnesoate through feed: a growth manipulator in female crab Oziothelphusa senex senex. Journal of FisheriesSciences.com, 13(1):??, ???? 2019. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/methyl-
farnesoate-through-feed-
a-growth-manipulator-in-
female-crab-oziothelphusa-
senex-senex.php?aid=24195.

Parenti:2013:PLW

[Par13] Paolo Parenti. Perca luth Walbaum, 1792, a senior synonym of the meagre, Argyrosomus regius (Asso, 1801) (Perciformes: Sciaenidae), invalidated by “Reversal of precedence”. Journal of FisheriesSciences.com, 7(1):

??, ????  2013.  CO-

DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/perca-luth-
walbaum-1792-a-senior-
synonym-of-the-meagre-
argyrosomus-regius-asso-
1801-perciformes-sciaenidae-
invalidated-by-reversal-
of-precedence.php?aid=
541.

Paital:2016:RMF

[PC16] Biswaranjan Paital and Gagan Bihari Nityananda Chainy. Redox metabolism in fishes under thermal stress warrants more attention. Journal of FisheriesSciences.com, 10(1): ??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/redox-metabolism-
in-fishes-under-thermal-
stress-warrants-more-attention.
php?aid=8212.

Parlak:2015:DFSa

[PÇD^{+}15a] Akif Evren Parlak, Metin Çalta, Mustafa Düşükcan, Mücahit Eroğlu, and Ökke Yılmaz. The determination of fat-soluble vitamins, cholesterol content and the fatty acid compositions of shabut (Arabibarbus grypus, Heckel 1843) from Keban Dam Lake, Elazig, Turkey. Journal of FisheriesSciences.com, 9 (1):??, ???? 2015. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
determination-of-fatsoluble-
vitamins-cholesterol-contentand-
the-fatty-acid-compositions-
of-shabut-arabibarbus-
grypusheckel-184.php?aid=
5816.

Parlak:2015:DFSb

[PÇD^{+}15b] Akif Evren Parlak, Metin Çalta, Mustafa Düşükcan, Mücahit Eroğlu, and Ökke Yılmaz. The determination of fat-soluble vitamins, cholesterol content and the fatty acid compositions of shabut (Arabibarbus grypus, Heckel 1843) from Keban Dam Lake, Elazig, Turkey. Journal of FisheriesSciences.com, 9 (3):??, ???? 2015. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
determination-of-fatsoluble-
vitamins-cholesterolcontent-
and-the-fatty-acd-compostons-
of-shabutarabibarbus-grypus-
heckel-1843-from-keban-
dam-lakeelazig-turkey.
php?aid=6676.

Paulino:2016:AGT

[PCV^{+}16] M. S. Paulino, G. C., Veras, V. O., Felizardo, L. D., Solis-Murgas, and Freitas R. T. F. Assessment of gametes in tilapia Oreochromis niloticus: Effects of body weight in a new lineage. Journal of FisheriesSciences.com, 10 (2):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/assessment-
                                        of-gametes-in-tilapia-
                                        oreochromis-niloticus-
                                        effects-of-body-weight-
                                        in-a-new-lineage.php?aid=
                                        9103.

Pepping:2021:CRF

[Pep21] Michelle Y. Pepping. Commentary on “Rapture Facilitates Inexpensive and High-Throughput Parent-Based Tagging in Salmonids”. Journal of FisheriesSciences.com, 15(2):14–15, ???? 2021. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/commentary-
                                        on-rapture-facilitates-
                                        inexpensive-and-highthroughput-
                                        parentbased-tagging-in-
                                        salmonids.php?aid=36172.

Percn:2010:NHT

[Per10] Fatih Perçn. A new harvesting technique for tunas: Lupara. Journal of FisheriesSciences.com, 4(2): ??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/a-new-harvesting-
                                        technique-for-tunas-lupara.
                                        php?aid=1030.

Polat:2009:KSY

[PiYB09] Nazmi Polat, Yasemin inceismail, Savaş Yılmaz, and Derya Bostancı. Karadeniz (samsun)’de yaşayan zargana (Belone belone L., 1761)’da yaş tayini, yaş-boy ve boyağirlik ilişkileri. (Turkish) [Age determination, age– length and length–weight relationships of garfish (Belone belone L., 1761) in the Black Sea (Samsun)]. Journal of FisheriesSciences.com, 3(3):187–198, ???? 2009. CODEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/age-
determination-agelength-
and-lengthweight-relationships-
of-garfish-belone-belone-
l-1761-in-the-black-sea-
samsun.php?aid=1463.

Pappa:2017:ASG

[PKMB17] Valentini A. Pappa, Panagiotis Kapsis, Elena Mente, and Panagiotis Berillis. Aquaponics software in Greece. Journal of FisheriesSciences.com, 11(3): ??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/aquaponics-
software-in-greece.php?
aid=19155.

Plaza:2017:SLL

[PLS^{+}17] Julian Plaza, Yanett Leyton, Camila Sayes, Cristian Mejias, and Carlos Riquelme. Seriola lalandi larviculture with probiotic supplements in mesocosm systems. Journal of FisheriesSciences.com, 11(3):??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/seriola-
                                        lalandi-larviculture-with-
                                        probiotic-supplements-
                                        in-mesocosm-systems.php?
                                        aid=20088.

Pires:2018:EMF

[PMOBF18] Antonio Adami Pires, Renata Manelli-Oliveira, and Julio Boschini-Filho. Evaluation of micronuclei frequencies and other nuclear alterations in erythrocytes of Tilapia rendalli (Perciformes: Cichlidae) Induced by Roundup herbicide. Journal of FisheriesSciences.com, 12(1):??, ???? 2018. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/evaluation-
                                        of-micronuclei-frequencies-
                                        and-other-nuclearalterations-
                                        in-erythrocytes-of-tilapia-
                                        rendalli-perciformescichlidae-
                                        induce.php?aid=22140.

Polat:2008:ADS

[Pol08a] Nazmi Polat. Age determination with some bony structures and length-frequency method of sprat (Sprattus sprattus L.,1758) in the Black Sea. Journal of FisheriesSciences.com, 2(2): ??, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/age-
determination-with-some-
bony-structures-and-lengthfrequency-
method-of-sprat-sprattus-
sprattus-l1758-in-the-
black-sea.php?aid=1648.

Polat:2008:DSD

[Pol08b] Sevim Polat. The distribution and seasonal dynamics of diatom and dinoflagellates in the İskenderun Bay (Northeastern Mediterranean). Journal of FisheriesSciences.com, 2(2):??, ???? 2008. CODEN JFOIAQ. ISSN 1307-234X. URL https:/

/www.fisheriessciences.
com/fisheries-aqua/the-
distribution-and-seasonal-
dynamics-of-diatom-and-
-dinoflagellates-in-the-
304skenderun-bay-northeastern-
mediterranean.php?aid=
1645.

Polat:2012:IKK

[Pol12] Sevim Polat. iskenderun körfezi (kuzeydoğu akdeniz) kıyısalında dağılım gösteren bazı kahverengi ve kırmızı makroalg türlerinin protein, lipit ve yağ asiti içerikleri. Journal of FisheriesSciences.com, 6(2):??, ???? 2012. CODEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/protein-
lipid-and-fatty-acid-composition-

                                        of-some-brown-and-red-
                                        seaweeds-from-the-coast-
                                        of-skenderun-bay-northeastern-
                                        mediterranean.php?aid=
                                        654.

Panichev:2016:IDC

[PP16] N. A. Panichev and S. E. Panicheva. Influence of different cooking procedure on the Hg concentration in fish. Journal of FisheriesSciences.com, 10(1): ??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/influence-
                                        of-different-cooking-procedure-
                                        on-the-hg-concentration-
                                        in-fish.php?aid=8227.

Paital:2016:HSU

[PRI16] Biswaranjan Paital and Georgina A. Rivera-Ingraham. High speed urbanization and its effects on aquatic food chain especially on fish in Bata River of Odisha, India. Journal of FisheriesSciences.com, 10(4): ??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

                                        //www.fisheriessciences.
                                        com/fisheries-aqua/high-
                                        speed-urbanization-and-
                                        its-effects-on-aquatic-
                                        food-chain-especially-
                                        on-fish-in-bata-river-
                                        of-odisha-india.php?aid=
                                        17235.

Rahman:2017:FSS

[RAB^{+}17] Md. Ashikur Rahman, Shirin Akter, Shuva Bhowmik, Md. Mehedi Hasan Pramanik, and A. K. M. Nowsad Alam. Fish in super-shops: a new dimension of fish marketing system in Dhaka (Bangladesh). Journal of FisheriesSciences.com, 11 (1):??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/fish-in-
supershops-a-new-dimension-
of-fish-marketing-system-
in-dhaka-bangladesh.php?
aid=17799.

Rad:2013:PIM

[Rad13] Ferit Rad. Preliminary investigation on morphometric and biometric characteristics of female silver and yellow, Anguilla anguilla, from Eastern Mediterranean (Gøksu Delta/Turkey). Journal of FisheriesSciences.com, 7(3):??, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/preliminary-
investigation-on-morphometric-
and-biometric-characteristics-
of-female-silver-and-yellow-
anguilla-anguilla-from-
eastern.php?aid=445.

Ramachandran:2017:PSP

[Ram17] Karthik Ramachandran. Present status and prospect trends of probiotics in shrimp aquaculture. Journal of FisheriesSciences.com, 11(2): ??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/present-
status-and-prospect-trends-
of-probiotics-in-shrimp-
aquaculture.php?aid=18966.

Ray:2017:AGM

[RBT^{+}17] Candis Ray, Noemi Bujan, Andrea Tarnecki, Allen D. Davis, C. Browdy, and C. R. Arias. Analysis of the gut microbiome of Nile tilapia Oreochromis niloticus L. fed diets supplemented with Previda and Saponin. Journal of FisheriesSciences.com, 11 (2):36–45, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/analysis-
of-the-gut-microbiome-
of-nile-tilapia-oreochromis-
niloticus-l-fed-diets-
supplemented-with-previda-
and-saponin.php?aid=18456.

Rasal:2016:STF

[RCP^{+}16] Kiran Dashrath Rasal, Vemulawada Chakrapani, Swagat Kumar Patra, Arun S. Ninawe, Jitendra Kumar Sundaray, Pallipuram Jayasankar, and Hirak Kumar Barman. Status of transgenic fish production with emphasis on development of food fishes and novel color varieties of ornamental fish: Implication and future perspectives. Journal of FisheriesSciences.com, 10(3):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/status-
of-transgenic-fish-production-
with-emphasis-on-development-
of-food-fishes-and-novel-
color-varieties-of-ornamental-
fish-impl.php?aid=11084.

Ramos:2016:SAR

[RCR^{+}16] G. B. Ramos, R. Corcino, H. Rivera, N. Alfonso, G. Sia Su, and R. Arcilla. Some aspects of the reproductive biology of two freshwater fish species, Puntius tumba and Anabas testudineus, of Lake Lanao, Philippines. Journal of FisheriesSciences.com, 10(3):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/some-
aspects-of-the-reproductive-
biology-of-two-freshwater-
fish-species-puntius-tumba-
-and-anabas-testudineus-
of-lake-lanao-philip.php?
aid=9376.

Raimundo:2017:SRB

[RCV^{+}17] Joana Raimundo, Miguel Caetano, Carlos Vale, Rui Coelho, Mário Mil-Homens, and Miguel Neves Dos Santos. Searching relationships between tissue elemental concentrations and geographical distribution of bigeye tuna (Thunnus obesus) from the South Atlantic Ocean. Journal of FisheriesSciences.com, 11 (2):??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/searching-
                                        relationships-between-
                                        tissue-elemental-concentrations-
                                        and-geographical-distribution-
                                        of-bigeye-tunathunnus-
                                        obesus-from-th.php?aid=
                                        18937.

Roloson:2016:FIG

[RGB^{+}16] Scott D. Roloson, Rachel L. Gould, Dave R. Barton, Frederick, W. Goetz, Andrew J. Jasonowicz, Christopher Beierling, and Michael R. van den Heuvel. Factors influencing growth variability in three Northern Alberta populations of yellow perch (Perca flavescens). Journal of FisheriesSciences.com, 10(4):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/factors-
                                        influencing-growth-variability-
                                        in-three-northern-alberta-
                                        populations-of-yellow-
                                        perch-perca-flavescens.
                                        php?aid=13161.

Rabbaniha:2013:STP

[RGO13] Mahnaz Rabbaniha, Javad Ghasemzadeh, and Fereydoon Owfi. Spatial and temporal patterns of fish larvae assemblage in the northern coastal waters of Persian Gulf along the Bushehr Province shoreline. Journal of FisheriesSciences.com, 7 (2):141–151, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/spatial-
and-temporal-patterns-
of-fish-larvae-assemblage-
in-the-northern-coastal-
waters-of-persian-gulf-
along-the-bushehr-province-
shoreline.php?aid=499.

Razak:2017:GUI

[RMKM17] Nurul Asmaa Abd Razak, Hizmawati Madzin, Fatimah Khalid, and Mas Rina Mustaffa. Gonad ultrasonography image preprocessing for mahseer (Tor tombroides). Journal of FisheriesSciences.com, 11 (3):??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/gonad-ultrasonography-
image-preprocessing-for-
mahseer-tor-tombroides.
php?aid=19573.

R:2015:ECC

[RPCC15] Karthik R., Angelin C.

Pushpam, Yashika Chelvan, and Vanitha M. C. Efficacy of cheatoceros calcitrans, enriched Artemia salina, Bacillus stratosphericus (amet1601) nov., and nitrifier bacterial consortium as probiotics on Litopenaeus vannamei culture. Journal of FisheriesSciences.com, 9 (4):??, ???? 2015. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/efficacy-
of-cheatoceros-calcitrans-
enriched-artemia-salina-
bacillus-stratosphericus-
amet1601-nov-and-nitrifier-
bacterial-consortiu.php?
aid=8190.

Radwan:2021:SPN

[RSA^{+}21] Mahmoud Kh Radwan, Sabry M. Shehata, Yasser M. Abdelhadi, Ramadan A. Mohammed, and Mohamed H. Mohamed. Survey parasitic of Nile tilapia (Oreochromis niloticus) with special reference to their immune status, histopathological aspects and nutritional values. Journal of FisheriesSciences.com, 15 (1):1–7, ???? 2021. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/survey-
parasitic-of-nile-tilapia-
oreochromis-niloticus-
with-special-reference-
to-their-immune-status-
histopathological-aspects-
and.php?aid=34740.

Rad:2016:THM

[RSSG16] Faezeh Hedayati Rad, Arezoo Solimani, Ali Dadolahi Sohrab, and Mohammad Hasan Gerami. Tracking heavy metals in sediments, muscle and skeleton of cynoglossus arel with application of new CSI index for assessing contamination in sediments. Journal of FisheriesSciences.com, 10(1):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/tracking-
heavy-metals-in-sediments-
muscle-and-skeleton-of-
cynoglossus-arel-with-
application-of-new-csi-
index-for-assessing-contami.
php?aid=8224.

Sodamola:2017:ENM

[SAA^{+}17] M. O. Sodamola, Y. A.

Adejola, K. T. Akanni, I. K. Ibrahim, A. K. Fajobi, and D. C. Aniche. Effect of nutmeg (Myristica fragrans) as an additive on the growth performance of juvenile catfish (Clarias gariepinus). Journal of FisheriesSciences.com, 11(2):??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/effect-
of-nutmeg-myristica-fragrans-
as-an-additive-on-the-
growth-performance-of-
juvenile-catfish-clarias-
gariepinus.php?aid=18938.

Sac:2014:FNR

[Saç14] Gülah Saç. Five new records for the fish fauna of Durusu Lake Basin (Istanbul). Journal of FisheriesSciences.com, 8(4):??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/five-new-
                                        records-for-the-fish-fauna-
                                        of-durusu-lake-basin-istanbul.
                                        php?aid=4.

Saglam:2011:PSE

[Sağ11] Naim Sağlam. Protection and sustainability, exportation of some species of medicinal leeches (Hirudo medicinalis L., 1758 and Hirudo verbana Carena, 1820). Journal of FisheriesSciences.com, 5(1):??, ???? 2011. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/protection-
                                        and-sustainability-exportation-
                                        of-some-species-of-medicinal-
                                        leeches-hirudo-medicinalis-
                                        l-1758-and-hirudo-verbana-
                                        carena.php?aid=828.

Sahan:2009:DLC

[Şah09]  Aysel Şahan.  Determi-

nation of leukocyte cell types in some cyprinidae caught from different regions of Ceyhan River (Adana– Turkey). Journal of FisheriesSciences.com, 3(2):??,

????  2009.  CODEN

JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/determination-
of-leukocyte-cell-types-
in-some-cyprinidae-caught-
from-different-regions-
of-ceyhan-river-adanaturkey.
php?aid=1480.

Sahin:2010:SVL

[Şah10] Temel Şahin. Seasonal variation in length, weight, and sex distribution of flounder (Platichtys flesus luscus Pallas, 1871) in the South-Eastern Black Sea. Journal of FisheriesSciences.com, 4(3):??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/seasonal-
variation-in-length-weight-
and-sex-distribution-of-
flounder-platichtys-flesus-
luscus-pallas-1871-in-
the-southeastern-blac.
php?aid=1021.

Sakai:2015:FI

[Sak15] Masahiro Sakai. Fish immunostilumants. Journal of FisheriesSciences.com, 9 (4):??, ???? 2015. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/fish-
immunostilumants.php?aid=
8189.

Saler:2009:CCC

[Sal09]  Serap Saler.  Cladocera

and Copepoda (Crustacea) fauna of Selı Stream (Elazıg– Turkey). Journal of FisheriesSciences.com, 3(4):??, ???? 2009. CODEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/cladocera-
                                        and-copepoda-crustacea-
                                        fauna-of-sel-stream-elaziturkey.
                                        php?aid=1161.

Saler:2011:KCE

[Sal11]  Serap Saler.  Kürk çayi

(Elaziğ-Türkiye) zooplanktonu. Journal of Fisheries-Sciences.com, 5(??):??, 2011. CODEN JFOIAQ. ISSN 1307-234X.

Saler:2014:ZUD

[Sal14]  Serap Saler.  Zooplank-

ton of Uzuncayir Dam Lake (Tuncelli–Turkey). Journal of FisheriesSciences.com, 8 (1):??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/zooplankton-
                                        of-uzuncayir-dam-lake-
                                        tuncelliturkey.php?aid=
                                        387.

Sanjay:2016:LWR

[SAMN16] Dakua Sanjay, Santosh Kumar Abujam, Islam Mofidul, and Basumatary Nilam. Length–weight relationship and condition factor of Parluciosoma daniconius (Hamilton) from the Upper Assam, India. Journal of FisheriesSciences.com, 10 (1):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/lengthweight-
relationship-and-condition-
factor-of-parluciosoma-
daniconius-hamilton-from-
the-upper-assam-india.
php?aid=8222.

Stathopoulou:2018:BNT

[SBL^{+}18] P. Stathopoulou, P. Berillis, E. Levizou, M. Sakellariou-Makrantonaki, A. K. Kormas, A. Aggelaki, P. Kapsis, N. Vlahos, and E. Mente. Basil and Nile tilapia production in a small scale aquaponic system. Journal of FisheriesSciences.com, 12 (4):??, ???? 2018. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/basil-and-
nile-tilapia-production-
in-a-small-scale-aquaponic-
system.php?aid=23582.

Schmidt:2017:EBG

[SBW^{+}17] J. Schmidt, A. A. Bischoff, M. Weiß, S. K. Kim, S. Frickenhaus, M. J. Slater, and B. H. Buck. Effect of beta-1-3-glucan and mannans on growth and fitness of starry flounder (Platichthys stellatus): a potential new candidate for aquaculture in temperate regions. Journal of FisheriesSciences.com, 11 (3):??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/effect-
                                        of-beta13glucan-and-mannans-
                                        on-growth-and-fitness-
                                        of-starry-flounder-platichthys-
                                        stellatus-a-potential-
                                        new-candidate-for-aq.php?
                                        aid=19538.

Soydemir:2008:ELC

[SCEA08] Nuray Çiftçi (Soydemir), Bedii Cicik, Cahit Erdem, and Özcan Ay. Effects of lead concentrations on sera parameters and hematocrit levels in Anguilla anguilla (Linnaeus, 1758). Journal of FisheriesSciences.com, 2 (3):??, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/effects-
                                        of-lead-concentrations-
                                        on-sera-parameters-and-
                                        hematocrit-levels-in-anguilla-
                                        anguilla-linnaeus-1758.
                                        php?aid=1605.

Sedat:2014:UBT

[Sed14] Gündodu Sedat. The usage of Bayes theory in fisheries sciences. Journal of FisheriesSciences.com, 8(1): ??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

                                        //www.fisheriessciences.
                                        com/fisheries-aqua/the-
                                        usage-of-bayes-theory-

in-fisheries-sciences.
php?aid=390.

Selcuk:2014:RHM

[Sel14] Arif Selçuk. A rapid HPLC method for determination of 4-hexylresorcinol residues in shrimp. Journal of FisheriesSciences.com, 8(2): ??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/a-rapid-
hplc-method-for-determination-
of-4hexylresorcinol-residues-
in-shrimp.php?aid=262.

Sen:2009:SED

[Şen09] Halil Şen. Spawning, egg development and incubation of cuttlefish (Sepia officinlis L. 1758) in controlled conditions. Journal of FisheriesSciences.com, 3 (3):??, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/spawning-
egg-development-and-incubation-
of-cuttlefish-sepia-officinlis-
l-1758-in-controlled-conditions.
php?aid=1465.

Sen:2011:PSF

[Şen11]  Halil Şen.  A prelimi-

nary study on feeding preference of the musky octopus, Eledone moschata, (Cephalopoda: Octopodidae) in Izmir Bay, Aegean Sea. Journal of FisheriesSciences.com, 5(2):??, ???? 2011. CODEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/a-preliminary-
study-on-feeding-preference-
of-the-musky-octopus-eledone-
moschata-cephalopoda-octopodidae-
in-izmir-bay-aegean-sea.
php?aid=768.

Sen:2013:ESD

[Şen13a] Halil Şen. Effects of string diameter preference during Sepia officinalis (L. 1758) spawning in captivity. Journal of FisheriesSciences.com, 7(4):??, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/effects-
of-string-diameter-preference-
of-sepia-officinalis-during-
l-1758-spawning-in-captivity.
php?aid=422.

Sen:2013:KKK

[Şen13b] Halil Şen. Kültür koşullarında kafadanbacaklıların beslenmesi. (Turkish) [Feeding of cephalopods under culture conditions]. Journal of FisheriesSciences.com, 7(3): ??, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/feeding-
of-cephalopods-under-culture-
conditions.php?aid=447.

Sen:2014:KLV

[Şen14] Halil Şen. Kalamar (Loligo vulgaris Lamarck, 1798) ın dış görünüşünden cinsiyet tayini. (Turkish) [Sex distinction by its morphologic view of European squid (Loligo vulgaris Lamarck, 1798). Journal of FisheriesSciences.com, 8(1):??, ???? 2014. CODEN JFOIAQ. ISSN 1307-234X. URL https:/

/www.fisheriessciences.
com/fisheries-aqua/sex-
distinction-by-its-morphologic-
view-of-european-squid-
loligo-vulgaris-lamarck-
1798.php?aid=402.

Serezli:2009:STS

[Ser09] Ramazan Serezli. Spawning time and spawning time repeatability of rainbow trout (Oncorhynchus mykiss). Journal of FisheriesSciences.com, 3(3):??, ???? 2009. CODEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/spawning-
time-and-spawning-time-
repeatability-of-rainbow-
trout-oncorhynchus-mykiss.
php?aid=1461.

Soltani:2016:CDE

[SGH^{+}16] Zohre Soltani, Rasool Ghorbani, Seyed Aliakbar Hedayati, Hamed Ghafari Farsani, and Mohammad Hasan Gerami. Comparative destructive effect of waterborne zinc nanoparticles and zinc sulfate on Capoeta capoeta gracilis hematological indices. Journal of FisheriesSciences.com, 10 (3):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/comparative-
destructive-effect-of-
waterborne-zinc-nanoparticles-
and-zinc-sulfate-on-capoeta-
capoeta-gracilis-hematological-
indices.php?aid=9873.

Sahin:2008:DKB

[ŞHG^{+}08] Cemalettin Şahın, Necati Hacımurtezaoğlu, Ahmet Mutlu Gözler, Ferhat Kalaycı, and Ertuğrul Ağırbaş. Doğu karadeniz bölgesinde girgir ağlarinda hedef dişi av kompozisyonunun araştirilmasi Üzerine bir Ön çalişma. (Turkish) [A preliminary study on investigation of purse seine bycatch composition in the Southeastern Black Sea]. Journal of FisheriesSciences.com, 2(4): 677–683, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/a-prelimenary-
study-on-investigation-
of-purse-seine-bycatch-
composition-in-the-southeastern-
black-sea.php?aid=1548.

Sossoukpe:2016:GME

[SIAF16] Edmond Sossoukpe, Rachad Sidi Imorou, Alphonse Adite, and Emile Didier Fiogbe. Growth, mortality and exploitation of the African lesser threadfin Galeoides decadactylus (Pisces, Polynemidae) fishing by the gill net “Soovi” in Benin nearshore waters. Journal of FisheriesSciences.com, 10(3): ??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/growth-
mortality-and-exploitation-
of-the-african-lesser-
threadfin-galeoides-decadactylus-
pisces-polynemidae-fishing-
by-the-gill-ne.php?aid=
9876.

Sibel:2008:GPR

[Sib08] Ergüden A. Sibel. Growth properties of roach (Rutilus rutilus L., 1758) in Seyhan Dam Lake (Adana). Journal of FisheriesSciences.com, 2(1):??, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/growth-
properties-of-roach-rutilus-
rutilus-l-1758-in-seyhan-
dam-lake-adana.php?aid=
1678.

Siddique:2013:PCC

[Sid13] Mohammad Abdul Momin Siddique. Proximate chemical composition and amino acid profile of two red seaweeds (Hypnea pannosa and Hypnea musciformis) collected from St. Martin’s Island, Bangladesh. Journal of FisheriesSciences.com, 7 (2):??, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/proximate-
chemical-composition-and-
amino-acid-profile-of-
two-red-seaweeds-hypnea-
pannosa-and-hypnea-musciformis-
collected-from-st.php?
aid=513.

Simonovi:2014:ESB

[Sim14] Predrag Simonović. Effects of stocking with brood fish to manage resident stream dwelling brown trout Salmo cf. trutta L. stock. Journal of FisheriesSciences.com, 8(2):??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/effects-
of-stocking-with-brood-
fish-to-manage-resident-
stream-dwelling-brown-
trout-salmo-cf-trutta-l-
stock.php?aid=264.

Singh:2015:AIC

[Sin15] Atul K. Singh. Advances in Indian coldwater fisheries and aquaculture. Journal of FisheriesSciences.com, 9 (3):??, ???? 2015. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/advances-
in-indian-coldwater-fisheries-
and-aquaculture.php?aid=
6890.

Srisakultiew:2013:IMD

[SK13] Penpun Srisakultiew and Wongpathom Kamonrat. Immersion of 17α-methyltestosterone dose & duration on tilapia masculinization. Journal of FisheriesSciences.com, 7(4): 302–308, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/immersion-
of-17945methyltestosterone-
doseduration-on-tilapia-
masculinization.php?aid=
435.

S:2018:AMS

[SKB^{+}18] Adhikari S., Chaudhury Ajit Keshav, Gangadhar Barlaya, Ramesh Rathod, Mandal R. N., Sarosh Ikmail, Saha G. S., De H. K., Sivaraman I., Mahapatra A. S., Sarkar S., Routray P., Bindu R. Pillai, and Sundaray Jk. Adaptation and mitigation strategies of climate change impact in freshwater aquaculture in some states of India. Journal of FisheriesSciences.com, 12(1):??, ???? 2018. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/adaptation-
and-mitigation-strategies-

                                        of-climate-change-impact-
                                        in-freshwater-aquaculture-
                                        in-some-states-of-india.
                                        php?aid=22020.

Singh:2016:CWI

[SKK16] A. K. Singh, Biju Sam Kamalam, and Prem Kumar. Charting ways to invigorate rainbow trout production in India. Journal of FisheriesSciences.com, 10 (2):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/charting-
                                        ways-to-invigorate-rainbow-
                                        trout-production-in-india.
                                        php?aid=9402.

Shehata:2018:ISE

[SMG^{+}18] Sabry M. Shehata, Ramadan A. Mohammed, Mohamed H. Ghanem, Yasser M. Abdelhadi, and Mahmoud Kh. Radwan. Impact of the stresses environmental condition on the prevalence of parasite in fresh water aquaculture. Journal of FisheriesSciences.com, 12 (2):??, ???? 2018. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/impact-
                                        of-the-stresses-environmental-
                                        condition-on-the-prevalence-
                                        of-parasite-in-fresh-water-
                                        aquaculture.php?aid=22615.

Simonetti:2015:ZDR

[SMJO15] Rajla Bressan Simonetti, Lis Santos Marques, Danilo Pedro Streit Jr, and Eneder Rosana Oberst. Zebrafish (Danio rerio): the future of animal model in biomedical research. Journal of FisheriesSciences.com, 9(3): ??, ???? 2015. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/zebrafish-
danio-rerio-the-future-
of-animal-model-inbiomedical-
research.php?aid=6681.

Soyinka:2015:HNT

[SOA^{+}15] Soyinka, Olufemi O., Ayoola, Simeon. O., Ifedayo, and Samuel O. Haematology of Nile tilapia, Oreochromis niloticus fed Mytilus edulis shell meal substituted for dicalcium phosphate. Journal of FisheriesSciences.com, 9(1):??, ???? 2015. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/haematology-
of-nile-tilapia-oreochromis-
niloticus-fed-mytilus-
edulis-shell-meal-substituted-
for-dicalcium-phosphate.
php?aid=5817.

Solomon:2015:ESP

[SOO15] Shola Gabriel Solomon, Victor Tosin Okomoda, and Joshua Andrew Oloche. Evaluation of sweet potato (Ipomea batatas) peel as a replacement for maize meal in the diet of Clarias gariepinus fingerling. Journal of FisheriesSciences.com, 9 (4):??, ???? 2015. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/evaluation-
of-sweet-potato-ipomea-
batatas-peel-as-a-replacement-
for-maize-meal-in-the-
diet-of-clarias-gariepinus-
fingerling.php?aid=8193.

Soylu:2010:EAS

[Soy10] Elif Neyran Soylu. Epipelic algae and seasonal variation of Gıcı Lake (Samsun– Bafra). Journal of FisheriesSciences.com, 4(4):??, ???? 2010. CODEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/epipelic-
algae-and-seasonal-variation-
of-g305c305-lake-samsunbafra.
php?aid=842.

SolisMurgas:2017:UMA

[SPC^{+}17] Luis David Solis Murgas, Michelle Sampaio Paulino, Priscila Cotta Palhares, Alessio Batista Miliorini, Eduardo Alves, and Viviane de Oliveira Felizardo. Ultrastructural and morphometric analysis of gametes in neotropical teleost fishes. Journal of FisheriesSciences.com, 11(1):

??, ????  2017.  CO-

DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/ultrastructural-
and-morphometric-analysis-
of-gametes-in-neotropical-
teleost-fishes.php?aid=
18121.

Su:2015:LBI

[SRB^{+}15] Glenn L. Sia Su, Gliceria B.

Ramos, Elleia Corinne B. Barcelon, Reyna Marie C. Federo, Maria Lilibeth L. Sia Su, and Kimberly Beltran-Benjamin. Lead bioaccumulation and the imposex effect of Volema (Pugilina) cochlidium in Bacoor Bay, Philippines. Journal of FisheriesSciences.com, 9(3): ??, ???? 2015. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/lead-
bioaccumulation-and-the-
imposex-effect-of-volemapugilina-
cochlidium-in-bacoor-bay-
philippines.php?aid=6677.

Sharma:2013:EPH

[SSG13] Parvati Sharma, Ram Chander Sihag, and Suresh Kumar Gahlawat. Effect of probiotic on haematogical parameters of diseased fish (Cirrihinus mrigala). Journal of FisheriesSciences.com, 7(4): 323–328, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/

                                        fisheries-aqua/effect-
                                        of-probiotic-on-haematogical-
                                        paramaters-of-diseased-
                                        fish-cirrihinus-mrigal.
                                        php?aid=432.

Solomon:2015:CGP

[STFO15] S. G. Solomon, L. O. Tiamiyu, A. Fada, and V. T. Okomoda. Comparative growth performance of common carp (Cyprinuscarpio) fry fed dried quail egg and other starter diets in indoor hatchery. Journal of FisheriesSciences.com, 9 (2):??, ???? 2015. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/comparative-
                                        growth-performance-of-
                                        common-carp-cyprinuscarpio-
                                        fry-fed-dried-quail-egg-
                                        and-other-starter-diets-
                                        in-indoor-hatchery.php?
                                        aid=5821.

Stratev:2014:AHS

[Str14] Deyan Stratev. Aeromonas hydrophila sensitivity to disinfectants. Journal of FisheriesSciences.com, 8(4): ??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/aeromonas-
                                        hydrophila-sensitivity-
                                        to-disinfectants.php?aid=
                                        52.

Surengil:2011:GAS

[Sür11]  Göknur Sürengil.  Gida-

ambalaj sektöründe nanoteknolojik uygulamalar ve su Ürünleri açisindan Önemi. Journal of FisheriesSciences.com, 5(??):??, 2011. CODEN JFOIAQ. ISSN 1307-234X.

Sivaci:2008:BAS

[SYG^{+}08] E. Rıdvan Sıvacı, Öztekin Yardım, Arif Gönülol, Levent Bat, and Fatih Gümüsş. Benthic algae of Sarıkum (Sinop– Turkey) Lagoon. Journal of FisheriesSciences.com, 2 (3):??, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/benthic-
algae-of-sarkum-sinopturkey-
lagoon.php?aid=1613.

Tademr:2009:CKC

[Tad09] Aye Tademr. Contributions to the knowledge of Chironomidae (Diptera) fauna of the Gediz Delta (Menemen-İzmir–Turkey). Journal of FisheriesSciences.com, 3(1): ??, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/contributions-
to-the-knowledge-of-chironomidae-
diptera-fauna-of-the-gediz-
delta-menemenzmir-turkey.
php?aid=1506.

Tanrikul:2012:FRR

[Tan12] T. Tansel Tanrıkul. First report on reproduction of Lessepsian ragged sea hare, Bursatella leachii (de Blainville, 1817) (Mollusca: Gastropoda) in Izmir Bay (Aegean Sea, Turkey). Journal of FisheriesSciences.com, 6(2):??, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/first-report-
on-reproduction-of-lessepsian-
ragged-sea-hare-bursatella-
leachii-de-blainville-
1817-mollusca-gastropoda-
in-izmir-bay-aegean-sea-
turkey.php?aid=656.

Tarkan:2008:BRB

[Tar08] Ali Serhan Tarkan. biometric relationship between body size and bone lengths of Carassius gibelio and Rutilus frisii from Iznik Lake. Journal of FisheriesSciences.com, 2(2):??, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/biometric-
relationship-between-body-
size-and-bone-lengths-
of-carassius-gibelio-and-
rutilus-frisii-from-iznik-
lake.php?aid=1646.

Tarkan:2010:ESD

[Tar10] Ali Serhan Tarkan. Effects of streams on drinkable water reservoir: an assessment of water quality, physical habitat and some biological features of the streams. Journal of FisheriesSciences.com, 4(1):??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/effects-
of-streams-on-drinkable-
water-reservoir-an-assessment-
of-water-quality-physical-
habitat-and-some-biological-
features-of-the-streams.
php?aid=1093.

Tas:2007:AVB

[Taş07a]  Beyhan Taş.  Aquacul-

ture in the Vona Bay (Southern Black Sea, Ordu, Turkey). Journal of FisheriesSciences.com, 1(4):176– 183, ???? 2007. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/aquaculture-
in-the-vona-bay-southern-
black-sea-ordu-turkey.
php?aid=1721.

Tas:2007:PAD

[Taş07b] Beyhan Taş. Planktonic algae of Derbent Dam Lake (Samsun, Turkey). Journal of FisheriesSciences.com, 1 (3):111–123, ???? 2007. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/planktonic-
algae-of-derbent-dam-lake-
samsun-turkey.php?aid=
1747.

Tas:2010:SPC

[Taş10a] Beyhan Taş. Some physicochemical features of Lake Ulugöl (Ordu–Turkey). Journal of FisheriesSciences.com, 4(3):??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

                                        //www.fisheriessciences.
                                        com/fisheries-aqua/some-
                                        physicochemical-features-
                                        of-lake-ulugl-orduturkey.
                                        php?aid=1017.

Tasdemir:2010:MOC

[Taş10b] Aye Taşdemır. The macrobenthic (Oligocheta, Chironomidae, Amphipoda) fauna of Tahtali Reservoir (Izmir). Journal of FisheriesSciences.com, 4(4): ??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

                                        //www.fisheriessciences.
                                        com/fisheries-aqua/the-
                                        macrobenthic-oligocheta-
                                        chironomidae-amphipoda-
                                        fauna-of-tahtali-reservoir-
                                        izmir.php?aid=850.

Taskavak:2011:IKE

[Taş11] Ertan Taşkavak. izmir körfezi (ege denizi)’nde gümüş balığı Atherina boyeri Risso, 1810’nın biyometrik Özellikleri. Journal of FisheriesSciences.com, 5(??):??, 2011. CODEN JFOIAQ. ISSN 1307-234X.

Tas:2012:ASA

[Taş12] Menekşe Taş. The aquatic and semi-aquatic Oligochaeta fauna of Turkish Thrace. Journal of FisheriesSciences.com, 6(1):??, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
aquatic-and-semiaquatic-
oligochaeta-fauna-of-turkish-
thrace.php?aid=686.

Taylan:2011:IKN

[Tay11] Burcu Taylan. izmir körfezi’ ndeki gobiidae familyası türlerine ait larvaların bolluk ve dağılımı. Journal of FisheriesSciences.com, 5(??):??, 2011. CODEN JFOIAQ. ISSN 1307-234X.

Teklu:2015:DFW

[Tek15] Demeke Teklu. Determinant factors for wasted fish during harvesting at Amerti and Fichawa Reservoirs Oromia/Ethiopia. Journal of FisheriesSciences.com, 9(4): ??, ???? 2015. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/determinant-
factors-for-wasted-fish-
during-harvesting-at-amerti-
and-fichawa-reservoirs-
oromiaethiopia.php?aid=
8185.

Terbiyik:2008:MCU

[Ter08] Tuba Terbiyik. Morphological characters for using species identification of Chaetognaths. Journal of FisheriesSciences.com, 2(4): ??, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.

                                        fisheriessciences.com/
                                        fisheries-aqua/morphological-
                                        characters-for-using-species-
                                        identification-of-chaetognaths.
                                        php?aid=1546.

Tiwari:2019:YRF

[Tiw19]  Yamini Tiwari.  Young

Researchers Forum 12th Global Summit on Aquaculture & Fisheries March 30–31, 2020 — Hong Kong, Hong Kong. Journal of FisheriesSciences.com, 13 (4):??, ???? 2019. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/young-researchers-
                                        forum-12th-global-summit-
                                        on-aquaculture-fisheries-
                                        march-3031-2020-hong-kong-
                                        hong-kong.pdf.

Tessema:2016:AFS

[TM16] Assefa Tessema and Adem Mohamed. Abundance and fish species composition of Gerado and Dirma Rivers, South Wollo, Ethiopia. Journal of FisheriesSciences.com, 10(2):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/abundance-
                                        and-fish-species-composition-
                                        of-gerado-and-dirma-rivers-
                                        south-wollo-ethiopia.php?
                                        aid=9165.

Tureli:2016:GMB

[TMGY16] Canan Tureli, Thomas J. Miller, Sedat Gundogdu, and Irem Nur Yesilyurt. Growth and mortality of blue crab (Callinectes sapidus) in the North-Eastern Mediterranean Sea. Journal of FisheriesSciences.com, 10 (2):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/growth-
and-mortality-of-blue-
crab-callinectes-sapidus-
in-the-northeastern-mediterranean-
sea.php?aid=9005.

Tantu:2017:TLV

[TNS17] Andi Gusti Tantu, Nurkaidah, and Suryawati Salam. Tempe Lake and various problems. Journal of FisheriesSciences.com, 11(4): ??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/tempe-lake-
and-various-problems.php?
aid=20912.

Tokur:2008:EDT

[Tok08]  Bahar Tokur.  The ef-

fects of different thawing methods on protein quality of frozen fish. Journal of FisheriesSciences.com, 2 (1):??, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
effects-of-different-thawing-
methods-on-protein-quality-

                                        of-frozen-fish.php?aid=
                                        1674.

Topkara:2009:CAI

[Top09] Esat Tark Topkara. Contributions to the aquatic insect (Insecta) fauna of some mountain lakes in the Taurus Range (Turkey). Journal of FisheriesSciences.com, 3(1):??, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/contributions-
                                        to-the-aquatic-insect-
                                        insecta-fauna-of-some-
                                        mountain-lakes-in-the-
                                        taurus-range-turkey.php?
                                        aid=1514.

Tosunoglu:2007:TCD

[Tos07] Zafer Tosunoğlu. Technical characteristics of demersal trawl nets recently used in the Turkish Coast of the Aegean Sea. Journal of FisheriesSciences.com, 1(4): 184–187, ???? 2007. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/technical-
                                        characteristics-of-demersal-
                                        trawl-nets-recently-used-
                                        in-the-turkish-coast-of-
                                        the-aegean-sea.php?aid=
                                        1720.

Tosun:2013:CFP

[Tos13] Deniz D. Tosun. Crocodile farming and its present state in global aquaculture. Journal of FisheriesSciences.com, 7(1):??, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/crocodile-
farming-and-its-present-
state-in-global-aquaculture.
php?aid=549.

Turan:2008:FFL

[TTÇY08] Davut Turan, Beyhan Taş, Muammer Çilek, and Zeynep Yılmaz. Fish fauna of the lower part of River Melet (Ordu, Turkey). Journal of FisheriesSciences.com, 2 (4):698–703, ???? 2008. CODEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/fish-
fauna-of-the-lower-part-
of-river-melet-ordu-turkey.
php?aid=1542.

Turan:2007:FAP

[Tur07] Hulya Turan. Fatty acid profile and proximate composition of the thornback ray (Raja clavata, L. 1758) from the Sinop coast in the Black Sea. Journal of FisheriesSciences.com, 1(2):97– 103, April 2007. CODEN JFOIAQ. ISSN 1307-234X.

Turker:2009:FRI

[Tür09] Hakan Türker. Farklı renklerdeki işıkların nil tilapyası’nın (Oreochromis niloticus L.) büyümesine etkileri. (Turkish) [effect of different color lights on growth of Nile tilapia (Oreochromis niloticus L.)]. Journal of FisheriesSciences.com, 3(3):231– 236, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/effect-
of-different-color-lights-
on-growth-of-nile-tilapia-
-oreochromis-niloticus-
l.php?aid=1454.

Turan:2011:AKC

[Tur11] Funda Turan. Androstenedionun karabalık’ta (Clarias gariepinus (Burchell, 1822)) cinsiyet dönüşümü ve gonad gelişimi üzerine etkileri. (Turkish) [Effects of androstenedione on sex reversal and gonadal development in the African catfish (Clarias gariepinus (Burchell, 1822)]. Journal of FisheriesSciences.com, 5 (2):??, ???? 2011. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/effects-
of-androstenedione-on-
sex-reversal-and-gonadal-
development-in-the-african-
catfish-clarias-gariepinus-
burchell-1822.php?aid=
775.

Tut:2019:DLW

[TWT19] Gatriay Tut, Mulugeta Wakjira, and Girum Tamire. Diversity, length–weight relationship and condition factor of fishes in Gilo River and its nearby wetlands in Akobo District, Gambella Region, Ethiopia. Journal of FisheriesSciences.com, 13 (1):??, ???? 2019. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/diversity-
lengthweight-relationship-
and-condition-factor-of-
fishes-in-gilo-river-and-
its-nearby-wetlands-in-
akobo-district-gambell.
php?aid=24252.

Turk:2013:DBD

[TYB^{+}13] Necla Türk, Murat Yabanlır, Esin Baba, Canan Ontaş, and Mehmet Ali Aydın. Detection of bacterial diseases and determination of antibacterial susceptibilities of rainbow trout (Oncorhynchus mykiss Walbaum, 1792) in Turkey. Journal of FisheriesSciences.com, 7(4):351– 359, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/detection-
of-bacterial-diseases-
and-determination-of-antibacterial-
susceptibilities-of-rainbow-
trout-oncorhynchus-mykiss-
walbaum-1792-in-turkey.
php?aid=426.

Unal:2008:IRL

[ÜA08] Hatice Ünal and Orhan Aral.

Investigation of reproduction and larvae development on peppered cory (Corydoras paleatus Jenys, 1842). Journal of FisheriesSciences.com, 2(1):??, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/investigation-
                                        of-reproduction-and-larvae-
                                        development-on-peppered-
                                        cory-corydoras-paleatus-
                                        jenys1842.php?aid=1671.

Ucar:2013:EAS

[UACÖ13] Arzu Uçar, Muhammed Atamanalp, Murat Cankaya, and Hasan Özdemir. Effects of anesthetic substances on some antioxidan enzyme activities of trouts. Journal of FisheriesSciences.com, 7(2): 152–160, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/effects-
                                        of-anesthetic-substances-
                                        on-some-antioxidan-enzyme-
                                        activities-of-trouts.php?
                                        aid=503.

Uyan:2007:ERA

[UAH^{+}07] Orhan Uyan, Orhan Aral, Fatma Burcu Harmantepe, Simla Uyan, and Muammer Erdem. Effect of raw anchovy as wet feed on growth performances and production cost of rainbow trout (Oncorhynchus mykiss) during winter season in the Black Sea. Journal of FisheriesSciences.com, 1(3):104– 110, ???? 2007. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/effect-
of-raw-anchovy-as-wet-
feed-on-growth-performances-
and-production-cost-of-
rainbow-trout-oncorhynchus-
mykiss-during-winter-se.
php?aid=1748.

Unlusayin:2007:CSA

[ÜBİG07] Mustafa Ünlüsayın, Sengül Bilgin, Levent İzci, and Ani Günlü. Chemical and sensory assessment of hot-smoked fish pâté. Journal of FisheriesSciences.com, 1(1):20– 25, February 2007. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/chemical-
and-sensory-assessment-
of-hotsmoked-fish-pt.php?
aid=1766.

Ukwe:2018:GMI

[UEA18] O. I. K. Ukwe, O. M.

Edun, and O. A. Akinrotimi. Growth and microbial indices in African catfish (Clarias gariepinus) larvae fed formulated and commercial diets. Journal of FisheriesSciences.com, 12 (2):??, ???? 2018. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/growth-
and-microbial-indices-

                                        in-african-catfish-clarias-
                                        gariepinuslarvae-fed-formulated-
                                        and-commercial-diets.php?
                                        aid=22467.

Ugurlu:2007:EFS

[Ugu07] Selma Ugurlu. Exotic fish species inhabiting in freshwater sources within the province of Samsun. Journal of FisheriesSciences.com, 1 (3):139–151, ???? 2007. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/exotic-
                                        fish-species-inhabiting-
                                        in-freshwater-sources-
                                        within-the-province-of-
                                        samsun.php?aid=1745.

Unlusayin:2009:RMU

[Ünl09] Mustafa Ünlüsayın. The recent methods on using for purification of fish proteins. Journal of FisheriesSciences.com, 3(4):??, ???? 2009. CODEN JFOIAQ. ISSN 1307-234X. URL https:/

                                        /www.fisheriessciences.
                                        com/fisheries-aqua/the-
                                        recent-methods-on-using-
                                        for-purification-of-fish-
                                        proteins.php?aid=1163.

Unver:2012:TGS

[Ünv12]  Bülent Ünver.  Tödürge

gölü’ndeki (sivas) cyprinid hibritlerinin [S. cephalus (L.) × A. chalcoides (G.)] üreme biyolojisi. (Turkish) [Reproductive biology of cyprinid hybrids [S. cephalus (L.) × A. chalcoides (G.)] in T˘1693crge Lake (Sivas)]. Journal of FisheriesSciences.com, 6(1): ??, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/reproductive-
biology-of-cyprinid-hybrids-
s-cephalus-l-x-a-chalcoides-
g-in-tdrge-lake-sivas.
php?aid=675.

Urku:2013:CSI

[Ürk13a]  Çidem Ürkü.  Capillaria

sp. infestation and bacterial septicemia in the angel fish (Pterophyllum scalare). Journal of FisheriesSciences.com, 7(3):??, ???? 2013. CODEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/capillaria-
sp-infestation-and-bacterial-
septicemia-in-the-angel-
fish-pterophyllum-scalare.
php?aid=452.

Urku:2013:MBP

[Ürk13b]  Çiğdem Ürkü.  Melek

balıklarında (Pterophyllum scalare) capilaria sp. enfestasyonu ve bakteriyel septisemi. Journal of Fisheries-Sciences.com, 7(??):??, 2013. CODEN JFOIAQ. ISSN 1307-234X.

Uslu:2011:ECC

[Usl11] Leyla Uslu. The effects of the climatic conditions of the Çukurova (Adana–Turkey) on the C-phycocyanin pigments of Spi-rulina platensis (Cyanophyta). Journal of FisheriesSciences.com, 5 (2):??, ???? 2011. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

                                        //www.fisheriessciences.
                                        com/fisheries-aqua/the-
                                        effects-of-the-climatic-
                                        conditions-of-the-ukurova-
                                        adanaturkey-on-the-cphycocyanin-
                                        pigments-of-spirulina-
                                        platensis-cyanophyta.php?
                                        aid=767.

Uslu:2012:BEK

[Usl12a] Leyla Uslu. Besleyici element kompozisyonundaki değişikliklerin mikroalglerde lipid içeriğine etkisi. (Turkish) [The effects of the changes in the composition of nutrient on lipid content of microalgae]. Journal of FisheriesSciences.com, 6(3):??, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

                                        //www.fisheriessciences.
                                        com/fisheries-aqua/the-
                                        effects-of-the-changes-
                                        in-the-composition-of-
                                        nutrient-on-lipid-content-
                                        of-microalgae.php?aid=
                                        610.

Uslu:2012:ELT

[Usl12b] Leyla Uslu. The effects of the light, temperature, nutrient deficiency and aeration on the growth and astaxanthin quantity of Haematococcus pluvialis Flotow. Journal of FisheriesSciences.com, 6(4):??, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
effects-of-the-light-temperature-
nutrient-deficiency-and-
aeration-on-the-growth-
and-astaxanthin-quantity-
of-haematococcus-pluvialis-
flotow.php?aid=574.

Utku:2008:ECC

[Utk08] Güner Utku. Effects of copper and cadmium interaction on total protein levels in liver of Carassius carassius. Journal of FisheriesSciences.com, 2(1):??, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/effects-
of-copper-and-cadmium-
interaction-on-total-protein-
levels-in-liver-of-carassius-
carassius.php?aid=1682.

Uzmanoglu:2008:EAA

[Uzm08] Selçuk Uzmanoğlu. Economic analysis of aquatic products enterprises on the Yene stream in Balkaya-Kırklareli Region. Journal of FisheriesSciences.com, 2 (2):??, ???? 2008. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/economic-
                                        analysis-of-aquatic-products-
                                        enterprises-on-the-yene-
                                        stream-in-balkayak305rklareli-
                                        region.php?aid=1644.

VonEschen:2019:FSM

[VBR19] Aaron J. Von Eschen, Michael L. Brown, and Kurt Rosentrater. Fermented soybean meal as a protein source in diets for yellow perch (Perca flavescens). Journal of FisheriesSciences.com, 13(4):??, ???? 2019. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/fermented-
                                        soybean-meal-as-a-protein-
                                        source-in-diets-for-yellow-
                                        perch-perca-flavescens.
                                        php?aid=25180.

Velloso:2015:USC

[VdFHLAA15] Marcos Antonio Correa Velloso, Lucia de Fatima Henriques Lourenco, Cleidiane Araujo, and Eder Augusto Furtado Araujo. Use of starter cultures with or without nisin in fishburgertype products. Journal of FisheriesSciences.com, 9 (3):??, ???? 2015. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/use-of-
                                        starter-cultures-with-

or-without-nisin-infishburgertype-
products.php?aid=6680.

VanGinneken:2016:WOH

[VdV16] Vincent Van Ginneken and Evert de Vries. Will our health come from our oceans

[in] the 21st century? Journal of FisheriesSciences.com, 10(2):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/will-our-
health-come-from-our-oceans-
the-21st-century.php?aid=
8845.

vanGinneken:2016:TSB

[vGdV16] Vincent van Ginneken and Evert de Vries. Towards a seaweed based economy: the global ten billion people issue at the midst of the 21st Century. Journal of FisheriesSciences.com, 10 (2):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/towards-
a-seaweed-based-economy-
the-global-ten-billion-
people-issue-at-the-midst-
of-the-21st-century.php?
aid=8565.

vanGinneken:2016:SSP

[vGdVW16] Vincent van Ginneken, Evert de Vries, and Tim Wijgerde. A suggested “Seaweed-Plantation model” to tackle the looming phosphorus crises in the 21st Century at the Rhine… North-Sea system. Journal of FisheriesSciences.com, 10 (1):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/a-suggested-
seaweedplantation-model-
to-tackle-the-looming-
phosphorus-crises-in-the-
21st-century-at-the-rhinenorthsea-
system.php?aid=8217.

Vadhel:2017:RTC

[VGP^{+}17] Nirali P. Vadhel, Akolkar N. G., Juned Pathan, Jitendrakumar T. Tandel, Smit Lende, and Vivek Shrivastava. Red tilapia: a candidate euryhaline species for aqua farming in Gujarat. Journal of FisheriesSciences.com, 11(1): ??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/red-
tilapia-a-candidate-euryhaline-
species-for-aqua-farming-
in-gujarat.php?aid=17965.

Voorhees:2020:IDE

[VHKB20] Jill M. Voorhees, Nathan Huysman, Eric Krebs, and Michael E. Barnes. Impacts of different exercise routines on rainbow trout rearing performance. Journal of FisheriesSciences.com, 14(2):??, ???? 2020. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/impacts-
of-different-exercise-
routines-on-rainbow-trout-
rearing-performance.pdf.

Yabanl:2009:MEI

[Yab09a] Murat Yabanl. Monitoring the environmental impacts of marine aquaculture activities on the water column and the sediment in vicinity of the Karaburun Peninsula (Turkey–Eastern Aegean Sea). Journal of FisheriesSciences.com, 3(3): ??, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/monitoring-
the-environmental-impacts-
of-marine-aquaculture-
activities-on-the-water-
column-and-the-sediment-
in-vicinity-of-the-kara.
php?aid=1459.

Yabanli:2009:RIF

[Yab09b]  Murat Yabanli.  A re-

view of integrated finfishseaweed farming. Journal of FisheriesSciences.com, 3 (2):??, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/a-review-
of-integrated-finfishseaweed-
farming.php?aid=1485.

Yaci:2009:RPP

[Yac09] Meral Apaydin Yaci. Reproduction properties of pike (Esox lucius L., 1758) population in the Işıklı Dam Lake (Çivril-Denizli/ Turkey). Journal of FisheriesSciences.com, 3(3):??, ???? 2009. CODEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/reproduction-
                                        properties-of-pike-esox-
                                        lucius-l1758-population-
                                        in-the-ikl-dam-lake-ivrildenizliturkey.
                                        php?aid=1455.

Yelcmen:2011:MCE

[Yel11] H. Özgür Yelçmen. Monthly change of economic fish species caught by bottom trawl fishing from Antalya Bay. Journal of FisheriesSciences.com, 5(2):??, ???? 2011. CODEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/monthly-
                                        change-of-economic-fish-
                                        species-caught-by-bottom-
                                        trawl-fishing-from-antalya-
                                        bay.php?aid=773.

Yesilcimen:2011:AKD

[Yeş11] H. Özgür Yeşilçimen. Antalya körfezi’nde dip trolü balikçiliği ile avlanilan ekonomik balik türlerinin aylik değişimi. Journal of FisheriesSciences.com, 5(??):??, 2011. CODEN JFOIAQ. ISSN 1307-234X.

Yesilsu:2013:UKV

[Yeş13] Ahmet Faruk Yeşilsu. Su Ürünlerinin kalite ve güvenliği için Türkiye ve dünyada uygulanan mevzuatlar. (Turkish) [Turkish and global legislations for seafood safety and quality]. Journal of FisheriesSciences.com, 7(1): ??, ???? 2013. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/turkish-
and-global-legislations-
for-seafood-safety-and-
quality.php?aid=547.

Yuksel:2008:ANO

[YG08a] Eşref Yüksel and Muhammet Gaffaroğlu. The analysis of nucleolar organizer regions in Chalcalburnus mossulensis (Pisces: Cyprinidae). Journal of FisheriesSciences.com, 2(3):587–591, ???? 2008. CODEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
analysis-of-nucleolar-
organizer-regions-in-chalcalburnus-
mossulensis-pisces-cyprinidae.
php?aid=1615.

Yuksel:2008:NPC

[YG08b] Eşref Yüksel and Muhammet Gaffaroğlu. NOR phenotype of Cyprinion macrostomus (Osteichthyes, Cyprinidae). Journal of FisheriesSciences.com, 2(2): 11400116, ???? 2008. CODEN JFOIAQ. ISSN 1307-234X. URL https:

                                        //www.fisheriessciences.
                                        com/fisheries-aqua/nor-
                                        phenotype-of-cyprinion-
                                        macrostomus-osteichthyes-
                                        cyprinidae.php?aid=1638.

Yigit:2011:BYE

[Yiğ11a] Nalan Özgür Yiğit. Balik yemlerinde enzim kullanimi. Journal of FisheriesSciences.com, 5(??):??, 2011. CODEN JFOIAQ. ISSN 1307-234X.

Yigit:2011:KGA

[Yiğ11b] Nalan Özgür Yiğit. Klinoptilolit’in gökkuşaği alabaliği (Oncorhynchus mykiss) yavrularinin büyümesi Üzerine etkisi. Journal of FisheriesSciences.com, 5(??):??, 2011. CODEN JFOIAQ. ISSN 1307-234X.

Yilmaz:2010:EDS

[Yıl10] Hilal Kargın Yılmaz. The effects of different salinity rates on fat and fatty acid composition of Spirulina platensis. Journal of FisheriesSciences.com, 4(??):??, 2010. CODEN JFOIAQ. ISSN 1307-234X.

Yildiz:2011:SPR

[Yıl11] Mustafa Yıldız. A study on profitability of rainbow trout (Oncorhynchus mykiss) farms in Marmara Region, Turkey. Journal of Fisheries-Sciences.com, 5(??):??, 2011. CODEN JFOIAQ. ISSN 1307-234X.

Yildirim:2014:FAA

[Yil14] der Yildirim. Feed additives in aquafeeds. Journal of FisheriesSciences.com, 8(4):??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/feed-
additives-in-aquafeeds.
php?aid=3.

Yesilcicek:2015:LWR

[YKŞ15] Tuncay Yesilcicek, Ferhat Kalayci, and Cemalettin Şahin. Length–weight relationships of 10 fish species from the Southern Black Sea, Turkey. Journal of FisheriesSciences.com, 9(1): ??, ???? 2015. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/lengthweight-
relationships-of-10-fish-
species-from-the-southern-
black-sea-turkey.php?aid=
5818.

Yildirim:2010:MIC

[YÖ10] Sevgi Yıldırım and Selmin Özer. Mersin ili çağlarca köyündeki gökkuşaği alabaliği (Oncorhynchus mykiss, Walbaum, 1792) kuluçkahanelerinde Flavobacterium spp. Varlığı. (Turkish) [The existence of Flavobacterium spp. at rainbow trout (Oncorhynchus mykiss, Walbaum, 1792) hatcheries in Caglarca village, Mersin–Turkey]. Journal of FisheriesSciences.com, 4(1):112–122, ???? 2010. CODEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
existence-of-flavobacterium-
spp-at-rainbow-trout-oncorhynchus-
mykiss-walbaum-1792-hatcheries-
in-caglarca-village-mersinturkey.
php?aid=1077.

Yumi:2016:DPM

[YOH^{+}16] Yumi, Ogata, Hiroki, Koike, Ikuo, Kimura, and Chunhong Yuan. Delaying post-mortem changes in the muscle of spotted mackerel killed by an instantaneous way of neckbreaking and bleeding. Journal of FisheriesSciences.com, 10(2):??, ???? 2016. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/delaying-
postmortem-changes-in-
the-muscle-of-spotted-
mackerel-killed-by-an-
instantaneous-way-of-neckbreaking-
and-bleeding.php?aid=9289.

Yonar:2010:SAS

[Yon10] Muhammet Enis Yonar. A study of Acinetobacter sp. infection in some cultured rainbow trout (Oncorhynchus mykiss) in Kahraman-maras.

Journal of FisheriesSciences.com, 4(4):??, ???? 2010. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/a-study-
of-acinetobacter-sp-infection-
in-some-cultured-rainbow-
trout-oncorhynchus-mykiss-
in-kahramanmaras.php?aid=
870.

Yonar:2014:ISH

[Yon14] Serpil Me Yonar. The investigation of some hematological and antioxi-dant parameters in rainbow trout (Oncorhynchus mykiss) treated formaldehyde. Journal of FisheriesSciences.com, 8(4): ??, ???? 2014. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/the-
investigation-of-some-
hematological-and-antioxidant-
parameters-in-rainbow-
trout-oncorhynchus-mykiss-
treated-formaldehyde.php?
aid=50.

Yilmaz:2007:ABG

[YPY07] Savaş Yılmaz, Nazmi Polat, and Mahmut Yılmaz. Altinkaya baraj gölü (Samsun, Türkiye)’Ndeki sudak baliği (Sander lucioperca (Linnaeus, 1758))’nin yaş tayini için en güvenilir kemiksi yapinin belirlenmesi. (Turkish) [Determination of most reliable bony structure for ageing of pikeperch, Sander lucioperca (Linnaeus, 1758) in Altınkaya Dam Lake (Samsun, Turkey). Journal of FisheriesSciences.com, 1(1): 34–40, February 2007. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/determination-
                                        of-most-reliable-bony-
                                        structure-for-ageing-of-
                                        pikeperch-sander-lucioperca-
                                        linnaeus-1758-in-alt305nkaya-
                                        dam-lake-samsun-turkey.
                                        php?aid=1764.

Yt:2009:IIP

[Yt09]  Murat Yt.  Influence of

increased photoperiods on growth, feed consumption and survival of juvenile mirror carp (Cyprinus carpio Linnaeus, 1758). Journal of FisheriesSciences.com, 3 (2):??, ???? 2009. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/influence-
                                        of-increased-photoperiods-
                                        on-growth-feed-consumption-
                                        and-survival-of-juvenile-
                                        mirror-carp-cyprinus-carpio-
                                        linnaeus-1758.php?aid=
                                        1478.

Yt:2012:UCG

[Yt12] Murat Yt. Utilization of corn gluten meal as a protein source in diets for gilthead sea bream (Sparus aurata L.) juveniles. Journal of FisheriesSciences.com, 6 (1):??, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/utilization-
of-corn-gluten-meal-as-
a-protein-source-in-diets-
for-gilthead-sea-bream-
sparus-aurata-l-juveniles.
php?aid=679.

Yue:2018:EET

[Yue18] Mengmeng Yue. Effects of estradiol and testosterone on the expression of estrogen receptor and androgen receptor genes in female and male Nile tilapia, Oreochromis niloticus. Journal of FisheriesSciences.com, 12(3):??, ???? 2018. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/effects-
of-estradiol-and-testosterone-
on-the-expression-of-estrogen-
receptor-and-androgen-
receptor-genes-in-female-
and-male-nile-t.php?aid=
23063.

Yuksel:2011:KBG

[Yük11]  Fahrettin Yüksel.  Ke-

ban Baraj gölü kerevit (Astacus leptodactylus Eschscholtz, 1823) populasyon büyüklüğünün araştirilmasi. Journal of FisheriesSciences.com, 5(??):??, 2011. CODEN JFOIAQ. ISSN 1307-234X.

Yuksel:2012:ISM

[Yük12] Fahrettin Yüksel. An investigation on some morphological characteristics of crayfish in Keban Dam Lake. Journal of FisheriesSciences.com, 6(4):??, ???? 2012. CO-DEN JFOIAQ. ISSN 1307-234X. URL https:

//www.fisheriessciences.
com/fisheries-aqua/an-
investigation-on-some-
morphological-characteristics-
of-crayfish-in-keban-dam-
lake.php?aid=582.

Yildiz:2007:PFS

[YV07] Hijran Yavuzcan Yıldız and Erdinç Veske. Pain in fish: Study on pain in guppy (Poecilia reticulata) with a painkiller. Journal of FisheriesSciences.com, 1(4):188– 190, ???? 2007. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/pain-in-
fish-study-on-pain-in-
guppy-poecilia-reticulata-
with-a-painkiller.php?
aid=1718.

Yilmaz:2007:BBG

[YYBP07] Mahmut Yılmaz, Savas Yılmaz, Derya Bostanci, and Nazmi Polat. Bafra balik golleri’nde yasayan havuz baligi (Carassius gibelio, Bloch 1782)’nin beslenme rejimi. (Turkish) [Feeding dietary of Prussian carp (Carassius gibelio, Bloch 1782) inhabiting Bafra fish lakes (Samsun)]. Journal of FisheriesSciences.com, 1(2): 48–57, April 2007. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/bafra-balik-
gollerinde-yasayan-havuzbaligi-
carassius-gibelio-bloch-
1782ninbeslenme-rejimi.
php?aid=1753.

Zlateva:2017:MBL

[Zla17] Ivelina Zlateva. Matlabbased length/weight relationship analysis of commercial fishery samples taken from the Black Sea (Bulgaria). Journal of FisheriesSciences.com, 11(4): ??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/matlabbased-
lengthweight-relationship-
analysis-of-commercial-
fishery-samples-taken-
from-the-black-sea-bulgaria.
php?aid=20582.

Zoe:2011:CEP

[Zoe11]  Roupa Zoe.  Cumula-

tive effect of psychotherapy in remission of symptomatology of major depressive disorder. Journal of FisheriesSciences.com, 5 (2):??, ???? 2011. CO-DEN JFOIAQ. ISSN 1307-

234X. URL https://www.
fisheriessciences.com/
fisheries-aqua/cumulative-
effect-of-psychotherapy-
in-remission-of-symptomatology-

                                        of-major-depressive-disorder.
                                        php?aid=780.

Zhang:2017:SRR

[ZPW^{+}17] Yuan-Wei Zhang, Xiao-Fu Pan, Xiao-Ai Wang, Wan-Sheng Jiang, Kun-Feng Yang, Qian Liu, and Jun-Xing Yang. Selection of reliable reference genes for quantitative real-time PCR in golden-line barbell (Sinocyclocheilus grahami) during juvenile and adult stages. Journal of FisheriesSciences.com, 11 (2):??, ???? 2017. CO-DEN JFOIAQ. ISSN 1307-

                                        234X. URL https://www.
                                        fisheriessciences.com/
                                        fisheries-aqua/selection-
                                        of-reliable-reference-
                                        genes-for-quantitative-
                                        realtime-pcr-in-goldenline-
                                        barbell-sinocyclocheilus-
                                        grahami-during-juvenile.
                                        php?aid=19078.